### v1

# Conditional PED-ANOVA: Hyperparameter Importance in Hierarchical & Dynamic Search Spaces

Abstract.
We propose conditional PED-ANOVA (condPED-ANOVA), a principled framework for estimating hyperparameter importance (HPI) in conditional search spaces, where the presence or domain of a hyperparameter can depend on other hyperparameters. Although the original PED-ANOVA provides a fast and efficient way to estimate HPI within the top-performing regions of the search space, it assumes a fixed, unconditional search space and therefore cannot properly handle conditional hyperparameters. To address this, we introduce a conditional HPI for top-performing regions and derive a closed-form estimator that accurately reflects conditional activation and domain changes. Experiments show that naive adaptations of existing HPI estimators yield misleading or uninterpretable importance estimates in conditional settings, whereas condPED-ANOVA consistently provides meaningful importances that reflect the underlying conditional structure.
Analysis of Variance, Hyperparameter Importance, Conditional Search Space, Automated Machine Learning
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Machine learning approachesccs: Mathematics of computing Mathematical analysis
1. Introduction
Analysis of parameter importance (Sobol, 1993, 2001; Hooker, 2007; Huang, 1998) is a central tool for understanding and designing optimization procedures. Over the past decade, techniques for measuring the importance of parameters have attracted considerable attention in the automated machine learning (AutoML) community (Hutter et al., 2014; Biedenkapp et al., 2017; Van Rijn and Hutter, 2018), particularly for improving the efficiency of hyperparameter optimization (HPO) (Bergstra et al., 2011; Bergstra and Bengio, 2012; Snoek et al., 2012; Lindauer et al., 2022). Given a hyperparameter search space and a set of evaluated configurations, methods for hyperparameter importance (or HPI methods for short) aim to quantify how strongly each hyperparameter influences the observed performance. Such information is useful for interpreting optimization runs, identifying critical design choices, detecting irrelevant or redundant parameters, and guiding subsequent search or manual tuning (Melis et al., 2018; Henderson et al., 2018). As a result, HPI is widely used across various settings, including HPO of machine learning models and HPO algorithm design, where understanding hyperparameter sensitivity is often essential (Van Rijn and Hutter, 2018; Zimmer et al., 2021; Probst et al., 2019; Theodorakopoulos et al., 2024).
Classic HPI methods such as functional analysis of variance (f-ANOVA) (Hutter et al., 2014) quantify HPI by measuring how much variance in the objective can be attributed to each hyperparameter. However, their existing local variants mix contributions from unpromising regions, leading to inaccurate estimates of HPI in top-performing subspaces (Hutter et al., 2014; Biedenkapp et al., 2019). PED-ANOVA (Watanabe et al., 2023) mitigates this by defining local HPI within a top-performing subset of the search space and efficiently approximating it using the Pearson divergence between one-dimensional distributions of each hyperparameter. Moreover, PED-ANOVA computes both global and local HPIs substantially faster than f-ANOVA while maintaining accurate HPI estimates. Thanks to this combination of speed and reliability, it has been adopted for HPI analysis in practical HPO frameworks, such as Optuna (Akiba et al., 2019), demonstrating its utility in real-world workflows.
However, we argue that PED-ANOVA has a practical limitation: it cannot properly handle hierarchical/dynamic search spaces, where the presence/domain of a hyperparameter depends on other hyperparameters (hereafter referred to as conditional search spaces). In practice, many real-world problems involve such conditional search spaces (Hutter et al., 2019). Typical examples include model selection, where hyperparameters are dependent on a chosen algorithm (Thornton et al., 2013; Kotthoff et al., 2017; Feurer et al., 2015; Komer et al., 2014), and neural network architecture design, where active parameters depend on the chosen layers (Bergstra et al., 2013; Swersky et al., 2014). Unfortunately, PED-ANOVA cannot be directly applied to such settings, since it assumes a fixed, unconditional search space and defines distributions over a single, fixed domain for each hyperparameter. We observe that naive extensions of PED-ANOVA, such as filtering inactive configurations, imputing missing values, or artificially expanding parameter domains, do not resolve this issue: these heuristics often violate the conditional semantics of the search space and yield unnatural or uninterpretable HPI estimates (cf. Section 5.4).
To address this limitation, we propose conditional PED-ANOVA (condPED-ANOVA), a principled framework for estimating HPI in conditional search spaces. We identify a critical flaw in the original definition of local HPI in PED-ANOVA when applied to conditional search spaces: the importance of hyperparameters that induce conditionality is spuriously attributed to the hyperparameters whose domains or activation depend on them, leading to systematically misleading importance estimates. Consequently, merely constructing a method that correctly computes local HPI in conditional search spaces is still insufficient. We introduce conditional local HPI, which measures importance within top-performing regions without confounding from conditional activation or domain changes. Building on this definition, we derive a closed-form PED-ANOVA-style estimator that remains efficient and yields an interpretable HPI in conditional settings. Experiments confirm that condPED-ANOVA reliably produces a meaningful HPI that reflects the underlying conditional structure, while naive extensions of existing methods fail to do so.
Our contributions are summarized as follows:
  • We identify a fundamental pitfall of original local HPI in conditional search spaces: the contribution of conditioning hyperparameters is spuriously attributed to conditionally active (or domain-shifted) hyperparameters, yielding misleading importance estimates.
  • We introduce conditional local HPI, a definition that measures importance within top-performing regions while accounting for conditional activation and domain changes.
  • We derive a closed-form PED-ANOVA-style estimator for conditional local HPI, retaining the efficiency of the original PED-ANOVA via Pearson divergence of one-dimensional distributions estimated with KDE.
  • We empirically demonstrate that condPED-ANOVA consistently produces meaningful importances in conditional settings, whereas naive extensions of existing methods, including PED-ANOVA, produce unnatural results.
2. Related Work
Here, we review prior work most relevant to ours. Additional related work is provided in Appendix A.
2.1. Variance-Based HPI and f-ANOVA
HPI aims to quantify how strongly each hyperparameter (and their interactions) influences performance, typically to support search-space design, debugging, and interpretation of HPO runs. Early work on parameter importance analysis builds on variance-based sensitivity analysis such as Sobol indices (Sobol, 1993, 2001) and ANOVA decompositions (Hooker, 2007; Huang, 1998), which attribute the global variance of an objective to individual coordinates and their interactions. In the context of HPO, f-ANOVA (Hutter et al., 2014) applies this idea to HPO objectives by fitting a surrogate model such as a random forest over the hyperparameter space and then estimating main and interaction effects by Monte Carlo integration over the surrogate. This yields importance values that reflect how much of the global variation in performance across the entire search space can be explained by each hyperparameter. f-ANOVA includes efficient marginal computation over random-forest predictions and produces a global ranking of important hyperparameters, making it a standard post-hoc analysis tool in AutoML and algorithm configuration.
2.2. PED-ANOVA
PED-ANOVA (Watanabe et al., 2023) extends HPI to settings where the region of interest is not an entire search space. It defines local HPI as the variance contribution of each hyperparameter within the top-$\gamma$ subset of the search space, thereby measuring importance directly in the most relevant region. It further shows that this local HPI can be approximated in closed form using the Pearson divergence between the one-dimensional distributions of each hyperparameter of two top-performing subsets, estimated via one-dimensional KDEs. Empirically, this leads to much faster HPI computation than f-ANOVA while maintaining accurate importance estimates in top-performing subspaces. The details of PED-ANOVA are reviewed in Section 3.
Our approach follows a PED-ANOVA-style framework, while enabling effective HPI estimation in conditional search spaces that arise in many real-world problem settings.
2.3. Conditional Hyperparameters in HPO
Many HPO algorithms furthermore operate on conditional search spaces, where certain hyperparameters are active only under specific parent choices, and their domains may also change depending on those choices (e.g., algorithm selection or internal mode switches). This view is made explicit in practical AutoML systems, such as Auto-WEKA (Thornton et al., 2013), auto-sklearn (Feurer et al., 2015, 2022), which describe conditional hyperparameters that are only relevant when a parent hyperparameter takes a particular value. Among widely used optimizers, SMAC (Lindauer et al., 2022) is designed to handle such conditional hierarchies in CASH-style spaces (Thornton et al., 2013), while TPE (Bergstra et al., 2011) formulates the search space as a tree-structured generative process where some variables are only well-defined under specific upstream choices. BOHB (Falkner et al., 2018) combines Hyperband (Li et al., 2018) with a Bayesian-optimization component built on TPE, thereby optimizing over the same tree-structured conditional spaces while exploiting cheap low-fidelity evaluations. Beyond these algorithms, a complementary line of work develops BO surrogates tailored to conditional domains to better share information across branches (Swersky et al., 2014; Jenatton et al., 2017; Ma and Blaschko, 2020).
2.4. HPI under Conditional Search Spaces
In conditional search spaces, applying standard HPI pipelines generally requires mapping configurations to a fixed representation, typically by imputing inactive hyperparameters with default values as done in SMAC-style toolchains and related preprocessing workflows (Lindauer et al., 2022). Practical analysis suites such as CAVE (Biedenkapp et al., 2019) (and its integration in BOAH (Lindauer et al., 2019)) then compute importance post-hoc using methods like f-ANOVA and local parameter importance, including for multi-fidelity runs. Optuna’s importance API highlights a common limitation of conditional spaces: by default, it evaluates only parameters that appear and have the same domain in all samples, excluding conditional parameters (Akiba et al., 2019). Recent work on multi-objective HPI likewise treats missing values arising from conditional hyperparameters in preprocessing before applying surrogate-based importance measures such as f-ANOVA and ablation paths (Theodorakopoulos et al., 2024).
Consequently, despite the fact that conditional hyperparameters arise frequently in practical HPO settings, prior HPI methods have lacked a principled way to handle conditional hyperparameters directly, without resorting to default-value imputation or discarding them during preprocessing. This limitation has prevented existing approaches from faithfully capturing the importance of hyperparameters within the conditional structure of the search space.
3. Preliminaries
3.1. Notations
Let $\mathcal{X}\coloneqq\mathcal{X}^{(1)}\times\cdots\times\mathcal{X}^{(D)}$ be the $D$-dimensional search space for hyperparameters $\bm{x}\coloneqq(x^{(1)},\ldots,x^{(D)})\in\mathcal{X}$. Let $f\colon\mathcal{X}\to\mathbb{R}$ be an objective function to be minimized. We observe a finite evaluation set $\{(\bm{x}_{n},f(\bm{x}_{n}))\}_{n=1}^{N}$, where each $\bm{x}_{n}$ is sampled from a distribution $p_{0}(\bm{x})$. For $\gamma\in(0,1]$, let $f_{\gamma}\in\mathbb{R}$ denote the lower (i.e., better) empirical $\gamma$-quantile of the observed values, i.e., $f_{\gamma}$ satisfies $\{n\mid f(\bm{x}_{n})\leq f_{\gamma}\}=\lfloor\gamma N\rfloor.$ The corresponding top-$\gamma$ set is represented as $\mathcal{X}_{\gamma}:=\{\bm{x}\in\mathcal{X}\mid f(\bm{x})\leq f_{\gamma}\}.$
3.2. Local HPI
To quantify how strongly the $d$-th hyperparameter influences the performance in the top-$\gamma$ set, a local HPI based on a local marginal variance is introduced (Watanabe et al., 2023). Let $b_{\gamma}(\bm{x})\coloneqq\mathbf{1}\{\bm{x}\in\mathcal{X}_{\gamma}\}=\mathbf{1}\{f(\bm{x})\leq f_{\gamma}\}$ be the indicator function of whether hyperparameters $\bm{x}$ belong to the top-$\gamma$ performers. The local mean $m_{\gamma}$ of the objective and the local marginal mean $f_{\gamma}^{(d)}(x^{(d)})$ for the $d$-th hyperparameter $x^{(d)}$ at level $\gamma$ are defined as:
$m_{\gamma}$ $\coloneqq\mathbb{E}_{\bm{x}\sim p_{0}}[f(\bm{x})\mid b_{\gamma}(\bm{x})]=\frac{\mathbb{E}_{\bm{x}\sim p_{0}}[f(\bm{x})\,b_{\gamma}(\bm{x})]}{\mathbb{E}_{\bm{x}\sim p_{0}}[b_{\gamma}(\bm{x})]},$
$f_{\gamma}^{(d)}(x^{(d)})$ $\coloneqq\mathbb{E}_{\bm{x}^{(\bar{d}\,)}\sim p_{0}^{(\bar{d}\,)}}[f(\bm{x})\mid b_{\gamma}(\bm{x}),x^{(d)}]$
$\,=\frac{\mathbb{E}_{\bm{x}^{(\bar{d}\,)}\sim p_{0}^{(\bar{d}\,)}}[f(\bm{x})\,b_{\gamma}(\bm{x})\mid x^{(d)}]}{\mathbb{E}_{\bm{x}^{(\bar{d}\,)}\sim p_{0}^{(\bar{d}\,)}}[b_{\gamma}(\bm{x})\mid x^{(d)}]},$
where $\bm{x}^{(\bar{d}\,)}\!\coloneqq\!(x^{(1)},\ldots,x^{(d-1)},x^{(d+1)},\ldots,x^{(D)})$ denotes all hyperparameters except for the $d$-th one. The local marginal mean $f_{\gamma}^{(d)}(x^{(d)})$ is obtained by taking the mean over all hyperparameters $\bm{x}^{(\bar{d}\,)}$ within the top-$\gamma$ region while keeping the $d$-th hyperparameter fixed at $x^{(d)}$, in a manner similar to marginalization.
The local marginal variance for the $d$-th hyperparameter at level $\gamma$ is then defined as follows:
$v_{\gamma}^{(d)}$ $\coloneqq\mathrm{Var}_{\bm{x}\sim p_{0}}\bigl(f_{\gamma}^{(d)}(x^{(d)})\mid b_{\gamma}(\bm{x})\bigr)$
$\,=\frac{\mathbb{E}_{\bm{x}\sim p_{0}}[(f_{\gamma}^{(d)}(x^{(d)})-m_{\gamma})^{2}\,b_{\gamma}(\bm{x})]}{\mathbb{E}_{\bm{x}\sim p_{0}}[b_{\gamma}(\bm{x})]}.$
Using this quantity, the local HPI for the $d$-th hyperparameter at level $\gamma$ is then defined as its normalized form:
$\text{Local HPI}\coloneqq v_{\gamma}^{(d)}\Big/\sum_{d^{\prime}=1}^{D}v_{\gamma}^{(d^{\prime})}.$
3.3. PED-ANOVA
PED-ANOVA (Watanabe et al., 2023) introduces an efficient approximation of the local HPI of the objective $f$ by computing the local HPI of a tighter level-set indicator function $b_{\gamma^{\prime}}(\bm{x})\coloneqq\mathbf{1}\{\bm{x}\in\mathcal{X}_{\gamma^{\prime}}\},$ where $0<\gamma^{\prime}<\gamma$. A key result in PED-ANOVA shows that the local marginal variance for $b_{\gamma^{\prime}}$ admits the following closed-form expression:
(1) $v_{\gamma}^{(d)}=\left(\frac{\gamma^{\prime}}{\gamma}\right)^{\!2}D_{\mathrm{PE}}\left(p_{\gamma^{\prime}}^{(d)}\,\middle\|\,p_{\gamma}^{(d)}\right),$
where $p_{\gamma}^{(d)}$ and $p_{\gamma^{\prime}}^{(d)}$ are the empirical one-dimensional probability density functions (PDFs) of the $d$-th hyperparameter induced by the samples in the top-$\gamma$ and top-$\gamma^{\prime}$ subsets, respectively. In practice, the PDFs are estimated via KDE using the observed samples. Here, $D_{\mathrm{PE}}(p\|q)$ denotes the Pearson ($\chi^{2}$) divergence between two PDFs $p$ and $q$, defined as:
$D_{\mathrm{PE}}(p\|q)\coloneqq\int\left(\frac{p(x)}{q(x)}-1\right)^{2}q(x)\mathrm{d}x.$
Note that this divergence is well-defined in the present setting, since $\mathcal{X}_{\gamma^{\prime}}\subset\mathcal{X}_{\gamma}$ implies that the density $p_{\gamma^{\prime}}^{(d)}$ has support only where $p_{\gamma}^{(d)}$ is nonzero, so the ratio $p_{\gamma^{\prime}}^{(d)}/p_{\gamma}^{(d)}$ is well-defined.
PED-ANOVA uses this closed-form expression to compute the local HPI solely from one-dimensional PDFs. In the original PED-ANOVA experiments, this approach was shown to yield dramatically faster HPI calculation compared to f-ANOVA, while maintaining comparable quality in the resulting importance measures.
4. condPED-ANOVA: Hyperparameter Importance in Conditional Search Spaces
In this section, we present condPED-ANOVA, a principled approach for efficient HPI computation in conditional search spaces. Rigorous details are provided in Appendix B.
4.1. Regime-Based Representation of Conditional Hyperparameters
In practice, many hyperparameters are conditional: the presence of a hyperparameter and its domain can change depending on other hyperparameters (Hutter et al., 2019; Thornton et al., 2013; Kotthoff et al., 2017; Feurer et al., 2015; Komer et al., 2014; Bergstra et al., 2013; Swersky et al., 2014).
We now formalize such behavior. To handle such conditional structure, we partition the behavior of the $d$-th hyperparameter into a finite number of regimes. Each regime $i^{(d)}\in\{1,\ldots,K^{(d)}\}$ has its own domain $\mathcal{Z}_{i}^{(d)}$ for $x^{(d)}$, which may differ across regimes. When $x^{(d)}$ is inactive in regime $i^{(d)}$, we represent this by a special symbol $\bot$ and set $\mathcal{Z}_{i}^{(d)}=\{\bot\}$; otherwise, $\mathcal{Z}_{i}^{(d)}$ is the usual domain (e.g., an interval or a discrete set), whose range may depend on $i^{(d)}$.
For example, suppose $x^{(0)}$ selects a learning algorithm, such as a neural network or a tree-based model. If $x^{(0)}=\texttt{neural}$, $x^{(1)}$ may correspond to the number of layers with domain $\mathcal{Z}^{(1)}_{\texttt{neural}}=\{1,\ldots,L_{\max}\}$, while $x^{(2)}$ is inactive, i.e., $\mathcal{Z}^{(2)}_{\texttt{neural}}$$\{\bot\}$. Conversely, if $x^{(0)}=\texttt{tree}$, $x^{(1)}$ is inactive, i.e., $\mathcal{Z}^{(1)}_{\texttt{tree}}$$\{\bot\}$, and $x^{(2)}$ may represent the minimum split gain with domain $\mathcal{Z}^{(2)}_{\texttt{tree}}$$\mathbb{R}_{+}$. Here, the choice of $x^{(0)}$ induces regimes $i^{(d)}\in\{\texttt{neural},\texttt{tree}\}$ under which $x^{(1)}$ and $x^{(2)}$ have regime-specific domains or become inactive.
Algorithm 1 condPED-ANOVA Computation for the $d$-th Hyperparameter
1:Evaluation set $\mathcal{D}=\{(\bm{x}_{n},f(\bm{x}_{n}))\}_{n=1}^{N}$, hyperparameter index $d$, quantile levels $0<\gamma^{\prime}\leq\gamma\leq 1$
2:The within-regime local marginal variance $v_{\gamma,\mathrm{within}}^{(d)}$ for the $d$-th hyperparameter $x^{(d)}$
3:Compute empirical $\gamma^{\prime}$- and $\gamma$- quantiles $f_{\gamma^{\prime}}$ and $f_{\gamma}$ of $\{f(\bm{x}_{n})\}_{n=1}^{N}$
4:Define top-performing subsets: $\mathcal{D}_{\gamma^{\prime}}\leftarrow\{(\bm{x}_{n},f(\bm{x}_{n}))\mid f(\bm{x}_{n})\leq f_{\gamma^{\prime}}\}$, $\mathcal{D}_{\gamma}\leftarrow\{(\bm{x}_{n},f(\bm{x}_{n}))\mid f(\bm{x}_{n})\leq f_{\gamma}\}$
5:for $i=1,\ldots,K^{(d)}$ do $\triangleright$ process each regime $i$
6:  $\mathcal{D}_{\gamma^{\prime},\,i}^{(d)}\,,\,\mathcal{D}_{\gamma,\,i}^{(d)}\leftarrow$ Extract samples in regime $i$ from $\mathcal{D}_{\gamma^{\prime}}$ and $\mathcal{D}_{\gamma}$
7:  $p_{\gamma^{\prime},\,i}^{(d)}\,,\,p_{\gamma,\,i}^{(d)}\leftarrow$ Estimate via one-dimensional KDE from the samples $\mathcal{D}_{\gamma^{\prime},\,i}^{(d)}$ and $\mathcal{D}_{\gamma,\,i}^{(d)}$
8:  $L^{\text{within}}_{i}\leftarrow D_{\mathrm{PE}}\left(p_{\gamma^{\prime},\,i}^{(d)}\,\middle\|\,p_{\gamma,\,i}^{(d)}\right)\triangleright$ within-regime Pearson divergence
9:  $\alpha_{i}^{(d)}\leftarrow|\mathcal{D}_{\gamma^{\prime},\,i}|\,/\,|\mathcal{D}_{\gamma^{\prime}}|,\;\beta_{i}^{(d)}\leftarrow|\mathcal{D}_{\gamma,\,i}|\,/\,|\mathcal{D}_{\gamma}|$ $\triangleright$ regime probabilities in top $\gamma^{\prime}$ and $\gamma$ subsets
10:end for
11:return $\left(\frac{\gamma^{\prime}}{\gamma}\right)^{2}\sum_{i=1}^{K^{(d)}}\frac{\big(\alpha_{i}^{(d)}\big)^{2}}{\beta_{i}^{(d)}}L^{\text{within}}_{i}$ $\triangleright$ weighted aggregation of within-regime divergences
4.2. Conditional Local HPI via Within-Regime Variance
Let $I^{(d)}$ and $Z^{(d)}$ denote random variables taking values in $\{1,\ldots,\allowbreak K^{(d)}\}$ and $\mathcal{Z}^{(d)}_{I^{(d)}}$, respectively. We write the local marginal mean $f_{\gamma}^{(d)}(x^{(d)})$ as $g_{\gamma}^{(d)}(I^{(d)},Z^{(d)}).$
We observe that the standard local HPI based on the local marginal variance $v_{\gamma}^{(d)}$ mixes two distinct sources of variation: differences within each regime and differences between (inter) regimes. This can be seen through the following variance decomposition:
(2) $v_{\gamma}^{(d)}$ $=\mathrm{Var}_{I^{(d)},\,Z^{(d)}}\left(g_{\gamma}^{(d)}(I^{(d)},Z^{(d)})\right)$
$=\underbrace{\mathbb{E}_{I^{(d)}}\left[\mathrm{Var}_{Z^{(d)}}\left(g_{\gamma}^{(d)}(I^{(d)},Z^{(d)})\,\middle|\,I^{(d)}\right)\right]}_{\text{within-regime variance}}$
$\hskip 25.00003pt+\underbrace{\mathrm{Var}_{I^{(d)}}\left(\mathbb{E}_{Z^{(d)}}\left[g_{\gamma}^{(d)}(I^{(d)},Z^{(d)})\,\middle|\,I^{(d)}\right]\right)}_{\text{inter-regime variance}},$
where the expectations and variances are taken under the empirical distribution restricted to the top-$\gamma$ region $\mathcal{X}_{\gamma}$.
The inter-regime variance captures differences across regimes, weighted by their frequency in the top-$\gamma$ region. However, for conditional hyperparameters, the regime $I^{(d)}$ is determined by other hyperparameters (e.g., parent choices) rather than by the $d$-th hyperparameter itself. As a result, attributing this inter-regime effect to the $d$-th hyperparameter is often inappropriate: it can reflect changes in upstream decisions rather than the effect of the value of $x^{(d)}$ itself. In fact, we can show that the standard local marginal variance of a conditioned hyperparameter includes the conditioning variable’s local marginal variance, and is therefore always no smaller than it (cf. Theorem E.1 in Section E.1). In particular, this leakage implies that even inactive hyperparameters can be assigned non-negligible importance solely due to upstream choices. Such behavior is undesirable because tuning an inactive hyperparameter has no effect on the objective and is therefore meaningless.
Motivated by this consideration, we define the local HPI for conditional hyperparameters using only the within-regime component:
Definition 4.1 (Conditional Local HPI).
For a conditional hyperparameter $x^{(d)}$, we define the conditional local HPI at level $\gamma$ as:
(3) $\text{Conditional Local HPI}\coloneqq v_{\gamma,\mathrm{within}}^{(d)}\Big/\sum_{d^{\prime}=1}^{D}v_{\gamma,\mathrm{within}}^{(d^{\prime})},$
where $v_{\gamma,\mathrm{within}}^{(d)}$ is the within-regime local marginal variance defined as:
(4) $v_{\gamma,\mathrm{within}}^{(d)}\coloneqq\mathbb{E}_{I^{(d)}}\left[\mathrm{Var}_{Z^{(d)}}\left(g_{\gamma}^{(d)}(I^{(d)},Z^{(d)})\,\middle|\,I^{(d)}\right)\right].$
Intuitively, $v_{\gamma,\mathrm{within}}^{(d)}$ quantifies how much the value of the $d$-th hyperparameter changes the likelihood of being in the tighter top set $\mathcal{X}_{\gamma^{\prime}}$, after fixing the regime. When the hyperparameter is not conditional (i.e., $K^{(d)}=1$), the inter-regime term in Equation 2 vanishes and $v_{\gamma,\mathrm{within}}^{(d)}$ reduces to the standard local HPI. The empirical difference in HPI performance between $v_{\gamma,\mathrm{within}}^{(d)}$ and the original $v_{\gamma}^{(d)}$ is examined in detail in Section 5.5.
4.3. condPED-ANOVA
We now derive a closed-form condPED-ANOVA estimator that efficiently computes the within-regime local marginal variance of each hyperparameter from one-dimensional PDFs within each regime. The conditional local HPI then follows immediately by normalization, as defined in Equation 3. The pseudocode is provided in Algorithm 1.
For the $d$-th hyperparameter, let $\mathcal{D}_{\gamma^{\prime}}\!\coloneqq\!\{(\bm{x}_{n},f(\bm{x}_{n}))\!\mid\!f(\bm{x}_{n})\!\leq f_{\gamma^{\prime}}\}$ and $\mathcal{D}_{\gamma}\!\coloneqq\!\{(\bm{x}_{n},f(\bm{x}_{n}))\!\mid\!f(\bm{x}_{n})\!\leq f_{\gamma}\}$ denote the top-$\gamma^{\prime}$ and top-$\gamma$ evaluation subsets, respectively. For each regime $i^{(d)}\!\in\!\{1,\ldots,K^{(d)}\}$, let $\mathcal{D}_{\gamma^{\prime},\,i}^{(d)}$ and $\mathcal{D}_{\gamma,\,i}^{(d)}$ be the samples in regime $i^{(d)}$ within $\mathcal{D}_{\gamma^{\prime}}$ and $\mathcal{D}_{\gamma}$, respectively. We define the regime probabilities:
$\alpha_{i}^{(d)}\coloneqq\frac{|\mathcal{D}_{\gamma^{\prime},\,i}^{(d)}|}{|\mathcal{D}_{\gamma^{\prime}}|},\quad\text{and}\quad\beta_{i}^{(d)}\coloneqq\frac{|\mathcal{D}_{\gamma,\,i}^{(d)}|}{|\mathcal{D}_{\gamma}|}.$
For each regime $i^{(d)}$, let $p_{\gamma^{\prime},\,i}^{(d)}$ and $p_{\gamma,\,i}^{(d)}$ denote the one-dimensional PDFs of the $d$-th hyperparameter. In practice, these PDFs are estimated using KDE from the samples in $\mathcal{D}_{\gamma^{\prime},\,i}^{(d)}$ and $\mathcal{D}_{\gamma,\,i}^{(d)}$, respectively.
With these definitions, we obtain the following closed-form expression for the within-regime local marginal variance:
Theorem 4.2 (condPED-ANOVA).
Let $0<\gamma^{\prime}<\gamma\leq 1$. The within-regime local marginal variance for the $d$-th hyperparameter at level $\gamma$, computed using the indicator function $b_{\gamma^{\prime}}\coloneqq\mathbf{1}\{x\in\mathcal{X}_{\gamma^{\prime}}\}$, is given by:
(5) $v_{\gamma,\mathrm{within}}^{(d)}=\left(\frac{\gamma^{\prime}}{\gamma}\right)^{\!2}\sum_{i=1}^{\;K^{(d)}}\frac{\big(\alpha_{i}^{(d)}\big)^{2}}{\beta_{i}^{(d)}}D_{\mathrm{PE}}\left(p_{\gamma^{\prime},\,i}^{(d)}\,\middle\|\,p_{\gamma,\,i}^{(d)}\right).$
By normalizing the variance across all hyperparameters as defined in Equation 3, we obtain the conditional local HPI for the $d$-th hyperparameter.
The proof of Theorem 4.2 is provided in Appendix B. Equation 5 aggregates within-regime distributional changes using regime-dependent weights $\big(\alpha_{i}^{(d)}\big)^{2}/\beta_{i}^{(d)}$. Intuitively, this weighting emphasizes regimes that are enriched in the tighter top set (i.e., with larger $\big(\alpha_{i}^{(d)}\big)^{2}/\beta_{i}^{(d)}$), while attributing importance only to changes within each regime. If a regime corresponds to an inactive configuration (so the value is always $\bot$), then $p_{\gamma^{\prime},\,i}^{(d)}$ and $p_{\gamma,\,i}^{(d)}$ are degenerate at $\bot$ and the corresponding divergence is zero. When the hyperparameter is not conditional, we have $K^{(d)}=1$ and $\alpha_{1}^{(d)}=\beta_{1}^{(d)}=1$, so Equation 5 reduces to the original PED-ANOVA in Equation 1.
(a) Conditional activation (Equation 6) with disjoint domains (Equation 7)
(b) Conditional activation (Equation 6) with overlapping domains (Equation 8)
(c) Regime-dependent domains (Equation 9 and Equation 10)
Figure 1. condPED-ANOVA ($\gamma=1.0$) HPI computed for the synthetic objectives. For the objective with conditional activation (Equation 6), the gating hyperparameter $c$ determines which branch is active: $x$ is present only when $c<0.5$, whereas $y$ is present only when $c\geq 0.5$. For the objective with regime-dependent domains (Equation 9), the gating hyperparameter $c$ determines the domain of $x$ and $y$. The lines denote the mean, and the shaded regions denote the standard error, both computed over ten independent runs with different random seeds. The results on other values of $\gamma$ are provided in Figure 5.
5. Numerical Experiments
Here, we first empirically validate condPED-ANOVA on synthetic problems to verify that it properly reflects conditional structure, including parameter activation and regime-dependent domains. We choose a synthetic setting because the expected behavior is clear and easy to diagnose, unlike real-world objectives where the “correct” importance patterns are often ambiguous. Further experimental results, such as runtime comparison, are available in Appendix C.
5.1. Common Experimental Setup
We use a fixed evaluation budget of N=1, [^0] N=1{,}000 and draw each sample independently from a uniform distribution over the search space. We consider three choices of $\gamma\in\{1.0,\,0.75,\,0.5\}$. For each choice of $\gamma$, we sweep $\gamma^{\prime}$ from $0.01$ up to $\gamma-0.01$ in increments of $0.01$, and run experiments for all resulting $(\gamma^{\prime},\gamma)$ quantile pairs. Our condPED-ANOVA implementation is built on top of the PED-ANOVA implementation in Optuna (Akiba et al., 2019). We repeat the HPI computation with 10 different random seeds and report the mean and the standard error across runs. Further details are provided in Section D.1.
5.2. Conditional Activation (Presence/Absence)
5.2.1. Problem Setting
We first consider a simple synthetic optimization problem with a conditional search space, where the active variable switches depending on a gating parameter. The objective function is defined as:
(6) $f(c,x,y)=\begin{cases}x,&\text{if }c<0.5,\\ y,&\text{if }c\geq 0.5,\end{cases}\quad\text{with }c\in[0,1].$
Our goal is to minimize $f$. Here, only one of $(x,y)$ is active depending on the value of $c$. That is, when $c<0.5$, the parameter $x$ is active and $y$ is inactive; when $c\geq 0.5$, vice versa.
We study the following two domain settings:
(7) (Disjoint domains) $x\in[-5,-2],~y\in[2,5],$
(8) (Overlapping domains) $x\in[-5,2],~y\in[-2,5].$
The first setting creates a clear separation between the two branches, whereas the second introduces overlap between the active-variable ranges, making the conditional structure less trivially identifiable from the returned objective values.
Figure 2. Baseline HPIs computed with naive extensions of existing methods for the synthetic objective with conditional activation (Equation 6) under the disjoint domain setting (Equation 7). “Filtering” computes HPI on the subset of samples where the target hyperparameter is active, whereas “Imputation” assigns a default value (the domain midpoint) to inactive samples before computing HPI. The lines and bars denote the mean, and the shaded regions and error bars denote the standard error, both computed over ten independent runs with different random seeds.
5.2.2. Results and Discussion
The results are shown in Figures 1 and 1. The value at each $\gamma^{\prime}$ can be interpreted as a target-dependent importance: it indicates which hyperparameters matter most if we aim to optimize the objective up to the top-$\gamma^{\prime}$ performance.
We first discuss the disjoint-domain setting (Equation 7) with $\gamma=1.0$ (Figure 1). In this figure, around $\gamma^{\prime}\approx 0.5$, the HPI of the gating variable $c$ is the largest, whereas for a tighter target such as $\gamma^{\prime}\approx 0.1$, the HPI shifts to $x$, and for looser targets such as $\gamma^{\prime}\approx 0.9$, the HPI is dominated by $y$. This behavior is intuitive. When $c<0.5$, the objective equals $x$ and is restricted to a low-valued range (i.e., better performance), while when $c\geq 0.5$, the objective equals $y$ and lies in a high-valued range (i.e., worse performance). Accordingly, whether a sample achieves at least median-level performance is largely determined by which branch is selected, which is reflected by the high importance of $c$ around $\gamma^{\prime}\approx 0.5$. On the other hand, to achieve further better performance (smaller $\gamma^{\prime}$), it is no longer sufficient to select the favorable branch; one must further optimize the active variable itself, which explains the increasing importance of $x$ as $\gamma^{\prime}$ decreases. Moreover, for $\gamma^{\prime}<0.5$, top-$\gamma^{\prime}$ performance requires the $c<0.5$ branch and $y$ to be inactive, which is consistent with the (near) zero importance of $y$. Conversely, for $\gamma^{\prime}>0.5$, whether a sample falls into the top-$\gamma^{\prime}$ region is determined by selecting the $c\geq 0.5$ branch, leading $x$ to be inactive and unimportant.
We next discuss the overlapping-domain setting (Equation 8) with $\gamma=1.0$ (Figure 1). In this setting, the domains of $x$ and $y$ overlap, so the objective values are no longer cleanly separated by the gating variable $c$ alone. Consequently, around $\gamma^{\prime}\approx 0.5$ the HPI of $c$ is no longer close to one, and both $x$ and $y$ also exhibit non-negligible importance. As $\gamma^{\prime}$ decreases, the contribution of $y$ persists up to approximately $\gamma^{\prime}\approx 0.2$, which corresponds to the best values attainable by the $y$-branch (near y=2y=-2). For tighter targets beyond this point, achieving top-$\gamma^{\prime}$ performance requires selecting the $c<0.5$ branch, consistent with the importance of $y$ dropping to (near) zero. Within the $c<0.5$ branch, achieving higher performance increasingly depends on the precise value of $x$, which explains why the HPI of $x$ grows as $\gamma^{\prime}$ becomes smaller.
Also, under both domain settings, the resulting HPI is approximately symmetric for $\gamma^{\prime}<0.5$ and $\gamma^{\prime}>0.5$, reflecting that the objective function is symmetric between the top and bottom regions. However, in the region $\gamma^{\prime}>0.5$, changing $c$ can still improve the objective value; consequently, a small but non-zero importance for $c$ remains even around $\gamma^{\prime}\approx 0.0$, leading to a slight asymmetry. This asymmetry arises from the fact that we minimize the objective, which breaks the top–bottom symmetry.
These results suggest that condPED-ANOVA yields sensible importance values for both the gating variable $c$ and the conditioned variables $x$ and $y$ across target levels $\gamma^{\prime}$. The results for varying $\gamma$ are presented and discussed in Section C.1, where we also observe condPED-ANOVA consistently shows reasonable behavior.
5.3. Regime-Dependent Domains
5.3.1. Problem Setting
We consider a synthetic optimization problem in which the domains of hyperparameters change depending on a gating variable.
The objective is to minimize:
(9) $f(c,x,y)=x+y,$
with the following regime-dependent domains:
(10) $c\in[0,1],~\begin{cases}x\in[-7,-2],~y\in[-5,-2],&\text{if }c<0.5,\\ x\in[2,7],~y\in[2,5],&\text{if }c\geq 0.5.\end{cases}$
Here, both $x$ and $y$ are always present, but their feasible ranges switch according to $c$.
5.3.2. Results and Discussion
The results are shown in Figure 1. Similar to the disjoint-domain setting (Equation 7) in the conditional activation experiment (Section 5.2), the gating variable $c$ determines whether a sample achieves top-half performance, which is reflected by the high importance of $c$ around $\gamma^{\prime}\approx 0.5$. On the other hand, unlike the conditional activation case, both $x$ and $y$ are always active and contribute to the objective, so their importance grows toward more extreme targets (smaller or larger $\gamma^{\prime}$). Moreover, $x$ is consistently more important than $y$, reflecting its ability to attain more extreme values (around $\pm 7$).
5.4. Comparison with Existing Methods
(a) Conditional activation (Equation 6) with disjoint domains (Equation 7)
(b) Conditional activation (Equation 6) with overlapping domains (Equation 8)
(c) Regime-dependent domains (Equation 9 and Equation 10)
Figure 3. Ablation study results obtained using standard local HPI instead of conditional local HPI we defined in Section 4.2. The lines denote the mean, and the shaded regions denote the standard error, both computed over ten independent runs with different random seeds.
5.4.1. Baselines
We compare condPED-ANOVA with existing HPI estimators. Since none of the standard HPI methods are designed to handle conditional hyperparameters in a principled manner, we evaluate them under common naive extensions that enable their application to conditional search spaces.
We consider the following four baselines:
  • PED-ANOVA (Watanabe et al., 2023), the original PED-ANOVA method;
  • f-ANOVA (Hutter et al., 2014), which estimates HPI via functional ANOVA decomposition on a random forest surrogate;
  • MDI (Louppe et al., 2013), the mean decrease impurity feature importance computed from a random forest surrogate; and
  • SHAP (Lundberg and Lee, 2017), which assigns importance based on Shapley values (Shapley, 1953) computed from a surrogate model, measuring each variable’s average marginal contribution.
For hyperparameters whose presence changes depending on other variables, we use the following two naive extensions:
  • Filtering: From the evaluation set $\{(\bm{x}_{n},f(\bm{x}_{n}))\}_{n=1}^{N}$, we keep only samples in which the target hyperparameter is present, and compute HPI on the filtered subset.
  • Imputation: For samples in which the target hyperparameter is inactive, we impute its value by a default value (the midpoint of its domain) and compute HPI on the resulting completed dataset.
For hyperparameters whose domain switches across regimes, we use the following naive extension:
  • Expansion: We treat the hyperparameter as if it were sampled from a single expanded domain that covers all regime-specific ranges, and compute HPI on this unified representation.
The pseudocodes of these naive extensions are provided in Section D.2. Consequently, for conditional activation, we obtain $4\times 2=8$ baseline variants (four HPI methods combined with two naive handling schemes), whereas for regime-dependent domains, we obtain $4\times 1=4$ variants (four HPI methods combined with the expansion scheme), which we compare against condPED-ANOVA. See Section D.2 for further details on the experimental setup.
5.4.2. Results and Discussion
We first analyze the results on the conditional activation objective (Equation 6) under the disjoint-domain setting (Equation 7). As shown in Figure 2, applying simple filtering to the original PED-ANOVA leads to almost identical HPI values for $x$ and $y$. This behavior is clearly undesirable: in order to achieve better performance (i.e., smaller $\gamma^{\prime}$), the hyperparameter $y$ is inactive, and tuning $y$ cannot affect the objective value at all. This failure arises because filtering discards information about how the presence or absence of the variable relates to the overall performance landscape.
When imputation is applied instead, PED-ANOVA assigns a higher HPI to $x$ in the small-$\gamma^{\prime}$ regime, which is qualitatively reasonable. However, the method still assigns a non-negligible importance to $y$ and it becomes comparable to that of $c$, even though $y$ should be inactive in this region. Moreover, for $\gamma^{\prime}>0.5$, the HPI of $x$ and $y$ exhibit substantial instability. These behaviors are caused by forcibly imputing an artificial value (the midpoint of the domain) for inactive configurations, which distorts the empirical sample distribution and introduces spurious dependencies.
Furthermore, in both filtering and imputation extensions of PED-ANOVA, the hyperparameter $c$, which solely determines whether the objective value belongs to the top-performing half, is not properly identified as important. Around $\gamma^{\prime}\approx 0.5$, all hyperparameters are assigned nearly identical HPI values, indicating a complete failure to capture the dominant role of $c$ in controlling performance.
Similar issues are observed for the other baseline estimators. For f-ANOVA and MDI with filtering, all hyperparameters receive identical importance values, while all the remaining variants still assign identical HPI values to $x$ and $y$.
These results demonstrate that naive extensions are insufficient for adapting existing HPI estimators to conditional search spaces and fail to produce meaningful or interpretable importance estimates. In contrast, condPED-ANOVA consistently yields meaningful importance values that reflect the underlying conditional structure (as shown in Section 5.2). Further discussion on the other objectives and domain settings is deferred to Section C.2, where we observe the baselines fail to yield meaningful HPI in conditional settings.
5.5. Ablation Study: Effect of Within-Regime Variance and Regime Weighting
In this section, we conduct ablation studies to examine the impact of key design choices in condPED-ANOVA, focusing on within-regime variance and regime-level weighting.
5.5.1. Effect of Within-Regime Variance for HPI
In Section 4.2, we defined the conditional local HPI using the within-regime local marginal variance $v_{\gamma,\mathrm{within}}^{(d)}$. We therefore compare the resulting HPI estimates with those obtained using the standard local marginal variance $v_{\gamma}^{(d)}$. The closed-form expression for the standard local marginal variance, computed using the indicator function $b_{\gamma^{\prime}}\coloneqq\mathbf{1}\{x\in\mathcal{X}_{\gamma^{\prime}}\}$, is given as follows, where we can also derive a decomposition of the Pearson divergence into within-regime and inter-regime components (cf. Lemma E.2 in Section E.3):
(11) $v_{\gamma}^{(d)}=\left(\frac{\gamma^{\prime}}{\gamma}\right)^{\!2}\Bigg($ $\underbrace{\!\sum_{i=1}^{\;K^{(d)}}\!\!\frac{\big(\alpha_{i}^{(d)}\big)^{2\!}}{\beta_{i}^{(d)}}D_{\mathrm{PE}}\!\left(p_{\gamma^{\prime},\,i}^{(d)}\middle\|p_{\gamma,\,i}^{(d)}\!\right)}_{\text{within-regime divergence}}$
$\hskip 30.00005pt+\underbrace{D_{\mathrm{PE}}\!\left(\!\bm{\alpha}^{(d)}\middle\|\bm{\beta}^{(d)}\!\right)}_{\text{inter-regime divergence}}\Bigg),$
where $\bm{\alpha}^{(d)}\coloneqq(\alpha_{1}^{(d)},\ldots,\alpha_{K^{(d)}}^{(d)})$ and $\bm{\beta}^{(d)}\coloneqq(\beta_{1}^{(d)},\ldots,\beta_{K^{(d)}}^{(d)})$.
(a) Conditional activation (Equation 6) with disjoint domains (Equation 7)
(b) Conditional activation (Equation 6) with overlapping domains (Equation 8)
(c) Regime-dependent domains (Equation 9 and Equation 10)
Figure 4. Ablation study results obtained by using the naive aggregation scheme (Equation 12) instead of our closed-form expression (Theorem 4.2). The bottom row shows the HPIs before normalization, i.e., $\tilde{v}_{\gamma,\mathrm{within}}^{(d)}$. The lines denote the mean, and the shaded regions denote the standard error, both computed over ten independent runs with different random seeds.
The results obtained using the standard local marginal variance are shown in Figure 3. In both the conditional-activation setting with disjoint domains (Figure 3) and the regime-dependent domain setting (Figure 3), the HPI of the gating variable $c$ is underestimated around $\gamma^{\prime}\approx 0.5$ and becomes lower than those of $x$ and $y$. This behavior is undesirable, since selecting the appropriate branch is crucial for achieving median-level performance. Moreover, in the conditional-activation setting (Figures 3 and 3), the importance of $x$ or $y$ does not vanish even in the regions where the variable is inactive and is assigned non-zero HPIs, which are even larger than that of $c$ (e.g., $y$ for $\gamma^{\prime}<0.5$ and $x$ for $\gamma^{\prime}>0.5$ in Figure 3, and $y$ for $\gamma^{\prime}\lesssim 0.2$ and $x$ for $\gamma^{\prime}\gtrsim 0.8$ in Figure 3). This happens because the inter-regime variance is primarily induced by the gating variable $c$. When HPI is computed using the standard local marginal variance that includes this inter-regime component, the effect of $c$ leaks into the HPI estimates for $x$ and $y$. As a result, $c$ is suppressed by $x$ and $y$ in the region where $c$ should be most important, and $x$ and $y$ receive spurious non-zero importance even in the regions where they should be inactive and have zero effect. This phenomenon can also be established mathematically. Theorem E.1 shows that, under the standard local marginal variance, the variance of a conditioned hyperparameter includes the gating variance as an additive term, which explains the observed leakage and the resulting misattribution of importance. Therefore, in conditional settings, it is essential to compute HPI using the within-regime local marginal variance $v_{\gamma,\mathrm{within}}^{(d)}$, rather than the standard local marginal variance $v_{\gamma}^{(d)}$.
5.5.2. Effect of Regime-Level Weighting for HPI
In Section 4.3, we derived a closed-form computation method for $v_{\gamma,\mathrm{within}}^{(d)}$ based on a mathematical derivation (Theorem 4.2). However, one might wonder whether such a derivation is necessary in practice, and instead consider a naive alternative that simply sums the within-regime Pearson divergences, i.e., using:
(12) $\tilde{v}_{\gamma,\mathrm{within}}^{(d)}=\left(\frac{\gamma^{\prime}}{\gamma}\right)^{\!2}\sum_{i=1}^{K^{(d)}}D_{\mathrm{PE}}\!\left(p_{\gamma^{\prime},\,i}^{(d)}\,\middle\|\,p_{\gamma,\,i}^{(d)}\right)$
in place of Equation 5. In this section, we empirically compare condPED-ANOVA against this naive aggregation scheme.
The results obtained using this naive aggregation scheme are shown in Figure 4. In this case, the resulting HPI curves become noticeably less stable and exhibit near-discontinuous changes around regime-switching boundaries. In particular, for the conditional-activation objective under the overlapping-domain setting (Equations 6 and 8), the importance of $y$ spikes immediately when $y$ becomes active, excessively dominating the HPI and suppressing the importances of $x$ and $c$, which should still contribute to the objective. Moreover, although the objective is symmetric between the top region ($\gamma^{\prime}<0.5$) and the bottom region ($\gamma^{\prime}>0.5$) across all settings, the naive scheme breaks this symmetry too severely, yielding clearly distorted importance patterns.
This instability is largely driven by abrupt scale changes at regime boundaries. As shown in the bottom row of Figure 4, $\tilde{v}_{\gamma,\mathrm{within}}^{(d)}$ for $x$ and $y$ exhibit implausible jumps at these boundaries, in sharp contrast to the regime-probability-weighted condPED-ANOVA (cf. Figure 9 in Section C.4). This failure occurs because removing the regime-probability ratios destroys the frequency-aware weighting across regimes, leading to an inappropriate aggregation that cannot properly account for regime prevalence.
6. Conclusion
We studied HPI estimation in conditional search spaces, where hyperparameters may appear, disappear, or change domain depending on other choices. We identified a fundamental pitfall of standard local HPI: conditioning effects can leak into conditioned hyperparameters, leading to misleading importances. To address this, we defined conditional local HPI and proposed condPED-ANOVA, a closed-form, PED-ANOVA-style estimator that correctly handles conditional search spaces while retaining efficiency. Empirically, condPED-ANOVA yielded meaningful importances consistent with the underlying conditional structure, whereas naive extensions of existing estimators, including PED-ANOVA, often violated conditional semantics and produced distorted importances. We believe condPED-ANOVA serves as a practical and principled foundation for reliable HPI analysis in HPO workflows with conditional search spaces.
Appendix A Additional Related Work
A.1. Contrastive Hyperparameter Importance
Beyond global variance-based rankings, several lines of work emphasize selection, contrastive explanations, or local sensitivity. Hutter et al. (2013) study forward selection to identify a small subset of influential inputs (parameters and instance features) by incrementally adding dimensions that most improve surrogate predictive accuracy, followed by a drop-one analysis to rank selected inputs. In parallel, ablation analysis (Fawcett and Hoos, 2016) explains performance differences between two high-quality configurations by constructing a path that sequentially ablates parameter changes and quantifies their contributions, yielding a contrastive, configuration-to-configuration notion of importance. Biedenkapp et al. (2017) accelerate this paradigm via ablation with surrogates, replacing expensive evaluations with model-based predictions to achieve large speedups while maintaining explanatory utility. To further focus on neighborhood behavior, CAVE (Biedenkapp et al., 2019) introduces local parameter importance, which measures variance in predicted performance when varying one parameter around a specific configuration, and empirically argues that local and global importance can disagree and thus complement each other.
A.2. Importance Across Datasets and Tunability
Another line of work studies importance beyond a single dataset or run. Van Rijn and Hutter (2018) aggregates evidence across many datasets (via OpenML metadata) to identify which hyperparameters tend to matter most for an algorithm in general, and to infer priors over promising values; this reframes HPI as a meta-learning problem and supports more data-driven space design and warm-starting. Complementarily, (Probst et al., 2019) introduce tunability measures that quantify how much performance can improve by tuning, providing a practitioner-facing notion of “which hyperparameters are worth tuning” and offering data-based defaults and empirical assessments across datasets and algorithms.
A.3. Explainable HPO via Partial Dependence and Symbolic Models
In parallel, a growing body of work tackles explainability of HPO using tools from interpretable machine learning. Moosbauer et al. (2021) study partial dependence plots (PDPs) for explaining BO-based HPO runs and show that naive PDPs can be biased due to the non-uniform, sequential sampling of BO. They propose uncertainty-aware PDPs based on the BO surrogate, together with a partitioning of the hyperparameter space into subregions where the surrogate is more reliable, to obtain more trustworthy visual summaries of marginal hyperparameter effects. Segel et al. (2023) go one step further and introduce symbolic explanations for HPO: using symbolic regression on meta-data collected from BO runs, they learn compact analytic formulas that approximate the relationship between hyperparameters and performance, thereby providing globally interpretable, human-readable explanations that complement fANOVA-style variance decompositions and PDP-based visualizations.
A.4. Tree-Based Feature Importances
A widely used family of importance measures arises from tree ensembles such as random forests (Breiman, 2001). In particular, the mean decrease impurity (MDI) ranks variables by aggregating their impurity reductions over splits across the ensemble. MDI is popular due to its simplicity and low computational cost, and has been theoretically characterized for randomized trees as a variance decomposition under asymptotic conditions (Louppe et al., 2013). At the same time, impurity-based importances can exhibit biases (e.g., with respect to variable scale or cardinality), motivating careful interpretation and alternative corrections (Strobl et al., 2007).
A.5. Game-Theoretic Explanations for HPI
Game-theoretic attribution methods provide another approach to importance. The Shapley value (Shapley, 1953) offers an axiomatic way to assign each feature its average marginal contribution to a model output. In machine learning, SHAP (Lundberg and Lee, 2017) leverages Shapley values to produce feature attributions for individual predictions, and global importance scores can be obtained by aggregating these local attributions across data. These Shapley-based scores are increasingly used as surrogate-based importance baselines in interpretability.
Within HPO, Rodemann et al. (2026) propose ShapleyBO, explaining Bayesian optimization proposals by attributing each parameter’s contribution to the acquisition function rather than to model predictions, and further decomposing contributions into exploration vs. exploitation terms for additive acquisition functions. More directly for HPI, HyperSHAP (Wever et al., 2026) uses Shapley values and Shapley interaction indices to provide additive decompositions of a performance measure over hyperparameters, aiming to support both local and global explanations, including interactions.
A.6. Multi-Objective HPI
For multi-objective HPO, Theodorakopoulos et al. (2024) propose a framework that computes HPI across objective trade-offs by applying surrogate-based importance measures (notably f-ANOVA and ablation paths) to scalarized objectives, thereby producing importance profiles that vary with the weighting of competing objectives.
Appendix B Rigorous Details of condPED-ANOVA
We briefly presented an overview of condPED-ANOVA in Section 4. Here, we provide the rigorous details of the method.
For simplicity, we omit the boldface notation for vectors $x$. We assume that the objective function $f$ is measurable.
B.1. Extended One-Dimensional Domain
Section 4.1 introduced a regime-based representation of conditional hyperparameters. Here, we provide a rigorous formulation of the corresponding hyperparameter search space. This formalization is necessary to define the empirical distributions $\mu_{\gamma^{\prime}}^{(d)}$ and $\mu_{\gamma}^{(d)}$ (and corresponding PDFs $p_{\gamma^{\prime}}$ and $p_{\gamma}$) over a conditional search space, which underlies the subsequent derivations.
We assume that the $d$-th hyperparameter can be in one of $K^{(d)}\in\mathbb{N}$ regimes. Formally, let $r^{(d)}:\mathcal{X}\to\{1,\ldots,K^{(d)}\}$ be a measurable regime function that specifies the regime of the $d$-th hyperparameter for each configuration $x\in\mathcal{X}$. Each regime $i\in\{1,\ldots,K^{(d)}\}$ is associated with its own domain $\mathcal{Z}_{i}^{(d)}$, which may differ across regimes.
To capture both regime information and regime-specific domain in a single one-dimensional object, we introduce an extended domain for the $d$-th hyperparameter as:
$\mathsf{S}^{(d)}\coloneqq\bigsqcup_{i=1}^{\;\;\;K^{(d)}}\left(\{i\}\times\mathcal{Z}_{i}^{(d)}\right),$
[…]
B.2. Rigorous Formulation of condPED-ANOVA
Let $\mu_{\gamma^{\prime}}^{(d)}$ and $\mu_{\gamma}^{(d)}$ denote the empirical distributions of the extended $d$-th hyperparameter on $\mathsf{S}^{(d)}$ induced by the top-$\gamma^{\prime}$ and top-$\gamma$ subsets, respectively. These correspond to densities $p_{\gamma^{\prime}}$ and $p_{\gamma}$ introduced in Section 4 with respect to a suitable reference measure.
For each regime $i$, recall the measurable component $A_{i}^{(d)}\subset\mathsf{S}^{(d)}$ and define the regime probabilities:
$\alpha_{i}^{(d)}\coloneqq\mu_{\gamma^{\prime}}^{(d)}(A_{i}^{(d)}),\quad\text{and}\quad\beta_{i}^{(d)}\coloneqq\mu_{\gamma}^{(d)}(A_{i}^{(d)}).$
Whenever $\alpha_{i}^{(d)}>0$ and $\beta_{i}^{(d)}>0$, define the conditional (restricted) measures:
$\mu_{\gamma^{\prime},\,i}^{(d)}\coloneqq\mu_{\gamma^{\prime}}^{(d)}(\cdot\mid A_{i}^{(d)}),\quad\text{and}\quad\mu_{\gamma,\,i}^{(d)}\coloneqq\mu_{\gamma}^{(d)}(\cdot\mid A_{i}^{(d)}).$
If $\alpha_{i}^{(d)}=0$, we do not need to define $\mu_{\gamma^{\prime},\,i}^{(d)}$ since its contribution vanishes in Equation 14 through the factor $\big(\alpha_{i}^{(d)}\big)^{2}$. Since $\mathsf{S}^{(d)}=\bigsqcup_{i=1}^{K^{(d)}}A_{i}^{(d)}=\bigsqcup_{i=1}^{K^{(d)}}\big(\{i\}\times\mathcal{Z}_{i}^{(d)}\big)$ forms a disjoint partition, we obtain the following regime-wise decomposition of the marginals:
(13) $\mu_{\gamma^{\prime}}^{(d)}=\sum_{i=1}^{\;K^{(d)}}\alpha_{i}^{(d)}\mu_{\gamma^{\prime},\,i}^{(d)}\,,\quad\text{and}\quad\mu_{\gamma}^{(d)}=\sum_{i=1}^{\;K^{(d)}}\beta_{i}^{(d)}\mu_{\gamma,\,i}^{(d)}.$
Note that we have the absolute continuity $\mu_{\gamma^{\prime}}^{(d)}\ll\mu_{\gamma}^{(d)}$, since $\mathcal{X}_{\gamma^{\prime}}\subset\mathcal{X}_{\gamma}$. The regime-wise $\mu_{\gamma^{\prime},\,i}^{(d)}\ll\mu_{\gamma,\,i}^{(d)}$ also holds for all regimes with $\beta_{i}>0$, since absolute continuity is preserved under restriction to measurable subsets.
With these definitions, we obtain the following closed-form expression for the within-regime local marginal variance:
Theorem B.1 (condPED-ANOVA (Restated)).
Let $0<\gamma^{\prime}<\gamma\leq 1$. The within-regime local marginal variance for the $d$-th hyperparameter at level $\gamma$, computed using the indicator function $b_{\gamma^{\prime}}\coloneqq\mathbf{1}\{x\in\mathcal{X}_{\gamma^{\prime}}\}$, is given by:
(14) $v_{\gamma,\mathrm{within}}^{(d)}=\left(\frac{\gamma^{\prime}}{\gamma}\right)^{\!2}\sum_{i:\,\beta_{i}^{(d)}>0}\frac{\big(\alpha_{i}^{(d)}\big)^{2}}{\beta_{i}^{(d)}}\,D_{\mathrm{PE}}\!\left(\mu_{\gamma^{\prime},i}^{(d)}\middle\|\mu_{\gamma,i}^{(d)}\right),$
Here, the Pearson ($\chi^{2}$) divergence is defined as:
$D_{\mathrm{PE}}(\nu\|\mu)\coloneqq\int\left(\frac{d\nu}{d\mu}-1\right)^{\!2}\,d\mu,$
for probability measures $\nu$ and $\mu$ with $\nu\ll\mu$. Note that the summation in Equation 14 is restricted to regimes with $\beta_{i}^{(d)}>0$; regimes absent from the top-$\gamma$ set contribute neither to the within-regime variance nor to the divergence. If $A_{i}^{(d)}$ corresponds to an inactive configuration (i.e., $\mathcal{Z}_{i}^{(d)}=\{\bot\}$), then $\mu_{\gamma,\,i}^{(d)}$ and $\mu_{\gamma^{\prime},\,i}^{(d)}$ are both degenerate on $A_{i}^{(d)}$, implying $D_{\mathrm{PE}}(\mu_{\gamma^{\prime},\,i}^{(d)}\|\mu_{\gamma,\,i}^{(d)})=0$. When $K^{(d)}=1$, we have $\alpha_{1}=\beta_{1}=1$ and Equation 14 reduces to the original PED-ANOVA expression.
In practice, the sample size within each regime $A_{i}^{(d)}$ must be large enough for KDE to be stable. When the domains $\mathcal{Z}_{i}^{(d)}$ vary continuously with other coordinates, it is often preferable to discretize them into fixed intervals so that sufficient samples can be pooled within each regime.
B.3. Proof of Theorem B.1
Proof.
Let $S$ be a random element distributed according to the empirical distribution $\mu_{\gamma}^{(d)}$ on $\mathsf{S}^{(d)}$ induced by the top-$\gamma$ subset. Let $Y\in\{0,1\}$ indicate whether the underlying sample comes from the tighter top set, i.e., $Y=1$ corresponds to membership in the top-$\gamma^{\prime}$ subset. By definition, the conditional law of $S$ given $Y=1$ is $\mu_{\gamma^{\prime}}^{(d)}$. Let $\kappa\coloneqq\mathbb{P}(Y=1)$. By construction, $\kappa=|\mathcal{D}_{\gamma^{\prime}}|/|\mathcal{D}_{\gamma}|$; under the quantile definition used in the main text, this equals $\gamma^{\prime}/\gamma$ up to the floor effect, and we use $\kappa=\gamma^{\prime}/\gamma$ for simplicity.
For any measurable set $B\subset\mathsf{S}^{(d)}$, we have:
$\mathbb{P}(S\in B,\,Y=1)$ $=\mathbb{P}(Y=1)\,\mathbb{P}(S\in B\mid Y=1)$
$=\kappa\,\mu_{\gamma^{\prime}}^{(d)}(B),$
and
$\mathbb{P}(S\in B)=\mu_{\gamma}^{(d)}(B).$
Since $\mathcal{X}_{\gamma^{\prime}}\subset\mathcal{X}_{\gamma}$, we have $\mu_{\gamma^{\prime}}^{(d)}\ll\mu_{\gamma}^{(d)}$, and thus the Radon–Nikodým derivative $d\mu_{\gamma^{\prime}}^{(d)}/d\mu_{\gamma}^{(d)}$ exists. Thus, by the definition of conditional probability, for $s\in\mathsf{S}^{(d)}$, we have:
(15) $\mathbb{P}\!\left(Y=1\mid S=s\right)\;=\;\kappa\,\frac{\mu_{\gamma^{\prime}}^{(d)}}{\mu_{\gamma}^{(d)}}(s).$
Since $\mathsf{S}^{(d)}=\bigsqcup_{i=1}^{K^{(d)}}A_{i}^{(d)}=\bigsqcup_{i=1}^{K^{(d)}}\big(\{i\}\times\mathcal{Z}_{i}^{(d)}\big)$ forms a disjoint partition (Equation 13), the following factorization holds for each $i$ with $\beta_{i}^{(d)}>0$, for $s\in A_{i}^{(d)}$:
(16) $\frac{d\mu_{\gamma^{\prime}}^{(d)}}{d\mu_{\gamma}^{(d)}}(s)=\frac{d(\alpha_{i}^{(d)}\,\mu_{\gamma^{\prime},\,i}^{(d)})}{d(\beta_{i}^{(d)}\,\mu_{\gamma,\,i}^{(d)})}(s)=\frac{\alpha_{i}^{(d)}}{\beta_{i}^{(d)}}\,\frac{d\mu_{\gamma^{\prime},\,i}^{(d)}}{d\mu_{\gamma,\,i}^{(d)}}(s),$
Using Equation 15 and Equation 16, for each $i$ with $\beta_{i}^{(d)}>0$, for $s\in A_{i}^{(d)}$, we obtain:
$\mathbb{P}(Y=1\mid S=s)\;=\;\kappa\,\frac{\alpha_{i}^{(d)}}{\beta_{i}^{(d)}}\,\frac{d\mu_{\gamma^{\prime},\,i}^{(d)}}{d\mu_{\gamma,\,i}^{(d)}}(s)$
Recall that the within-regime local HPI is defined as the within-regime variance of the conditional probability of being in the tighter top set, i.e.,
$v_{\gamma,\mathrm{within}}^{(d)}\coloneqq\mathbb{E}_{I}\!\left[\mathrm{Var}\!\left(\mathbb{P}(Y=1\mid S)\,\middle|\,I\right)\right].$
Taking expectation over $I$ and using $\mathbb{P}(I=i)=\beta_{i}$ yields:
$v_{\gamma,\mathrm{within}}^{(d)}$ $=\sum_{i:\,\beta_{i}^{(d)}>0}\beta_{i}^{(d)}\mathrm{Var}_{s\sim\mu_{\gamma,\,i}^{(d)}}\!\left(\kappa\frac{\alpha_{i}^{(d)}}{\beta_{i}^{(d)}}\frac{d\mu_{\gamma^{\prime},\,i}^{(d)}}{d\mu_{\gamma,\,i}^{(d)}}(s)\right)$
$=\kappa^{2}\sum_{i:\,\beta_{i}>0}\frac{\big(\alpha_{i}^{(d)}\big)^{2}}{\beta_{i}^{(d)}}\mathrm{Var}_{s\sim\mu_{\gamma,\,i}^{(d)}}\!\left(\frac{d\mu_{\gamma^{\prime},\,i}^{(d)}}{d\mu_{\gamma,\,i}^{(d)}}(s)\right).$
$=\kappa^{2}\sum_{i:\,\beta_{i}^{(d)}>0}\frac{\big(\alpha_{i}^{(d)}\big)^{2}}{\beta_{i}^{(d)}}\int\left(\frac{d\mu_{\gamma^{\prime},\,i}^{(d)}}{d\mu_{\gamma,\,i}^{(d)}}(s)-1\right)^{2}d\mu_{\gamma,\,i}^{(d)}(s)$
$=\kappa^{2}\sum_{i:\,\beta_{i}^{(d)}>0}\frac{\big(\alpha_{i}^{(d)}\big)^{2}}{\beta_{i}^{(d)}}D_{\mathrm{PE}}\!\left(\mu_{\gamma^{\prime},\,i}^{(d)}\middle\|\mu_{\gamma,\,i}^{(d)}\right).$
Here, we used the following equation, which holds since $\mu_{\gamma^{\prime},i}^{(d)}$ is a probability measure:
$\mathbb{E}_{s\sim\mu_{\gamma,i}^{(d)}}\left[\frac{d\mu_{\gamma^{\prime},\,i}^{(d)}}{d\mu_{\gamma,\,i}^{(d)}}(s)\right]=\int\frac{d\mu_{\gamma^{\prime},\,i}^{(d)}}{d\mu_{\gamma,\,i}^{(d)}}\,d\mu_{\gamma,\,i}^{(d)}=\mu_{\gamma^{\prime},\,i}^{(d)}(\mathsf{S}^{(d)})=1.$
This completes the proof. ∎
Appendix C Additional Results and Discussion on Numerical Experiments
(a) Conditional activation (Equation 6) with disjoint domains (Equation 7)
(b) Conditional activation (Equation 6) with overlapping domains (Equation 8)
(c) Regime-dependent domains (Equation 9 and Equation 10)
Figure 5. condPED-ANOVA HPI computed for the synthetic objectives. For the objective with conditional activation (Equation 6), the gating hyperparameter $c$ determines which branch is active: $x$ is present only when $c<0.5$, whereas $y$ is present only when $c\geq 0.5$. For the objective with regime-dependent domains (Equation 9), the gating hyperparameter $c$ determines the domain of the hyperparameters $x$ and $y$. The lines denote the mean, and the shaded regions denote the standard error, both computed over ten independent runs with different random seeds.
C.1. condPED-ANOVA Results on Synthetic Objectives with Different $\gamma$ Values
The results of condPED-ANOVA for the objective with conditional activation (Equation 6) and the objective with regime-dependent domains (Equation 9), with $\gamma$ set to $0.75$ and $0.5$, are shown in Figure 5.
The results share some similarities with the higher-$\gamma^{\prime}$ range in Figure 1, which reflects the fact that varying $\gamma$ changes the region over which HPI is computed. However, the behavior is not a simple zoom-in because the definition of the top-$\gamma$ region itself changes. In particular, with $\gamma=0.5$ in the disjoint-domain setting (Figure 5, right panel), once we restrict attention to the better half of the evaluations (i.e., the region where $c<0.5$ is already selected), adjusting $c$ no longer changes the objective value. Accordingly, the importance of $c$ becomes small in this regime, similar to the importance of $y$.
C.2. Comparison with Existing Methods
Figure 6. Baseline HPIs computed with naive extensions of existing methods for the synthetic objective with conditional activation (Equation 6) under the overlapping domain setting (Equation 8). The lines and bars denote the mean, and the shaded regions and error bars denote the standard error, both computed over ten independent runs with different random seeds.
Figure 7. Baseline HPIs computed with naive extensions of existing methods for the synthetic objective with regime-dependent domains (Equation 9 and Equation 10). The lines and bars denote the mean, and the shaded regions and error bars denote the standard error, both computed over ten independent runs with different random seeds.
The results for the baseline methods described in Section 5.4 for the conditional-activation objective with the overlapping-domain setting (Equations 6 and 8) and the regime-dependent domain setting (Equations 9 and 10) are shown in Figures 6 and 7, respectively.
First, for the conditional-activation objective, we observe the same trend as in the disjoint-domain setting (Equations 6 and 7). PED-ANOVA with filtering fails to distinguish between $x$ and $y$, while PED-ANOVA with imputation assigns $y$ an importance comparable to $c$ in the small-$\gamma^{\prime}$ regime, where $y$ should be inactive and thus cannot improve performance. Similarly, f-ANOVA with filtering and MDI with filtering again yield nearly identical HPI values for all hyperparameters, and the remaining baselines also tend to assign almost the same importance to $x$ and $y$.
Next, for the regime-dependent domain setting, PED-ANOVA with expansion also behaves undesirably: although $c$ determines whether a configuration falls into the top half, the importance of $c$ around $\gamma^{\prime}\approx 0.5$ becomes smaller than that of $x$ and $y$. This failure is caused by expansion, which collapses the regime-specific ranges of $x$ and $y$ into a single unified domain and thereby obscures the effect of range switching, preventing accurate estimation of $x$ and $y$ in the regimes where their domains change. Moreover, for f-ANOVA, MDI, and SHAP with expansion, $y$ is assigned larger importance than $x$, even though $x$ should contribute more to the objective. This is because forcibly expanding the domain distorts the relationship between the variables and the objective across regimes, leading to misleading HPI estimates.
These results indicate that the naive extensions do not provide meaningful HPI estimates that capture the difference between the contributions of $x$ and $y$ in this conditional setting, not only for the conditional-activation objective under the disjoint-domain setting (as shown in Section 5.4.2), but also under the overlapping-domain setting and in the regime-dependent domain setting.
C.3. Runtime Comparison
Figure 8. Runtime comparison against baseline HPI methods with naive extensions for the synthetic objective (Equation 9 and Equation 10). The lines denote the mean, and the shaded regions denote the standard error, both computed over ten independent runs with different random seeds.
We compare the runtime of condPED-ANOVA against baseline HPI methods equipped with naive extensions on the synthetic objective with regime-dependent domains (Equation 9 and Equation 10). For each sample size from $2^{9}$ to $2^{17}$, we measure the wall-clock time required to compute HPI, and report the results in Figure 8. f-ANOVA is reported only up to $2^{11}$, since it could not be executed at larger sample sizes due to excessive memory consumption. Due to its substantial memory consumption, f-ANOVA could not be executed for larger sample sizes and is reported only up to $2^{11}$. Similarly, SHAP is reported only up to $2^{15}$ because it was prohibitively slow at larger sample sizes.
In these results, condPED-ANOVA runs faster than all baseline methods across the evaluated ranges, indicating that the overhead introduced by our approach is small. Here, the fact that condPED-ANOVA is faster than PED-ANOVA indicates that the overhead of filtering or imputation is larger than the additional computation introduced by condPED-ANOVA.
(a) Conditional activation (Equation 6) with disjoint domains (Equation 7)
(b) Conditional activation (Equation 6) with overlapping domains (Equation 8)
(c) Regime-dependent domains (Equation 9 and Equation 10)
Figure 9. condPED-ANOVA ($\gamma=1.0$) HPI values before normalization, i.e., $v_{\gamma,\mathrm{within}}^{(d)}$, computed for the synthetic objectives. The lines denote the mean, and the shaded regions denote the standard error, both computed over ten independent runs with different random seeds.
C.4. Ablation Study on the Effect of Regime-Level Weighting
The ablation study results obtained by using a naive aggregation scheme (Equation 12) instead of the closed-form expression derived in Theorem 4.2 are shown in Figure 4. For reference, we also show the HPI values before normalization, i.e., $\tilde{v}_{\gamma,\mathrm{within}}^{(d)}$, in Figure 9.
Appendix D Detailed Experimental Setup
D.1. Details for Common Experimental Setup
For estimating the PDFs $p_{\gamma^{\prime},\,i}^{(d)}$ and $p_{\gamma,\,i}^{(d)}$, we use KDE with Optuna’s default settings, which employ Scott’s rule for bandwidth selection (Scott, 1992). All experiments are run on a single machine with an Intel® CoreTM Ultra 7 155H CPU (22 logical CPUs) and 16 GB RAM, running Arch Linux. We use Python version 3.13.11 and Optuna v4.7.0.
D.2. Implementation Details for Baselines
For all four baseline methods, we use the implementations provided in Optuna (Akiba et al., 2019) and follow Optuna’s default hyperparameter settings. The pseudocode for naive extension for these baselines is provided in Algorithms 2, 3 and 4
Algorithm 2 Naive Extension: Filtering (Conditional Presence)
1:Evaluation set $\mathcal{D}=\{(\bm{x}_{n},f(\bm{x}_{n}))\}_{n=1}^{N}$, baseline evaluator $\textsc{BaselineHPI}(\cdot)$
2:Naive HPI for the hyperparameters $\bm{x}$ via filtering
3:for $d=1,\ldots,D$ do
4:  $\mathcal{D}_{\mathrm{act}}^{(d)}\leftarrow\{(\bm{x}_{n},f(\bm{x}_{n}))\in\mathcal{D}\mid x_{n}^{(d)}\text{ is present}\}$ $\triangleright$ filtering
5:  if $|\mathcal{D}_{\mathrm{act}}^{(d)}|=0$ then
6:   $\text{HPI}^{(d)}\leftarrow 0$ $\triangleright$ no active trials; define HPI as zero
7:  else
8:   $\text{HPI}\leftarrow d\text{-th value of }\textsc{BaselineHPI}(\mathcal{D}_{\mathrm{act}}^{(d)})$
9:  end if
10:end for
11:return $\{\text{HPI}^{(d)}\}_{d=1}^{D}$
Algorithm 3 Naive Extension: Imputation (Conditional Presence)
1:Evaluation set $\mathcal{D}=\{(\bm{x}_{n},f(\bm{x}_{n}))\}_{n=1}^{N}$, hyperparameter domains $\mathcal{X}^{(1)}\times\cdots\times\mathcal{X}^{(D)}=[\ell^{(1)},u^{(1)}]\times\cdots\times[\ell^{(D)},u^{(D)}]$, baseline evaluator $\textsc{BaselineHPI}(\cdot)$
2:Naive HPI for the hyperparameters via imputation
3:Initialize $\tilde{\mathcal{D}}\leftarrow\emptyset$
4:for $n=1,\ldots,N$ do
5:  $\tilde{\bm{x}}_{n}\leftarrow\bm{x}_{n}$
6:  for $d=1,\ldots,D$ do
7:   if $x_{n}^{(d)}\text{ is inactive}$ in $\bm{x}_{n}$ then
8:     $\triangleright$ imputation with the midpoint of the domain
9:     Set $\tilde{x}_{n}^{(d)}\leftarrow(\ell^{(d)}+u^{(d)})/2$
10:   end if
11:  end for
12:  $\tilde{\mathcal{D}}\leftarrow\tilde{\mathcal{D}}\cup\{(\tilde{\bm{x}}_{n},f(\bm{x}_{n}))\}$
13:end for
[…]
Algorithm 4 Naive Extension: Expansion (Regime-Specific Ranges)
1:Evaluation set $\mathcal{D}=\{(\bm{x}_{n},f(\bm{x}_{n}))\}_{n=1}^{N}$, regime-specific domains for each $d$-th hyperparameter $\{\mathcal{X}^{(d)}_{i}=[\ell^{(d)}_{i},u^{(d)}_{i}]\}_{i=1}^{K^{(d)}}$, baseline evaluator $\textsc{BaselineHPI}(\cdot)$
2:Naive HPI for the hyperparameters via domain expansion
3:for $d=1,\ldots,D$ do
4:  $\ell^{(d)}_{\mathrm{min}}\leftarrow\min_{i\in\{1,\ldots,K^{(d)}\}}\ell^{(d)}_{i}$
5:  $u^{(d)}_{\mathrm{max}}\leftarrow\max_{i\in\{1,\ldots,K^{(d)}\}}u^{(d)}_{i}$
6:  $\mathcal{\tilde{X}}^{(d)}=[\ell^{(d)}_{\mathrm{min}},u^{(d)}_{\mathrm{max}}$] $\triangleright$ expanded domain
7:end for
8:Initialize $\tilde{\mathcal{D}}\leftarrow\emptyset$
9:for $n=1,\ldots,N$ do
10:  $\tilde{\bm{x}}_{n}\in\mathcal{\tilde{X}}^{(d)}\leftarrow\bm{x}_{n}$ $\triangleright$ treat each $x_{n}^{(d)}$ as a sample from $\mathcal{\tilde{X}}^{(d)}$
11:  $\tilde{\mathcal{D}}\leftarrow\tilde{\mathcal{D}}\cup\{(\tilde{\bm{x}}_{n},f(\bm{x}_{n}))\}$
12:end for
13:return $\textsc{BaselineHPI}(\tilde{\mathcal{D}},d)$
Appendix E Supplementary Analysis for the Ablation Study Using Standard Local HPI
E.1. Leakage of Gating Effects in Standard Local HPI
Let $b_{\gamma^{\prime}}\coloneqq\mathbf{1}\{\bm{x}\in\mathcal{X}_{\gamma^{\prime}}\}$ be the top-set indicator. Consider a gating hyperparameter $x^{(c)}$ and a conditional (child) hyperparameter $x^{(d)}$ whose regime is determined by $x^{(c)}$. Formally, assume that for a conditional hyperparameter $x^{(d)}$ there exists a measurable mapping $\phi^{(d)}\colon\mathcal{X}^{(c)}\to\{1,\ldots,K^{(d)}\}$ such that:
(17) $I^{(d)}=\phi^{(d)}(x^{(c)})\ \ \text{a.s.\hbox{}},$
where $I^{(d)}\in\{1,\ldots,K^{(d)}\}$ denotes the random variable representing the regime index of the $d$-th hyperparameter. Moreover, assume that $x^{(c)}$ affects $b_{\gamma^{\prime}}$ only through the induced regime (i.e., $x^{(c)}$ has no within-regime effect):
(18) $\mathbb{E}_{\gamma}\left[b_{\gamma^{\prime}}\mid x^{(c)}\right]=\mathbb{E}_{\gamma}\left[b_{\gamma^{\prime}}\mid I^{(d)}\right]\quad\text{a.s.\hbox{}},$
where the expectations are taken under the empirical distribution restricted to the top-$\gamma$ region $\mathcal{X}_{\gamma}$.
Here, $x^{(c)}$ corresponds to the gating variable $c$ in the synthetic objectives (Equations 6 and 9), while $x^{(d)}$ corresponds to the conditional variables $x$ and $y$. These assumptions are satisfied by the synthetic objectives used in our experiments.
Then, we have the following result regarding the standard local HPI of the conditional hyperparameter $x^{(d)}$.
Theorem E.1 (Leakage of gating effects under the standard local HPI).
Under the above assumptions (Equations 17 and 18), the standard local marginal variance of the conditional hyperparameter $x^{(d)}$ contains the local marginal variance of the gating hyperparameter $x^{(c)}$:
(19) $v_{\gamma}^{(d)}=v_{\gamma,\mathrm{within}}^{(d)}+v_{\gamma}^{(c)}.$
In particular, $v_{\gamma}^{(c)}\leq v_{\gamma}^{(d)}$.
The proof is provided in Section E.2. Consequently, when HPI is computed from $v_{\gamma}^{(d)}$, the effect of the gating variable $x^{(c)}$ is additively inherited by conditional hyperparameters through the inter-regime term.
This theorem is consistent with our empirical observations. In Figure 3, when using the standard local HPI, the importance of the gating variable $c$ is always no larger than those of $x$ and $y$. Moreover, in the regions of $\gamma^{\prime}$ where $x$ or $y$ is inactive, their HPI becomes almost identical to that of $c$. This is exactly what Theorem E.1 predicts: once $v_{\gamma,\mathrm{within}}^{(d)}$ vanishes for an inactive variable, the standard local variance $v_{\gamma}^{(d)}$ reduces to the additive gating term $v_{\gamma}^{(c)}$.
E.2. Proof of Theorem E.1
Proof.
In this section, all expectations and variances are taken under the empirical distribution restricted to the top-$\gamma$ region $\mathcal{X}_{\gamma}$, and we omit the subscript $\gamma$ on $\mathbb{E}$ and $\mathrm{Var}$ for notational simplicity.
Recall we defined $g_{\gamma}^{(d)}(I^{(d)},Z^{(d)})\coloneqq\mathbb{E}\!\left[b_{\gamma^{\prime}}\mid I^{(d)},Z^{(d)}\right]$ in Section 4.2. We also restate the variance decomposition in Equation 2 here for convenience:
(20) $v_{\gamma}^{(d)}$ $=\mathrm{Var}_{I^{(d)},\,Z^{(d)}}\left(g_{\gamma}^{(d)}(I^{(d)},Z^{(d)})\right)$
$=\underbrace{\mathbb{E}_{I^{(d)}}\left[\mathrm{Var}_{Z^{(d)}}\left(g_{\gamma}^{(d)}(I^{(d)},Z^{(d)})\,\middle|\,I^{(d)}\right)\right]}_{\text{within-regime variance: }v_{\gamma,\mathrm{within}}^{(d)}}$
$\hskip 30.00005pt+\underbrace{\mathrm{Var}_{I^{(d)}}\left(\mathbb{E}_{Z^{(d)}}\left[g_{\gamma}^{(d)}(I^{(d)},Z^{(d)})\,\middle|\,I^{(d)}\right]\right)}_{\text{inter-regime variance}}.$
By the tower property, we have:
$\mathbb{E}_{Z^{(d)}}\left[g_{\gamma}^{(d)}(I^{(d)},Z^{(d)})\,\middle|\,I^{(d)}\right]$ $=\mathbb{E}_{Z^{(d)}}\Big[\mathbb{E}\big[b_{\gamma^{\prime}}\,\big|\,I^{(d)},Z^{(d)}\big]\,\Big|\,I^{(d)}\Big]$
$=\mathbb{E}\!\left[b_{\gamma^{\prime}}\mid I^{(d)}\right].$
Thus, the inter-regime variance in Equation 20 equals:
(21) $v_{\gamma,\mathrm{inter}}^{(d)}$ $\coloneqq\mathrm{Var}_{I^{(d)}}\left(\mathbb{E}_{Z^{(d)}}\left[g_{\gamma}^{(d)}(I^{(d)},Z^{(d)})\,\middle|\,I^{(d)}\right]\right)$
$=\mathrm{Var}_{I^{(d)}}\!\left(\mathbb{E}\left[b_{\gamma^{\prime}}\mid I^{(d)}\right]\right).$
On the other hand, the local marginal mean of the top-set indicator $b_{\gamma^{\prime}}$ with respect to the gating coordinate $x^{(c)}$ is:
$f_{\gamma}^{(c)}(x^{(c)})$ $\coloneqq\frac{1}{Z_{\gamma}^{(c)}(x^{(c)})}\sum_{n=1}^{N}b_{\gamma^{\prime}}(x^{(c)},x^{(\bar{c}\,)}_{n})\,b_{\gamma}(x^{(c)},x^{(\bar{c}\,)}_{n})$,
$=\mathbb{E}\left[b_{\gamma^{\prime}}\mid x^{(c)}\right].$
Thus, by Equation 18, we have:
(22) $v_{\gamma}^{(c)}$ $\coloneqq\mathrm{Var}_{x^{(c)}}\left(f_{\gamma}^{(c)}(x^{(c)})\right)$
$=\mathrm{Var}_{x^{(c)}}\left(\mathbb{E}\left[b_{\gamma^{\prime}}\mid x^{(c)}\right]\right)$
$=\mathrm{Var}_{x^{(c)}}\left(\mathbb{E}\left[b_{\gamma^{\prime}}\mid I^{(d)}\right]\right).$
From Equation 17, the quantity $\mathbb{E}[b_{\gamma^{\prime}}\mid I^{(d)}]$ is a function of $x^{(c)}$ through $I^{(d)}$. Therefore,
(23) $\mathrm{Var}_{x^{(c)}}\!\left(\mathbb{E}\!\left[b_{\gamma^{\prime}}\mid I^{(d)}\right]\right)=\mathrm{Var}_{I^{(d)}}\!\left(\mathbb{E}\!\left[b_{\gamma^{\prime}}\mid I^{(d)}\right]\right).$
Comparing Equations 21, 22 and 23, we conclude that the inter-regime variance of the conditional hyperparameter $x^{(d)}$ equals the marginal variance of the gating hyperparameter $x^{(c)}$:
$v_{\gamma,\mathrm{inter}}^{(d)}=v_{\gamma}^{(c)}.$
Substituting this into Equation 20 completes the proof. ∎
E.3. Decomposition of PED over Regimes
Here, we use the notation defined in Sections B.1 and B.2. That is, on the extended domain $\mathsf{S}^{(d)}$, $\mu_{\gamma}^{(d)}$ and $\mu_{\gamma^{\prime}}^{(d)}$ denote the empirical distributions of the extended $d$-th hyperparameter induced by the samples in $\mathcal{X}_{\gamma}$ and $\mathcal{X}_{\gamma^{\prime}}$, respectively.
We can then show the following decomposition of the Pearson divergence between $\mu_{\gamma}^{(d)}$ and $\mu_{\gamma^{\prime}}^{(d)}$ on $\mathsf{S}^{(d)}$.
Lemma E.2 (Regime-wise decomposition of Pearson divergence).
Assume $\mu_{\gamma^{\prime}}^{(d)}\ll\mu_{\gamma}^{(d)}$ and $\mu_{\gamma^{\prime},\,i}^{(d)}\ll\mu_{\gamma,\,i}^{(d)}$ for every $i$ with $\beta_{i}^{(d)}>0$. Then the Pearson divergence between the extended marginals satisfies:
(24) $D_{\mathrm{PE}}\left(\mu_{\gamma^{\prime}}^{(d)}\,\middle\|\,\mu_{\gamma}^{(d)}\right)$ $=\underbrace{\sum_{i=1}^{\;K^{(d)}}\frac{\big(\alpha_{i}^{(d)}\big)^{2}}{\beta_{i}^{(d)}}D_{\mathrm{PE}}\left(\mu_{\gamma^{\prime},\,i}^{(d)}\,\middle\|\,\mu_{\gamma,\,i}^{(d)}\right)}_{\text{within-regime divergence}}$
$\hskip 50.00008pt+\underbrace{D_{\mathrm{PE}}\left(\alpha^{(d)}\middle\|\,\beta^{(d)}\right)}_{\text{inter-regime divergence}}\hskip-12.0pt,\hskip-2.0pt$
where $\alpha^{(d)}\coloneqq\big(\alpha_{i}^{(d)}\big)_{i=1}^{K^{(d)}}$ and $\beta^{(d)}\coloneqq\big(\beta_{i}^{(d)}\big)_{i=1}^{K^{(d)}}$ are the discrete distributions over regimes.
The proof is provided in Section E.4. The absolute continuity $\mu_{\gamma^{\prime}}^{(d)}\ll\mu_{\gamma}^{(d)}$ holds for the extended marginals in the same way as in Equation 1. The regime-wise $\mu_{\gamma^{\prime},\,i}^{(d)}\ll\mu_{\gamma,\,i}^{(d)}$ also holds for all regimes with $\beta_{i}^{(d)}>0$, since absolute continuity is preserved under restriction to measurable subsets. If a regime $i$ corresponds to an inactive configuration (i.e., $\mathcal{Z}_{i}^{(d)}=\{\bot\}$), then $\mu_{\gamma^{\prime},\,i}^{(d)}$ and $\mu_{\gamma,\,i}^{(d)}$ are both degenerate at $\bot$, so $D_{\mathrm{PE}}\left(\mu_{\gamma^{\prime},\,i}^{(d)}\,\middle\|\,\mu_{\gamma,\,i}^{(d)}\right)=0$ and only the inter-regime term contributes.
Using this decomposition, we obtain Equation 11.
E.4. Proof of Lemma E.2
Proof.
Here, we omit the superscript (d){(d)} for notational simplicity. Since $\mu_{\gamma^{\prime}}\ll\mu_{\gamma}$ and $\mu_{\gamma^{\prime},\,i}\ll\mu_{\gamma,\,i}$ holds for $\beta_{i}>0$, the Radon–Nikodým derivative $d\mu_{\gamma^{\prime}}/d\mu_{\gamma}$ exists and is given by:
$\frac{d\mu_{\gamma^{\prime}}}{d\mu_{\gamma}}(x)=\frac{d(\alpha_{i}\,\mu_{\gamma^{\prime},\,i})}{d(\beta_{i}\,\mu_{\gamma,\,i})}(x)=\frac{\alpha_{i}}{\beta_{i}}\,\frac{d\mu_{\gamma^{\prime},\,i}}{d\mu_{\gamma,\,i}}(x),$
for $x\in A_{i}$. We denote $s_{i}\coloneqq d\mu_{\gamma^{\prime},\,i}/d\mu_{\gamma,\,i}$.
The Pearson divergence can be expressed as follows:
(25) $D_{\mathrm{PE}}(\mu_{\gamma^{\prime}}\|\mu_{\gamma})$ $=\int\left(\frac{d\mu_{\gamma^{\prime}}}{d\mu_{\gamma}}-1\right)^{2}d\mu_{\gamma}$
$=\sum_{i=1}^{\;K^{(d)}}\int_{A_{i}^{(d)}}\!\Bigl(\frac{\alpha_{i}}{\beta_{i}}s_{i}-1\Bigr)^{2}d(\beta_{i}\,\mu_{\gamma,\,i})$
$=\sum_{i=1}^{\;K^{(d)}}\beta_{i}\int\!\Bigl(\frac{\alpha_{i}}{\beta_{i}}s_{i}-1\Bigr)^{2}d\mu_{\gamma,\,i}$
$=\sum_{i=1}^{\;K^{(d)}}\left(\frac{\alpha_{i}^{2}}{\beta_{i}}\int s_{i}^{2}d\mu_{\gamma,\,i}-2\alpha_{i}\int s_{i}d\mu_{\gamma,\,i}+\beta_{i}\right).$
Here, we have:
(26) $\int s_{i}d\mu_{\gamma,\,i}=\int\frac{d\mu_{\gamma^{\prime},\,i}}{d\mu_{\gamma,\,i}}d\mu_{\gamma,\,i}=\mu_{\gamma^{\prime},\,i}(A_{i})=1.$
Thus, we obtain:
(27) $\int s_{i}^{2}d\mu_{\gamma,\,i}$ $=\int s_{i}^{2}d\mu_{\gamma,\,i}-2\int s_{i}d\mu_{\gamma,\,i}+1+1$
$=\int(s_{i}-1)^{2}d\mu_{\gamma,\,i}+1$
$=D_{\mathrm{PE}}\left(\mu_{\gamma^{\prime},\,i}\,\middle\|\,\mu_{\gamma,\,i}\right)+1.$
By substituting Equations 26 and 27 into Equation 25, we have:
$D_{\mathrm{PE}}(\mu_{\gamma^{\prime}}\|\mu_{\gamma})$ $=\sum_{i=1}^{\;K^{(d)}}\left(\frac{\alpha_{i}^{2}}{\beta_{i}}\left(D_{\mathrm{PE}}\left(\mu_{\gamma^{\prime},\,i}\,\middle\|\,\mu_{\gamma,\,i}\right)+1\right)-2\alpha_{i}+\beta_{i}\right)$
$=\sum_{i=1}^{\;K^{(d)}}\left(\frac{\alpha_{i}^{2}}{\beta_{i}}D_{\mathrm{PE}}\left(\mu_{\gamma^{\prime},\,i}\,\middle\|\,\mu_{\gamma,\,i}\right)+\frac{(\alpha_{i}-\beta_{i})^{2}}{\beta_{i}}\right)$
$=\sum_{i=1}^{\;K^{(d)}}\frac{\alpha_{i}^{2}}{\beta_{i}}D_{\mathrm{PE}}\left(\mu_{\gamma^{\prime},\,i}\,\middle\|\,\mu_{\gamma,\,i}\right)+\sum_{i=1}^{\;K^{(d)}}\frac{(\alpha_{i}-\beta_{i})^{2}}{\beta_{i}}$
$=\sum_{i=1}^{\;K^{(d)}}\frac{\alpha_{i}^{2}}{\beta_{i}}D_{\mathrm{PE}}\left(\mu_{\gamma^{\prime},\,i}\,\middle\|\,\mu_{\gamma,\,i}\right)+D_{\mathrm{PE}}\left(\alpha\,\middle\|\,\beta\right).$
This completes the proof.

### v2

@@
 To address this, we introduce a conditional HPI for top-performing regions and derive a closed-form estimator that accurately reflects conditional activation and domain changes.
-Experiments show that naive adaptations of existing HPI estimators yield misleading or uninterpretable importance estimates in conditional settings, whereas condPED-ANOVA consistently provides meaningful importances that reflect the underlying conditional structure.
+Experiments show that naive adaptations of existing HPI estimators yield misleading or uninterpretable importances in conditional settings, whereas condPED-ANOVA consistently provides meaningful importances that reflect the underlying conditional structure.
+Our code is publicly available at https://github.com/kAIto47802/condPED-ANOVA.
 Analysis of Variance, Hyperparameter Importance, Conditional Search Space, Automated Machine Learning
-copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Machine learning approachesccs: Mathematics of computing Mathematical analysis
+copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: The 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 09–13, 2026; Jeju, Koreaisbn: 978-1-4503-XXXX-X/2026/02ccs: Computing methodologies Machine learning approachesccs: Mathematics of computing Mathematical analysis
 1. Introduction
 Analysis of parameter importance (Sobol, 1993, 2001; Hooker, 2007; Huang, 1998) is a central tool for understanding and designing optimization procedures.
-Over the past decade, techniques for measuring the importance of parameters have attracted considerable attention in the automated machine learning (AutoML) community (Hutter et al., 2014; Biedenkapp et al., 2017; Van Rijn and Hutter, 2018), particularly for improving the efficiency of hyperparameter optimization (HPO) (Bergstra et al., 2011; Bergstra and Bengio, 2012; Snoek et al., 2012; Lindauer et al., 2022).
-Given a hyperparameter search space and a set of evaluated configurations, methods for hyperparameter importance (or HPI methods for short) aim to quantify how strongly each hyperparameter influences the observed performance.
-Such information is useful for interpreting optimization runs, identifying critical design choices, detecting irrelevant or redundant parameters, and guiding subsequent search or manual tuning (Melis et al., 2018; Henderson et al., 2018).
+Over the past decade, techniques for measuring parameter importance have attracted considerable attention in the automated machine learning (AutoML) community (Hutter et al., 2014; Biedenkapp et al., 2017; van Rijn and Hutter, 2018), particularly for improving hyperparameter optimization (HPO) (Bergstra et al., 2011; Bergstra and Bengio, 2012; Snoek et al., 2012; Lindauer et al., 2022; Eggensperger et al., 2013).
+Given a hyperparameter search space and evaluated configurations, methods for hyperparameter importance (HPI) aim to quantify how strongly each hyperparameter influences the observed performance.
+Such information helps interpret optimization runs, identify critical design choices, detect irrelevant or redundant parameters, and guide subsequent search or manual tuning (Melis et al., 2018; Henderson et al., 2018).
 As a result, HPI is widely used across various settings, including HPO of machine learning models and HPO algorithm design, where understanding hyperparameter sensitivity is often essential (van Rijn and Hutter, 2018; Zimmer et al., 2021; Probst et al., 2019; Theodorakopoulos et al., 2024).
-Classic HPI methods such as functional analysis of variance (f-ANOVA) (Hutter et al., 2014) quantify HPI by measuring how much variance in the objective can be attributed to each hyperparameter.
+Classic HPI methods such as functional analysis of variance (f-ANOVA) (Hutter et al., 2014) quantify HPI by measuring how much objective variance is attributed to each hyperparameter.
 However, their existing local variants mix contributions from unpromising regions, leading to inaccurate estimates of HPI in top-performing subspaces (Hutter et al., 2014; Biedenkapp et al., 2019).
-PED-ANOVA (Watanabe et al., 2023) mitigates this by defining local HPI within a top-performing subset of the search space and efficiently approximating it using the Pearson divergence between one-dimensional distributions of each hyperparameter.
-Moreover, PED-ANOVA computes both global and local HPIs substantially faster than f-ANOVA while maintaining accurate HPI estimates.
+PED-ANOVA (Watanabe et al., 2023) addresses this by defining local HPI within a top-performing subset and efficiently approximating it using the Pearson divergence between one-dimensional distributions of each hyperparameter.
+Moreover, PED-ANOVA computes both global and local HPIs substantially faster than f-ANOVA while preserving accuracy.
 Thanks to this combination of speed and reliability, it has been adopted for HPI analysis in practical HPO frameworks, such as Optuna (Akiba et al., 2019), demonstrating its utility in real-world workflows.
-However, we argue that PED-ANOVA has a practical limitation: it cannot properly handle hierarchical/dynamic search spaces, where the presence/domain of a hyperparameter depends on other hyperparameters (hereafter referred to as conditional search spaces).
-In practice, many real-world problems involve such conditional search spaces (Hutter et al., 2019).
-Typical examples include model selection, where hyperparameters are dependent on a chosen algorithm (Thornton et al., 2013; Kotthoff et al., 2017; Feurer et al., 2015; Komer et al., 2014), and neural network architecture design, where active parameters depend on the chosen layers (Bergstra et al., 2013; Swersky et al., 2014).
-Unfortunately, PED-ANOVA cannot be directly applied to such settings, since it assumes a fixed, unconditional search space and defines distributions over a single, fixed domain for each hyperparameter.
-We observe that naive extensions of PED-ANOVA, such as filtering inactive configurations, imputing missing values, or artificially expanding parameter domains, do not resolve this issue: these heuristics often violate the conditional semantics of the search space and yield unnatural or uninterpretable HPI estimates (cf.
+However, we argue that PED-ANOVA has a practical limitation: it cannot properly handle hierarchical/dynamic search spaces (hereafter conditional search spaces), where the presence or domain of a hyperparameter depends on others.
+In practice, many real-world problems involve conditional search spaces (Hutter et al., 2019).
+Typical examples include model selection, where hyperparameters depend on the chosen algorithm (Thornton et al., 2013; Kotthoff et al., 2017; Feurer et al., 2015; Komer et al., 2014), and neural network architecture design, where active parameters depend on the chosen layers (Bergstra et al., 2013; Swersky et al., 2014).
+Unfortunately, PED-ANOVA cannot be directly applied to such settings, as it assumes a fixed, unconditional search space and defines distributions over a single, fixed domain for each hyperparameter.
+We observe that naive extensions of PED-ANOVA, such as filtering inactive samples, imputing missing values, or artificially expanding parameter domains, do not resolve this issue: these heuristics often violate the conditional semantics of the search space and yield unnatural or uninterpretable HPI estimates (cf.
 Section 5.4).
 To address this limitation, we propose conditional PED-ANOVA (condPED-ANOVA), a principled framework for estimating HPI in conditional search spaces.
-We identify a critical flaw in the original definition of local HPI in PED-ANOVA when applied to conditional search spaces: the importance of hyperparameters that induce conditionality is spuriously attributed to the hyperparameters whose domains or activation depend on them, leading to systematically misleading importance estimates.
+We identify a critical flaw in PED-ANOVA’s original definition of local HPI under conditionality: the importance of hyperparameters that induce conditionality is spuriously attributed to dependent hyperparameters, yielding systematically misleading HPI estimates (empirically shown in Section 5.5.1 and theoretically in Theorem F.1).
 Consequently, merely constructing a method that correctly computes local HPI in conditional search spaces is still insufficient.
-We introduce conditional local HPI, which measures importance within top-performing regions without confounding from conditional activation or domain changes.
-Building on this definition, we derive a closed-form PED-ANOVA-style estimator that remains efficient and yields an interpretable HPI in conditional settings.
+We thus introduce conditional local HPI, which measures importance within top-performing regions without confounding from conditional activation or domain changes.
+Building on this definition, we derive a closed-form PED-ANOVA-style estimator that yields interpretable HPIs in conditional settings.
 Experiments confirm that condPED-ANOVA reliably produces a meaningful HPI that reflects the underlying conditional structure, while naive extensions of existing methods fail to do so.
@@
   • We identify a fundamental pitfall of original local HPI in conditional search spaces: the contribution of conditioning hyperparameters is spuriously attributed to conditionally active (or domain-shifted) hyperparameters, yielding misleading importance estimates.
-  • We introduce conditional local HPI, a definition that measures importance within top-performing regions while accounting for conditional activation and domain changes.
-  • We derive a closed-form PED-ANOVA-style estimator for conditional local HPI, retaining the efficiency of the original PED-ANOVA via Pearson divergence of one-dimensional distributions estimated with KDE.
+  • We introduce conditional local HPI, a definition that measures importance without being confounded by conditional activation and domain changes.
+  • We derive a closed-form PED-ANOVA-style estimator, namely condPED-ANOVA, for conditional local HPI, retaining PED-ANOVA’s efficiency.
   • We empirically demonstrate that condPED-ANOVA consistently produces meaningful importances in conditional settings, whereas naive extensions of existing methods, including PED-ANOVA, produce unnatural results.
@@
 Early work on parameter importance analysis builds on variance-based sensitivity analysis such as Sobol indices (Sobol, 1993, 2001) and ANOVA decompositions (Hooker, 2007; Huang, 1998), which attribute the global variance of an objective to individual coordinates and their interactions.
-In the context of HPO, f-ANOVA (Hutter et al., 2014) applies this idea to HPO objectives by fitting a surrogate model such as a random forest over the hyperparameter space and then estimating main and interaction effects by Monte Carlo integration over the surrogate.
-This yields importance values that reflect how much of the global variation in performance across the entire search space can be explained by each hyperparameter.
-f-ANOVA includes efficient marginal computation over random-forest predictions and produces a global ranking of important hyperparameters, making it a standard post-hoc analysis tool in AutoML and algorithm configuration.
+In the context of HPO, f-ANOVA (Hutter et al., 2014) applies this idea to HPO objectives by fitting a surrogate model (e.g., a random forest) over the search space and estimating HPIs by integrating the surrogate over tree partitions using region-volume weights.
+This produces global importance scores that quantify how much of the performance variance is explained by each hyperparameter, and it has become a standard post-hoc analysis tool in AutoML.
 2.2. PED-ANOVA
@@
 It defines local HPI as the variance contribution of each hyperparameter within the top-$\gamma$ subset of the search space, thereby measuring importance directly in the most relevant region.
-It further shows that this local HPI can be approximated in closed form using the Pearson divergence between the one-dimensional distributions of each hyperparameter of two top-performing subsets, estimated via one-dimensional KDEs.
+It further shows that this local HPI can be approximated in closed form using the Pearson divergence between the one-dimensional distributions of each hyperparameter of the two top-performing subsets, estimated via one-dimensional kernel density estimators (KDEs) (Parzen, 1962).
 Empirically, this leads to much faster HPI computation than f-ANOVA while maintaining accurate importance estimates in top-performing subspaces.
@@
 2.3. Conditional Hyperparameters in HPO
-Many HPO algorithms furthermore operate on conditional search spaces, where certain hyperparameters are active only under specific parent choices, and their domains may also change depending on those choices (e.g., algorithm selection or internal mode switches).
-This view is made explicit in practical AutoML systems, such as Auto-WEKA (Thornton et al., 2013), auto-sklearn (Feurer et al., 2015, 2022), which describe conditional hyperparameters that are only relevant when a parent hyperparameter takes a particular value.
-Among widely used optimizers, SMAC (Lindauer et al., 2022) is designed to handle such conditional hierarchies in CASH-style spaces (Thornton et al., 2013), while TPE (Bergstra et al., 2011) formulates the search space as a tree-structured generative process where some variables are only well-defined under specific upstream choices.
-BOHB (Falkner et al., 2018) combines Hyperband (Li et al., 2018) with a Bayesian-optimization component built on TPE, thereby optimizing over the same tree-structured conditional spaces while exploiting cheap low-fidelity evaluations.
-Beyond these algorithms, a complementary line of work develops BO surrogates tailored to conditional domains to better share information across branches (Swersky et al., 2014; Jenatton et al., 2017; Ma and Blaschko, 2020).
+Many HPO algorithms operate on conditional search spaces, where some hyperparameters’ activation and domains depend on specific parent choices (e.g., algorithm selection).
+This conditional structure is explicitly modeled in practical AutoML systems, such as Auto-WEKA (Thornton et al., 2013) and auto-sklearn (Feurer et al., 2015, 2022).
+Among widely used optimizers, SMAC (Lindauer et al., 2022) is designed to handle such conditional hierarchies in CASH-style spaces (Thornton et al., 2013), while TPE (Bergstra et al., 2011) formulates the search space as a tree-structured generative process.
+BOHB (Falkner et al., 2018) combines Hyperband (Li et al., 2018) with a TPE-based Bayesian optimization (BO), optimizing over the same tree-structured conditional spaces while exploiting cheap low-fidelity evaluations.
+Beyond these algorithms, a complementary work develops BO surrogates tailored to conditional domains to share information across branches (Swersky et al., 2014; Jenatton et al., 2017; Ma and Blaschko, 2020).
 2.4. HPI under Conditional Search Spaces
-In conditional search spaces, applying standard HPI pipelines generally requires mapping configurations to a fixed representation, typically by imputing inactive hyperparameters with default values as done in SMAC-style toolchains and related preprocessing workflows (Lindauer et al., 2022).
-Practical analysis suites such as CAVE (Biedenkapp et al., 2019) (and its integration in BOAH (Lindauer et al., 2019)) then compute importance post-hoc using methods like f-ANOVA and local parameter importance, including for multi-fidelity runs.
-Optuna’s importance API highlights a common limitation of conditional spaces: by default, it evaluates only parameters that appear and have the same domain in all samples, excluding conditional parameters (Akiba et al., 2019).
+In conditional search spaces, applying standard HPI pipelines requires mapping configurations to a fixed representation, often by imputing inactive hyperparameters with default values as done in SMAC-style toolchains (Lindauer et al., 2022).
+Practical analysis suites such as CAVE (Biedenkapp et al., 2019) (and its integration in BOAH (Lindauer et al., 2019)) and DeepCAVE (Segel et al., 2025) then compute importance post-hoc using methods like f-ANOVA and local parameter importance, including for multi-fidelity runs.
+Current Optuna’s importance API highlights a common limitation of conditional spaces: by default, it evaluates only parameters that appear and have the same domain in all samples, excluding conditional parameters (Akiba et al., 2019).
 Recent work on multi-objective HPI likewise treats missing values arising from conditional hyperparameters in preprocessing before applying surrogate-based importance measures such as f-ANOVA and ablation paths (Theodorakopoulos et al., 2024).
@@
 The local mean $m_{\gamma}$ of the objective and the local marginal mean $f_{\gamma}^{(d)}(x^{(d)})$ for the $d$-th hyperparameter $x^{(d)}$ at level $\gamma$ are defined as:
-$m_{\gamma}$ $\coloneqq\mathbb{E}_{\bm{x}\sim p_{0}}[f(\bm{x})\mid b_{\gamma}(\bm{x})]=\frac{\mathbb{E}_{\bm{x}\sim p_{0}}[f(\bm{x})\,b_{\gamma}(\bm{x})]}{\mathbb{E}_{\bm{x}\sim p_{0}}[b_{\gamma}(\bm{x})]},$
-$f_{\gamma}^{(d)}(x^{(d)})$ $\coloneqq\mathbb{E}_{\bm{x}^{(\bar{d}\,)}\sim p_{0}^{(\bar{d}\,)}}[f(\bm{x})\mid b_{\gamma}(\bm{x}),x^{(d)}]$
-$\,=\frac{\mathbb{E}_{\bm{x}^{(\bar{d}\,)}\sim p_{0}^{(\bar{d}\,)}}[f(\bm{x})\,b_{\gamma}(\bm{x})\mid x^{(d)}]}{\mathbb{E}_{\bm{x}^{(\bar{d}\,)}\sim p_{0}^{(\bar{d}\,)}}[b_{\gamma}(\bm{x})\mid x^{(d)}]},$
+(1) $\hskip 27.11546ptm_{\gamma}\coloneqq\mathbb{E}_{\bm{x}\sim p_{0}}[f(\bm{x})\mid b_{\gamma}(\bm{x})]=\frac{\mathbb{E}_{\bm{x}\sim p_{0}}[f(\bm{x})\,b_{\gamma}(\bm{x})]}{\mathbb{E}_{\bm{x}\sim p_{0}}[b_{\gamma}(\bm{x})]},$
+(2) $\begin{aligned} f_{\gamma}^{(d)}(x^{(d)})&\coloneqq\mathbb{E}_{\bm{x}^{(\bar{d}\,)}\sim p_{0}^{(\bar{d}\,)}}[f(\bm{x})\mid b_{\gamma}(\bm{x}),x^{(d)}]\\ &\,=\frac{\mathbb{E}_{\bm{x}^{(\bar{d}\,)}\sim p_{0}^{(\bar{d}\,)}}[f(\bm{x})\,b_{\gamma}(\bm{x})\mid x^{(d)}]}{\mathbb{E}_{\bm{x}^{(\bar{d}\,)}\sim p_{0}^{(\bar{d}\,)}}[b_{\gamma}(\bm{x})\mid x^{(d)}]},\end{aligned}$
 where $\bm{x}^{(\bar{d}\,)}\!\coloneqq\!(x^{(1)},\ldots,x^{(d-1)},x^{(d+1)},\ldots,x^{(D)})$ denotes all hyperparameters except for the $d$-th one.
@@
 The local marginal variance for the $d$-th hyperparameter at level $\gamma$ is then defined as follows:
-$v_{\gamma}^{(d)}$ $\coloneqq\mathrm{Var}_{\bm{x}\sim p_{0}}\bigl(f_{\gamma}^{(d)}(x^{(d)})\mid b_{\gamma}(\bm{x})\bigr)$
+(3) $v_{\gamma}^{(d)}$ $\coloneqq\mathrm{Var}_{\bm{x}\sim p_{0}}\bigl(f_{\gamma}^{(d)}(x^{(d)})\mid b_{\gamma}(\bm{x})\bigr)$
 $\,=\frac{\mathbb{E}_{\bm{x}\sim p_{0}}[(f_{\gamma}^{(d)}(x^{(d)})-m_{\gamma})^{2}\,b_{\gamma}(\bm{x})]}{\mathbb{E}_{\bm{x}\sim p_{0}}[b_{\gamma}(\bm{x})]}.$
 Using this quantity, the local HPI for the $d$-th hyperparameter at level $\gamma$ is then defined as its normalized form:
-$\text{Local HPI}\coloneqq v_{\gamma}^{(d)}\Big/\sum_{d^{\prime}=1}^{D}v_{\gamma}^{(d^{\prime})}.$
+(4) $\text{Local HPI}\coloneqq v_{\gamma}^{(d)}\Big/\sum_{d^{\prime}=1}^{D}v_{\gamma}^{(d^{\prime})}.$
 3.3. PED-ANOVA
@@
 Here, $D_{\mathrm{PE}}(p\|q)$ denotes the Pearson ($\chi^{2}$) divergence between two PDFs $p$ and $q$, defined as:
-$D_{\mathrm{PE}}(p\|q)\coloneqq\int\left(\frac{p(x)}{q(x)}-1\right)^{2}q(x)\mathrm{d}x.$
+(6) $D_{\mathrm{PE}}(p\|q)\coloneqq\int\left(\frac{p(x)}{q(x)}-1\right)^{2}q(x)\mathrm{d}x.$
 Note that this divergence is well-defined in the present setting, since $\mathcal{X}_{\gamma^{\prime}}\subset\mathcal{X}_{\gamma}$ implies that the density $p_{\gamma^{\prime}}^{(d)}$ has support only where $p_{\gamma}^{(d)}$ is nonzero, so the ratio $p_{\gamma^{\prime}}^{(d)}/p_{\gamma}^{(d)}$ is well-defined.
@@
 Here, the choice of $x^{(0)}$ induces regimes $i^{(d)}\in\{\texttt{neural},\texttt{tree}\}$ under which $x^{(1)}$ and $x^{(2)}$ have regime-specific domains or become inactive.
-Algorithm 1 condPED-ANOVA Computation for the $d$-th Hyperparameter
+Algorithm 1 condPED-ANOVA Computation for the $d$-th Hyperparameter $x^{(d)}$
 1:Evaluation set $\mathcal{D}=\{(\bm{x}_{n},f(\bm{x}_{n}))\}_{n=1}^{N}$, hyperparameter index $d$, quantile levels $0<\gamma^{\prime}\leq\gamma\leq 1$
@@
 8:  $L^{\text{within}}_{i}\leftarrow D_{\mathrm{PE}}\left(p_{\gamma^{\prime},\,i}^{(d)}\,\middle\|\,p_{\gamma,\,i}^{(d)}\right)\triangleright$ within-regime Pearson divergence
-9:  $\alpha_{i}^{(d)}\leftarrow|\mathcal{D}_{\gamma^{\prime},\,i}|\,/\,|\mathcal{D}_{\gamma^{\prime}}|,\;\beta_{i}^{(d)}\leftarrow|\mathcal{D}_{\gamma,\,i}|\,/\,|\mathcal{D}_{\gamma}|$ $\triangleright$ regime probabilities in top $\gamma^{\prime}$ and $\gamma$ subsets
+9:  $\alpha_{i}^{(d)}\leftarrow|\mathcal{D}_{\gamma^{\prime},\,i}^{(d)}|\,/\,|\mathcal{D}_{\gamma^{\prime}}|,\;\beta_{i}^{(d)}\leftarrow|\mathcal{D}_{\gamma,\,i}^{(d)}|\,/\,|\mathcal{D}_{\gamma}|$ $\triangleright$ regime probabilities in top $\gamma^{\prime}$ and $\gamma$ subsets
 10:end for
@@
 We observe that the standard local HPI based on the local marginal variance $v_{\gamma}^{(d)}$ mixes two distinct sources of variation: differences within each regime and differences between (inter) regimes.
-This can be seen through the following variance decomposition:
+This can be seen through the following variance decomposition by the law of total variance:
 (7) $v_{\gamma}^{(d)}$ $=\mathrm{Var}_{I^{(d)},\,Z^{(d)}}\left(g_{\gamma}^{(d)}(I^{(d)},Z^{(d)})\right)$
 $=\underbrace{\mathbb{E}_{I^{(d)}}\left[\mathrm{Var}_{Z^{(d)}}\left(g_{\gamma}^{(d)}(I^{(d)},Z^{(d)})\,\middle|\,I^{(d)}\right)\right]}_{\text{within-regime variance}}$
-$\hskip 25.00003pt+\underbrace{\mathrm{Var}_{I^{(d)}}\left(\mathbb{E}_{Z^{(d)}}\left[g_{\gamma}^{(d)}(I^{(d)},Z^{(d)})\,\middle|\,I^{(d)}\right]\right)}_{\text{inter-regime variance}},$
+$\hskip 26.00009pt+\underbrace{\mathrm{Var}_{I^{(d)}}\left(\mathbb{E}_{Z^{(d)}}\left[g_{\gamma}^{(d)}(I^{(d)},Z^{(d)})\,\middle|\,I^{(d)}\right]\right)}_{\text{inter-regime variance}},$
 where the expectations and variances are taken under the empirical distribution restricted to the top-$\gamma$ region $\mathcal{X}_{\gamma}$.
@@
 However, for conditional hyperparameters, the regime $I^{(d)}$ is determined by other hyperparameters (e.g., parent choices) rather than by the $d$-th hyperparameter itself.
-As a result, attributing this inter-regime effect to the $d$-th hyperparameter is often inappropriate: it can reflect changes in upstream decisions rather than the effect of the value of $x^{(d)}$ itself.
-In fact, we can show that the standard local marginal variance of a conditioned hyperparameter includes the conditioning variable’s local marginal variance, and is therefore always no smaller than it (cf.
-Theorem E.1 in Section E.1).
-In particular, this leakage implies that even inactive hyperparameters can be assigned non-negligible importance solely due to upstream choices.
+Thus, attributing this inter-regime effect to the $d$-th hyperparameter is often inappropriate.
+In fact, we can show that the standard local marginal variance of a conditioned hyperparameter is lower-bounded by the conditioning variable’s local marginal variance, implying systematic leakage (cf.
+Theorem F.1 in Section F.1).
+Consequently, even inactive hyperparameters can receive non-negligible importance solely due to upstream choices.
 Such behavior is undesirable because tuning an inactive hyperparameter has no effect on the objective and is therefore meaningless.
-Motivated by this consideration, we define the local HPI for conditional hyperparameters using only the within-regime component:
+Motivated by this consideration, we define the local HPI for conditional hyperparameters using only the within-regime component.
 Definition 4.1 (Conditional Local HPI).
@@
 4.3. condPED-ANOVA
-We now derive a closed-form condPED-ANOVA estimator that efficiently computes the within-regime local marginal variance of each hyperparameter from one-dimensional PDFs within each regime.
+We now derive a closed-form condPED-ANOVA estimator that efficiently computes the within-regime local marginal variance of each hyperparameter.
+We adopt a PED-ANOVA-style construction since it enables fast HPI computation while maintaining accurate estimates (cf.
+runtime comparisons in Section D.1).
 The conditional local HPI then follows immediately by normalization, as defined in Equation 8.
@@
 We define the regime probabilities:
-$\alpha_{i}^{(d)}\coloneqq\frac{|\mathcal{D}_{\gamma^{\prime},\,i}^{(d)}|}{|\mathcal{D}_{\gamma^{\prime}}|},\quad\text{and}\quad\beta_{i}^{(d)}\coloneqq\frac{|\mathcal{D}_{\gamma,\,i}^{(d)}|}{|\mathcal{D}_{\gamma}|}.$
+(10) $\alpha_{i}^{(d)}\coloneqq\frac{|\mathcal{D}_{\gamma^{\prime},\,i}^{(d)}|}{|\mathcal{D}_{\gamma^{\prime}}|},\quad\text{and}\quad\beta_{i}^{(d)}\coloneqq\frac{|\mathcal{D}_{\gamma,\,i}^{(d)}|}{|\mathcal{D}_{\gamma}|}.$
 For each regime $i^{(d)}$, let $p_{\gamma^{\prime},\,i}^{(d)}$ and $p_{\gamma,\,i}^{(d)}$ denote the one-dimensional PDFs of the $d$-th hyperparameter.
 In practice, these PDFs are estimated using KDE from the samples in $\mathcal{D}_{\gamma^{\prime},\,i}^{(d)}$ and $\mathcal{D}_{\gamma,\,i}^{(d)}$, respectively.
-With these definitions, we obtain the following closed-form expression for the within-regime local marginal variance:
+With these definitions, we obtain the following closed-form estimation for the within-regime local marginal variance.
 Theorem 4.2 (condPED-ANOVA).
@@
 By normalizing the variance across all hyperparameters as defined in Equation 8, we obtain the conditional local HPI for the $d$-th hyperparameter.
-The proof of Theorem 4.2 is provided in Appendix B.
-Equation 5 aggregates within-regime distributional changes using regime-dependent weights $\big(\alpha_{i}^{(d)}\big)^{2}/\beta_{i}^{(d)}$.
+The proof is provided in Appendix B.
+Equation 11 aggregates within-regime Pearson divergence using regime-dependent weights $\big(\alpha_{i}^{(d)}\big)^{2}/\beta_{i}^{(d)}$.
 Intuitively, this weighting emphasizes regimes that are enriched in the tighter top set (i.e., with larger $\big(\alpha_{i}^{(d)}\big)^{2}/\beta_{i}^{(d)}$), while attributing importance only to changes within each regime.
@@
 When the hyperparameter is not conditional, we have $K^{(d)}=1$ and $\alpha_{1}^{(d)}=\beta_{1}^{(d)}=1$, so Equation 11 reduces to the original PED-ANOVA in Equation 5.
+5. Experiments on Synthetic Problems
 (a) Conditional activation (Equation 12) with disjoint domains (Equation 13)
@@
 The results on other values of $\gamma$ are provided in Figure 7.
-5. Numerical Experiments
 Here, we first empirically validate condPED-ANOVA on synthetic problems to verify that it properly reflects conditional structure, including parameter activation and regime-dependent domains.
 We choose a synthetic setting because the expected behavior is clear and easy to diagnose, unlike real-world objectives where the “correct” importance patterns are often ambiguous.
-Further experimental results, such as runtime comparison, are available in Appendix C.
-5.1. Common Experimental Setup
+After that, we confirm that condPED-ANOVA also properly works on real-world benchmarks, which is presented in Section 6.
+Further experimental results, such as runtime comparison and sensitivity analyses with respect to $N$, are available in Appendix D.
+5.1. Experimental Setup
 We use a fixed evaluation budget of N=1, [^0] N=1{,}000 and draw each sample independently from a uniform distribution over the search space.
@@
 We repeat the HPI computation with 10 different random seeds and report the mean and the standard error across runs.
-Further details are provided in Section D.1.
-5.2. Conditional Activation (Presence/Absence)
-5.2.1. Problem Setting
-We first consider a simple synthetic optimization problem with a conditional search space, where the active variable switches depending on a gating parameter. The objective function is defined as:
-(6) $f(c,x,y)=\begin{cases}x,&\text{if }c<0.5,\\ y,&\text{if }c\geq 0.5,\end{cases}\quad\text{with }c\in[0,1].$
-Our goal is to minimize $f$. Here, only one of $(x,y)$ is active depending on the value of $c$. That is, when $c<0.5$, the parameter $x$ is active and $y$ is inactive; when $c\geq 0.5$, vice versa.
-We study the following two domain settings:
-(7) (Disjoint domains) $x\in[-5,-2],~y\in[2,5],$
-(8) (Overlapping domains) $x\in[-5,2],~y\in[-2,5].$
-The first setting creates a clear separation between the two branches, whereas the second introduces overlap between the active-variable ranges, making the conditional structure less trivially identifiable from the returned objective values.
+To ensure reproducibility, we publicly release our condPED-ANOVA implementation and all experimental codes [^1] https://github.com/kAIto47802/condPED-ANOVA.
+Further details are provided in Section C.1.
 Figure 2.
@@
 The lines and bars denote the mean, and the shaded regions and error bars denote the standard error, both computed over ten independent runs with different random seeds.
+5.2. Conditional Activation (Presence/Absence)
+5.2.1. Problem Setting
+We first consider a simple synthetic problem with a conditional search space, where the active variable is selected by a gating parameter. The objective function is defined as:
+(12) $f(c,x,y)=\begin{cases}x,&\text{if }c<0.5,\\ y,&\text{if }c\geq 0.5,\end{cases}\quad\text{with }c\in[0,1].$
+Our goal is to minimize $f$. When $c<0.5$, the parameter $x$ is active and $y$ is inactive; when $c\geq 0.5$, vice versa.
+We study the following two domain settings:
+(13) (Disjoint domains) $x\in[-5,-2],~y\in[2,5],$
+(14) (Overlapping domains) $x\in[-5,2],~y\in[-2,5].$
+The first setting cleanly separates the two branches, whereas the second introduces overlap, making the conditional structure less trivially identifiable from the objective values.
+Experiments with multi-level conditional hierarchies are presented in Section E.1.
 5.2.2. Results and Discussion
@@
 We first discuss the disjoint-domain setting (Equation 13) with $\gamma=1.0$ (Figure 1).
-In this figure, around $\gamma^{\prime}\approx 0.5$, the HPI of the gating variable $c$ is the largest, whereas for a tighter target such as $\gamma^{\prime}\approx 0.1$, the HPI shifts to $x$, and for looser targets such as $\gamma^{\prime}\approx 0.9$, the HPI is dominated by $y$.
+Around $\gamma^{\prime}\approx 0.5$, the gating variable $c$ has the largest HPI, whereas for a tighter target (e.g., $\gamma^{\prime}\approx 0.1$), the HPI shifts to $x$, and for looser targets (e.g., $\gamma^{\prime}\approx 0.9$), the HPI is dominated by $y$.
 This behavior is intuitive.
-When $c<0.5$, the objective equals $x$ and is restricted to a low-valued range (i.e., better performance), while when $c\geq 0.5$, the objective equals $y$ and lies in a high-valued range (i.e., worse performance).
-Accordingly, whether a sample achieves at least median-level performance is largely determined by which branch is selected, which is reflected by the high importance of $c$ around $\gamma^{\prime}\approx 0.5$.
-On the other hand, to achieve further better performance (smaller $\gamma^{\prime}$), it is no longer sufficient to select the favorable branch; one must further optimize the active variable itself, which explains the increasing importance of $x$ as $\gamma^{\prime}$ decreases.
+When $c<0.5$, the objective equals $x$ and lies in a low (better) range; when $c\geq 0.5$, it equals $y$ and lies in a high (worse) range.
+Thus, achieving roughly median performance is mainly determined by the selected branch, explaining the high importance of $c$ near $\gamma^{\prime}\approx 0.5$.
+For smaller $\gamma^{\prime}$, selecting the favorable branch is not enough; one must optimize the active variable, which explains the increasing importance of $x$.
 Moreover, for $\gamma^{\prime}<0.5$, top-$\gamma^{\prime}$ performance requires the $c<0.5$ branch and $y$ to be inactive, which is consistent with the (near) zero importance of $y$.
-Conversely, for $\gamma^{\prime}>0.5$, whether a sample falls into the top-$\gamma^{\prime}$ region is determined by selecting the $c\geq 0.5$ branch, leading $x$ to be inactive and unimportant.
+Conversely, for $\gamma^{\prime}>0.5$, the $c\geq 0.5$ branch dominates, making $y$ important and $x$ inactive (hence unimportant).
 We next discuss the overlapping-domain setting (Equation 14) with $\gamma=1.0$ (Figure 1).
-In this setting, the domains of $x$ and $y$ overlap, so the objective values are no longer cleanly separated by the gating variable $c$ alone.
-Consequently, around $\gamma^{\prime}\approx 0.5$ the HPI of $c$ is no longer close to one, and both $x$ and $y$ also exhibit non-negligible importance.
-As $\gamma^{\prime}$ decreases, the contribution of $y$ persists up to approximately $\gamma^{\prime}\approx 0.2$, which corresponds to the best values attainable by the $y$-branch (near y=2y=-2).
-For tighter targets beyond this point, achieving top-$\gamma^{\prime}$ performance requires selecting the $c<0.5$ branch, consistent with the importance of $y$ dropping to (near) zero.
-Within the $c<0.5$ branch, achieving higher performance increasingly depends on the precise value of $x$, which explains why the HPI of $x$ grows as $\gamma^{\prime}$ becomes smaller.
-Also, under both domain settings, the resulting HPI is approximately symmetric for $\gamma^{\prime}<0.5$ and $\gamma^{\prime}>0.5$, reflecting that the objective function is symmetric between the top and bottom regions.
-However, in the region $\gamma^{\prime}>0.5$, changing $c$ can still improve the objective value; consequently, a small but non-zero importance for $c$ remains even around $\gamma^{\prime}\approx 0.0$, leading to a slight asymmetry.
-This asymmetry arises from the fact that we minimize the objective, which breaks the top–bottom symmetry.
-These results suggest that condPED-ANOVA yields sensible importance values for both the gating variable $c$ and the conditioned variables $x$ and $y$ across target levels $\gamma^{\prime}$.
-The results for varying $\gamma$ are presented and discussed in Section C.1, where we also observe condPED-ANOVA consistently shows reasonable behavior.
+Because the domains of $x$ and $y$ overlap, objective values are no longer cleanly separated by the gating variable $c$.
+Consequently, around $\gamma^{\prime}\approx 0.5$ the HPI of $c$ is no longer close to one, and both $x$ and $y$ also have non-negligible HPIs.
+As $\gamma^{\prime}$ decreases, the contribution of $y$ persists until about $\gamma^{\prime}\approx 0.2$, reflecting the best values attainable in the $y$-branch (near y=2y=-2).
+For further tighter targets, top-$\gamma^{\prime}$ performance requires the $c<0.5$ branch, consistent with the (near) zero importance of $y$.
+Within this branch, achieving higher performance increasingly depends on the precise value of $x$, leading to a growing HPI for $x$ as $\gamma^{\prime}$ decreases.
+Also, under both domain settings, the resulting HPI is approximately symmetric for $\gamma^{\prime}<0.5$ and $\gamma^{\prime}>0.5$, reflecting the objective’s top–bottom symmetry.
+However, for $\gamma^{\prime}>0.5$, changing $c$ can still improve the objective, so $c$ retains a non-zero HPI even near $\gamma^{\prime}\approx 1.0$, reflecting a slight asymmetry caused by minimization.
+These results suggest that condPED-ANOVA yields sensible HPIs for $c$, $x$, and $y$ across target levels $\gamma^{\prime}$.
+The results for varying $\gamma$ are presented and discussed in Section D.2, where condPED-ANOVA consistently shows reasonable behavior.
 5.3. Regime-Dependent Domains
 5.3.1. Problem Setting
-We consider a synthetic optimization problem in which the domains of hyperparameters change depending on a gating variable.
-The objective is to minimize:
+We next consider a synthetic optimization problem where hyperparameter domains change with a gating variable. The objective is to minimize:
 (15) $f(c,x,y)=x+y,$
@@
 (16) $c\in[0,1],~\begin{cases}x\in[-7,-2],~y\in[-5,-2],&\text{if }c<0.5,\\ x\in[2,7],~y\in[2,5],&\text{if }c\geq 0.5.\end{cases}$
-Here, both $x$ and $y$ are always present, but their feasible ranges switch according to $c$.
+Here, both $x$ and $y$ are always active, but their domains switch according to $c$. Additional experiments combining conditional activation with regime-dependent domains are provided in Section E.2.
 5.3.2. Results and Discussion
 The results are shown in Figure 1.
-Similar to the disjoint-domain setting (Equation 7) in the conditional activation experiment (Section 5.2), the gating variable $c$ determines whether a sample achieves top-half performance, which is reflected by the high importance of $c$ around $\gamma^{\prime}\approx 0.5$.
-On the other hand, unlike the conditional activation case, both $x$ and $y$ are always active and contribute to the objective, so their importance grows toward more extreme targets (smaller or larger $\gamma^{\prime}$).
+Similar to the disjoint-domain setting (Equation 13) in the conditional activation experiment (Section 5.2), the gating variable $c$ largely determines top-half performance, yielding high importance near $\gamma^{\prime}\approx 0.5$.
+Unlike that case, however, both $x$ and $y$ are always active and contribute to the objective, consistent with their importance grows toward more extreme targets (smaller or larger $\gamma^{\prime}$).
 Moreover, $x$ is consistently more important than $y$, reflecting its ability to attain more extreme values (around $\pm 7$).
 5.4. Comparison with Existing Methods
-(a) Conditional activation (Equation 6) with disjoint domains (Equation 7)
-(b) Conditional activation (Equation 6) with overlapping domains (Equation 8)
-(c) Regime-dependent domains (Equation 9 and Equation 10)
-Figure 3.
-Ablation study results obtained using standard local HPI instead of conditional local HPI we defined in Section 4.2.
-The lines denote the mean, and the shaded regions denote the standard error, both computed over ten independent runs with different random seeds.
 5.4.1. Baselines
-We compare condPED-ANOVA with existing HPI estimators.
-Since none of the standard HPI methods are designed to handle conditional hyperparameters in a principled manner, we evaluate them under common naive extensions that enable their application to conditional search spaces.
+We then compare condPED-ANOVA with existing HPI estimators.
+Since these HPI estimators do not handle conditional hyperparameters in a principled way, we evaluate them using common naive extensions for conditional search spaces.
 We consider the following four baselines:
@@
   • Filtering: From the evaluation set $\{(\bm{x}_{n},f(\bm{x}_{n}))\}_{n=1}^{N}$, we keep only samples in which the target hyperparameter is present, and compute HPI on the filtered subset.
-  • Imputation: For samples in which the target hyperparameter is inactive, we impute its value by a default value (the midpoint of its domain) and compute HPI on the resulting completed dataset.
+  • Imputation: For samples where the target hyperparameter is inactive, we impute it with a default value (the domain midpoint) and compute HPI on the completed dataset.
 For hyperparameters whose domain switches across regimes, we use the following naive extension:
-  • Expansion: We treat the hyperparameter as if it were sampled from a single expanded domain that covers all regime-specific ranges, and compute HPI on this unified representation.
-The pseudocodes of these naive extensions are provided in Section D.2.
-Consequently, for conditional activation, we obtain $4\times 2=8$ baseline variants (four HPI methods combined with two naive handling schemes), whereas for regime-dependent domains, we obtain $4\times 1=4$ variants (four HPI methods combined with the expansion scheme), which we compare against condPED-ANOVA.
-See Section D.2 for further details on the experimental setup.
-5.4.2. Results and Discussion
-We first analyze the results on the conditional activation objective (Equation 6) under the disjoint-domain setting (Equation 7).
-As shown in Figure 2, applying simple filtering to the original PED-ANOVA leads to almost identical HPI values for $x$ and $y$.
-This behavior is clearly undesirable: in order to achieve better performance (i.e., smaller $\gamma^{\prime}$), the hyperparameter $y$ is inactive, and tuning $y$ cannot affect the objective value at all.
-This failure arises because filtering discards information about how the presence or absence of the variable relates to the overall performance landscape.
-When imputation is applied instead, PED-ANOVA assigns a higher HPI to $x$ in the small-$\gamma^{\prime}$ regime, which is qualitatively reasonable.
-However, the method still assigns a non-negligible importance to $y$ and it becomes comparable to that of $c$, even though $y$ should be inactive in this region.
-Moreover, for $\gamma^{\prime}>0.5$, the HPI of $x$ and $y$ exhibit substantial instability.
-These behaviors are caused by forcibly imputing an artificial value (the midpoint of the domain) for inactive configurations, which distorts the empirical sample distribution and introduces spurious dependencies.
-Furthermore, in both filtering and imputation extensions of PED-ANOVA, the hyperparameter $c$, which solely determines whether the objective value belongs to the top-performing half, is not properly identified as important.
-Around $\gamma^{\prime}\approx 0.5$, all hyperparameters are assigned nearly identical HPI values, indicating a complete failure to capture the dominant role of $c$ in controlling performance.
-Similar issues are observed for the other baseline estimators.
-For f-ANOVA and MDI with filtering, all hyperparameters receive identical importance values, while all the remaining variants still assign identical HPI values to $x$ and $y$.
-These results demonstrate that naive extensions are insufficient for adapting existing HPI estimators to conditional search spaces and fail to produce meaningful or interpretable importance estimates.
-In contrast, condPED-ANOVA consistently yields meaningful importance values that reflect the underlying conditional structure (as shown in Section 5.2).
-Further discussion on the other objectives and domain settings is deferred to Section C.2, where we observe the baselines fail to yield meaningful HPI in conditional settings.
-5.5. Ablation Study: Effect of Within-Regime Variance and Regime Weighting
-In this section, we conduct ablation studies to examine the impact of key design choices in condPED-ANOVA, focusing on within-regime variance and regime-level weighting.
-5.5.1. Effect of Within-Regime Variance for HPI
-In Section 4.2, we defined the conditional local HPI using the within-regime local marginal variance $v_{\gamma,\mathrm{within}}^{(d)}$.
-We therefore compare the resulting HPI estimates with those obtained using the standard local marginal variance $v_{\gamma}^{(d)}$.
-The closed-form expression for the standard local marginal variance, computed using the indicator function $b_{\gamma^{\prime}}\coloneqq\mathbf{1}\{x\in\mathcal{X}_{\gamma^{\prime}}\}$, is given as follows, where we can also derive a decomposition of the Pearson divergence into within-regime and inter-regime components (cf.
-Lemma E.2 in Section E.3):
-(11) $v_{\gamma}^{(d)}=\left(\frac{\gamma^{\prime}}{\gamma}\right)^{\!2}\Bigg($ $\underbrace{\!\sum_{i=1}^{\;K^{(d)}}\!\!\frac{\big(\alpha_{i}^{(d)}\big)^{2\!}}{\beta_{i}^{(d)}}D_{\mathrm{PE}}\!\left(p_{\gamma^{\prime},\,i}^{(d)}\middle\|p_{\gamma,\,i}^{(d)}\!\right)}_{\text{within-regime divergence}}$
-$\hskip 30.00005pt+\underbrace{D_{\mathrm{PE}}\!\left(\!\bm{\alpha}^{(d)}\middle\|\bm{\beta}^{(d)}\!\right)}_{\text{inter-regime divergence}}\Bigg),$
-where $\bm{\alpha}^{(d)}\coloneqq(\alpha_{1}^{(d)},\ldots,\alpha_{K^{(d)}}^{(d)})$ and $\bm{\beta}^{(d)}\coloneqq(\beta_{1}^{(d)},\ldots,\beta_{K^{(d)}}^{(d)})$.
+  • Expansion: We treat the hyperparameter as if it were sampled from a single expanded domain covering all regime-specific ranges, and compute HPI on this unified domain.
+The pseudocodes of these naive extensions are provided in Section C.2.
+Consequently, for conditional activation we evaluate 8 baseline variants (4 HPI methods $\times$ 2 handling schemes), while for regime-dependent domains we evaluate 4 variants (4 methods $\times$ 1 expansion scheme), and compare them against condPED-ANOVA.
+See Section C.2 for further details on the experimental setup.
 (a) Conditional activation (Equation 12) with disjoint domains (Equation 13)
@@
 Figure 3.
-Ablation study results obtained by using the naive aggregation scheme (Equation 12) instead of our closed-form expression (Theorem 4.2).
+Ablation study results using standard local HPI instead of our conditional local HPI.
+The lines denote the mean, and the shaded regions denote the standard error, both computed over ten independent runs with different random seeds.
+(a) Conditional activation (Equation 12) with disjoint domains (Equation 13)
+(b) Conditional activation (Equation 12) with overlapping domains (Equation 14)
+(c) Regime-dependent domains (Equation 15 and Equation 16)
+Figure 4.
+Ablation study results using the naive aggregation scheme (Equation 18) instead of our expression (Theorem 4.2).
 The bottom row shows the HPIs before normalization, i.e., $\tilde{v}_{\gamma,\mathrm{within}}^{(d)}$.
 The lines denote the mean, and the shaded regions denote the standard error, both computed over ten independent runs with different random seeds.
+5.4.2. Results and Discussion
+We first analyze the results on the conditional activation objective (Equation 12) under the disjoint-domain setting (Equation 13), which are shown in Figure 2.
+As shown in this figure, applying simple filtering to PED-ANOVA leads to almost identical HPIs for $x$ and $y$.
+This behavior is clearly undesirable: in order to achieve better performance (i.e., smaller $\gamma^{\prime}$), $y$ is inactive, and tuning $y$ cannot affect the objective at all.
+This failure arises because filtering obscures how the variable relates to performance over the full search space.
+When imputation is applied instead, the HPIs of $x$ and $y$ exhibit substantial instability for $\gamma^{\prime}>0.5$.
+These behaviors are caused by forcibly imputing an artificial value for inactive samples, which distorts the sample distribution and introduces spurious dependencies.
+Furthermore, under both filtering and imputation, PED-ANOVA fails to identify $c$ as important: around $\gamma^{\prime}\approx 0.5$, it assigns nearly identical HPIs to all hyperparameters, even though $c$ largely determines whether a sample falls in the top-performing half.
+Similar issues are observed for the other baselines.
+For f-ANOVA and MDI with filtering, all hyperparameters receive identical HPIs, while all the remaining variants assign identical HPIs to $x$ and $y$.
+These results show that naive extensions are insufficient for applying existing HPI estimators to conditional search spaces, often yielding uninterpretable HPIs.
+In contrast, condPED-ANOVA consistently yields meaningful HPIs that reflect the underlying conditional structure (as shown in Section 5.2).
+Additional results for other objectives and domain settings are deferred to Section D.4, where the baselines similarly fail in conditional settings.
+5.5. Ablation Study: Effect of Within-Regime Variance and Regime Weighting
+In this section, we conduct ablation studies to examine the impact of key design choices in condPED-ANOVA.
+5.5.1. Effect of Within-Regime Variance for HPI
+In Section 4.2, we defined the conditional local HPI using the within-regime local marginal variance $v_{\gamma,\mathrm{within}}^{(d)}$.
+We therefore compare the resulting HPI estimates with those obtained using the standard local marginal variance $v_{\gamma}^{(d)}$.
+The closed-form estimation for the standard local marginal variance, computed using the indicator function $b_{\gamma^{\prime}}\coloneqq\mathbf{1}\{x\in\mathcal{X}_{\gamma^{\prime}}\}$, is given as follows, where we can also derive a decomposition of the Pearson divergence into within-regime and inter-regime components (cf.
+Lemma F.2 in Section F.3):
+(17) $v_{\gamma}^{(d)}\!=\left(\frac{\gamma^{\prime}}{\gamma}\right)^{\!2}\!\Bigg(\underbrace{\!\sum_{i=1}^{\;K^{(d)}}\!\!\frac{\big(\alpha_{i}^{(d)}\big)^{2\!}}{\beta_{i}^{(d)}}D_{\mathrm{PE}}\!\left(p_{\gamma^{\prime},\,i}^{(d)}\middle\|p_{\gamma,\,i}^{(d)}\!\right)}_{\text{within-regime divergence}}+\hskip-3.0pt\underbrace{D_{\mathrm{PE}}\!\left(\bm{\alpha}^{(d)}\middle\|\bm{\beta}^{(d)}\right)}_{\text{inter-regime divergence}}\hskip-3.0pt\Bigg),\!$
+where $\bm{\alpha}^{(d)}\coloneqq(\alpha_{1}^{(d)},\ldots,\alpha_{K^{(d)}}^{(d)})$ and $\bm{\beta}^{(d)}\coloneqq(\beta_{1}^{(d)},\ldots,\beta_{K^{(d)}}^{(d)})$.
 The results obtained using the standard local marginal variance are shown in Figure 3.
-In both the conditional-activation setting with disjoint domains (Figure 3) and the regime-dependent domain setting (Figure 3), the HPI of the gating variable $c$ is underestimated around $\gamma^{\prime}\approx 0.5$ and becomes lower than those of $x$ and $y$.
-This behavior is undesirable, since selecting the appropriate branch is crucial for achieving median-level performance.
-Moreover, in the conditional-activation setting (Figures 3 and 3), the importance of $x$ or $y$ does not vanish even in the regions where the variable is inactive and is assigned non-zero HPIs, which are even larger than that of $c$ (e.g., $y$ for $\gamma^{\prime}<0.5$ and $x$ for $\gamma^{\prime}>0.5$ in Figure 3, and $y$ for $\gamma^{\prime}\lesssim 0.2$ and $x$ for $\gamma^{\prime}\gtrsim 0.8$ in Figure 3).
+In both the conditional-activation with disjoint domain setting (Figure 3) and the regime-dependent domain setting (Figure 3), the HPI of the gating variable $c$ is underestimated around $\gamma^{\prime}\approx 0.5$ and even falls below those of $x$ and $y$, despite branch selection being crucial for median-level performance.
+Moreover, in the conditional-activation setting (Figures 3 and 3), $x$ and $y$ retain non-zero HPI even where they are inactive (e.g., $y$ for $\gamma^{\prime}<0.5$ and $x$ for $\gamma^{\prime}>0.5$ in Figure 3).
 This happens because the inter-regime variance is primarily induced by the gating variable $c$.
-When HPI is computed using the standard local marginal variance that includes this inter-regime component, the effect of $c$ leaks into the HPI estimates for $x$ and $y$.
-As a result, $c$ is suppressed by $x$ and $y$ in the region where $c$ should be most important, and $x$ and $y$ receive spurious non-zero importance even in the regions where they should be inactive and have zero effect.
-This phenomenon can also be established mathematically.
-Theorem E.1 shows that, under the standard local marginal variance, the variance of a conditioned hyperparameter includes the gating variance as an additive term, which explains the observed leakage and the resulting misattribution of importance.
-Therefore, in conditional settings, it is essential to compute HPI using the within-regime local marginal variance $v_{\gamma,\mathrm{within}}^{(d)}$, rather than the standard local marginal variance $v_{\gamma}^{(d)}$.
+When HPI is computed using the standard local marginal variance that includes this inter-regime component, the effect of $c$ leaks into the HPI of $x$ and $y$, suppressing $c$ and assigning spurious HPIs to $x$ and $y$.
+In fact, we can theoretically show that, under the standard local marginal variance, the variance of a conditioned hyperparameter includes the gating variance as an additive term, which explains the observed leakage and the resulting misattribution of importance (cf.
+Theorem F.1).
+Therefore, in conditional settings, it is essential to compute HPI using the within-regime local marginal variance $v_{\gamma,\mathrm{within}}^{(d)}$ rather than the standard local marginal variance $v_{\gamma}^{(d)}$.
 5.5.2. Effect of Regime-Level Weighting for HPI
-In Section 4.3, we derived a closed-form computation method for $v_{\gamma,\mathrm{within}}^{(d)}$ based on a mathematical derivation (Theorem 4.2).
+In Section 4.3, we derived a closed-form method to compute $v_{\gamma,\mathrm{within}}^{(d)}$ (Theorem 4.2).
 However, one might wonder whether such a derivation is necessary in practice, and instead consider a naive alternative that simply sums the within-regime Pearson divergences, i.e., using:
-(12) $\tilde{v}_{\gamma,\mathrm{within}}^{(d)}=\left(\frac{\gamma^{\prime}}{\gamma}\right)^{\!2}\sum_{i=1}^{K^{(d)}}D_{\mathrm{PE}}\!\left(p_{\gamma^{\prime},\,i}^{(d)}\,\middle\|\,p_{\gamma,\,i}^{(d)}\right)$
-in place of Equation 5. In this section, we empirically compare condPED-ANOVA against this naive aggregation scheme.
+(18) $\tilde{v}_{\gamma,\mathrm{within}}^{(d)}\coloneqq\left(\frac{\gamma^{\prime}}{\gamma}\right)^{\!2}\sum_{i=1}^{K^{(d)}}D_{\mathrm{PE}}\!\left(p_{\gamma^{\prime},\,i}^{(d)}\,\middle\|\,p_{\gamma,\,i}^{(d)}\right)$
+in place of Equation 11.
+Figure 5.
+condPED-ANOVA HPIs on a real-world problem from YAHPO Gym rbv2_super scenario (instance ID 1053) (Pfisterer et al., 2022; Binder et al., 2020).
+The task is model selection over ranger, aknn, svm, xgboost, rpart, and glmnet: the hyperparameter learner_id selects the learner, activating only the corresponding learner-specific hyperparameters (prefixed by the learner name).
+The bars denote the mean, and the error bars denote the standard error, both computed over ten independent runs with different random seeds.
 The results obtained using this naive aggregation scheme are shown in Figure 4.
-In this case, the resulting HPI curves become noticeably less stable and exhibit near-discontinuous changes around regime-switching boundaries.
-In particular, for the conditional-activation objective under the overlapping-domain setting (Equations 6 and 8), the importance of $y$ spikes immediately when $y$ becomes active, excessively dominating the HPI and suppressing the importances of $x$ and $c$, which should still contribute to the objective.
-Moreover, although the objective is symmetric between the top region ($\gamma^{\prime}<0.5$) and the bottom region ($\gamma^{\prime}>0.5$) across all settings, the naive scheme breaks this symmetry too severely, yielding clearly distorted importance patterns.
-This instability is largely driven by abrupt scale changes at regime boundaries.
-As shown in the bottom row of Figure 4, $\tilde{v}_{\gamma,\mathrm{within}}^{(d)}$ for $x$ and $y$ exhibit implausible jumps at these boundaries, in sharp contrast to the regime-probability-weighted condPED-ANOVA (cf.
-Figure 9 in Section C.4).
-This failure occurs because removing the regime-probability ratios destroys the frequency-aware weighting across regimes, leading to an inappropriate aggregation that cannot properly account for regime prevalence.
+The resulting HPI curves become noticeably less stable, with near-discontinuous jumps at regime-switching boundaries.
+In particular, for the conditional-activation objective under the overlapping-domain setting (Figure 4), the importance of $y$ spikes immediately when $y$ becomes active, excessively dominating the HPI and suppressing $x$ and $c$, which should still contribute to the objective.
+Moreover, although the objective is symmetric between the top ($\gamma^{\prime}<0.5$) and bottom ($\gamma^{\prime}>0.5$) region across all settings, the naive scheme breaks this symmetry too severely, yielding clearly distorted importance patterns.
+This instability is driven by abrupt scale changes at regime boundaries.
+As shown in the bottom row of Figure 4, $\tilde{v}_{\gamma,\mathrm{within}}^{(d)}$ for $x$ and $y$ exhibit implausible jumps at these boundaries, in contrast to the regime-probability-weighted condPED-ANOVA (cf.
+Figure 11 in Section D.5).
+This failure occurs because removing the regime-probability ratios destroys the frequency-aware weighting, leading to an inappropriate aggregation that ignores regime prevalence.
+6. Experiments on Real-World Benchmarks
+6.1. Experimental Setup
+We validate that condPED-ANOVA works properly in real-world settings using YAHPO Gym (Pfisterer et al., 2022).
+We use the rbv2_super scenario (Binder et al., 2020), as it defines a conditional search space in a single-objective setting.
+We use the instance ID 1053 from the benchmark suites proposed with YAHPO Gym.
+This benchmark represents a model selection problem over six learners: aknn (approximate kkNN), glmnet (regularized linear model), ranger (random forest), rpart (decision tree), svm (support vector machine), and xgboost (gradient-boosted trees).
+A categorical hyperparameter learner_id selects the learner, and only the corresponding learner-specific hyperparameters are active under that choice.
+We use a sample size of $N=1000$ and maximize the accuracy (acc) as the objective.
+We use $\gamma=1.0$ and $\gamma^{\prime}=0.1$.
+Further details of the experimental setup, including details of the YAHPO Gym, are provided in Section C.3.
+Table 1.
+Summary statistics of the objective value (acc) for each learner_id in rbv2_super (instance ID 1053).
+Values are reported as mean $\pm$ standard error, both computed over 10 independent runs with different random seeds.
+The learner_ids are sorted from best to worst max performance.
+Learner ID Min Mean Max
+svm $0.727\pm 0.004$ $0.796\pm 0.001$ $0.926\pm 0.005$
+xgboost $0.591\pm 0.011$ $0.765\pm 0.001$ $0.853\pm 0.004$
+rpart $0.736\pm 0.001$ $0.774\pm 0.001$ $0.846\pm 0.003$
+glmnet $0.731\pm 0.001$ $0.778\pm 0.001$ $0.842\pm 0.003$
+ranger $0.738\pm 0.002$ $0.777\pm 0.000$ $0.815\pm 0.002$
+aknn $0.174\pm 0.001$ $0.591\pm 0.006$ $0.809\pm 0.002$
+6.2. Results and Discussion
+The results are shown in Figure 5.
+We observe that trainsize and learner_id, as well as svm- and xgboost-related hyperparameters, have relatively large HPIs, indicating that these factors are most influential.
+For reference, we report summary statistics of the objective values (acc) for each selected learner in Table 1.
+The table shows large differences in attainable performance across learners, making it reasonable for larger HPI on learner_id.
+Moreover, the ordering of learners that appear in the computed HPIs (via learner-specific hyperparameters) in Figure 5 largely matches the learners’ performance ordering in Table 1 (e.g., svm higher and aknn lower), suggesting that condPED-ANOVA correctly captures each learner’s relative impact on performance.
+Furthermore, within each learner, condPED-ANOVA assigns high HPIs to hyperparameters that are empirically known to be important, such as xgboost’s booster, nrounds and subsample, and svm’s cost, which is reasonable.
+These results indicate that condPED-ANOVA can effectively identify important hyperparameters in real-world model selection problems with conditional hyperparameters.
+Experiments on other problems in YAHPO Gym are reported in Appendix G, where condPED-ANOVA constantly produces reasonable HPIs.
 7. Conclusion
-We studied HPI estimation in conditional search spaces, where hyperparameters may appear, disappear, or change domain depending on other choices.
-We identified a fundamental pitfall of standard local HPI: conditioning effects can leak into conditioned hyperparameters, leading to misleading importances.
-To address this, we defined conditional local HPI and proposed condPED-ANOVA, a closed-form, PED-ANOVA-style estimator that correctly handles conditional search spaces while retaining efficiency.
-Empirically, condPED-ANOVA yielded meaningful importances consistent with the underlying conditional structure, whereas naive extensions of existing estimators, including PED-ANOVA, often violated conditional semantics and produced distorted importances.
+We studied HPI estimation in conditional search spaces, where hyperparameter activation and domains depend on others.
+We identified a key pitfall of standard local HPI: conditioning effects can leak into conditioned hyperparameters, leading to misleading HPIs.
+To address this, we introduced conditional local HPI and proposed condPED-ANOVA, a closed-form, PED-ANOVA-style estimator that correctly handles conditionality while retaining efficiency.
+Experiments showed that condPED-ANOVA produces meaningful HPIs under conditional search spaces, whereas naive extensions of existing estimators often yield misleading HPIs.
 We believe condPED-ANOVA serves as a practical and principled foundation for reliable HPI analysis in HPO workflows with conditional search spaces.
@@
 (2017) accelerate this paradigm via ablation with surrogates, replacing expensive evaluations with model-based predictions to achieve large speedups while maintaining explanatory utility.
-To further focus on neighborhood behavior, CAVE (Biedenkapp et al., 2019) introduces local parameter importance, which measures variance in predicted performance when varying one parameter around a specific configuration, and empirically argues that local and global importance can disagree and thus complement each other.
+To further focus on neighborhood behavior, CAVE (Biedenkapp et al., 2019) introduces local parameter importance, which measures variance in predicted performance when varying one parameter around a specific configuration, and empirically argues that local and global importance can disagree and complement each other.
 A.2. Importance Across Datasets and Tunability
-Another line of work studies importance beyond a single dataset or run.
+Another line of work studies the importance beyond a single dataset or run.
 van Rijn and Hutter (2018) aggregates evidence across many datasets (via OpenML metadata) to identify which hyperparameters tend to matter most for an algorithm in general, and to infer priors over promising values; this reframes HPI as a meta-learning problem and supports more data-driven space design and warm-starting.
@@
 To capture both regime information and regime-specific domain in a single one-dimensional object, we introduce an extended domain for the $d$-th hyperparameter as:
-$\mathsf{S}^{(d)}\coloneqq\bigsqcup_{i=1}^{\;\;\;K^{(d)}}\left(\{i\}\times\mathcal{Z}_{i}^{(d)}\right),$
+(19) $\mathsf{S}^{(d)}\coloneqq\bigsqcup_{i=1}^{\;\;\;K^{(d)}}\left(\{i\}\times\mathcal{Z}_{i}^{(d)}\right),$
 […]
@@
 For each regime $i$, recall the measurable component $A_{i}^{(d)}\subset\mathsf{S}^{(d)}$ and define the regime probabilities:
-$\alpha_{i}^{(d)}\coloneqq\mu_{\gamma^{\prime}}^{(d)}(A_{i}^{(d)}),\quad\text{and}\quad\beta_{i}^{(d)}\coloneqq\mu_{\gamma}^{(d)}(A_{i}^{(d)}).$
+(20) $\alpha_{i}^{(d)}\coloneqq\mu_{\gamma^{\prime}}^{(d)}(A_{i}^{(d)}),\quad\text{and}\quad\beta_{i}^{(d)}\coloneqq\mu_{\gamma}^{(d)}(A_{i}^{(d)}).$
 Whenever $\alpha_{i}^{(d)}>0$ and $\beta_{i}^{(d)}>0$, define the conditional (restricted) measures:
-$\mu_{\gamma^{\prime},\,i}^{(d)}\coloneqq\mu_{\gamma^{\prime}}^{(d)}(\cdot\mid A_{i}^{(d)}),\quad\text{and}\quad\mu_{\gamma,\,i}^{(d)}\coloneqq\mu_{\gamma}^{(d)}(\cdot\mid A_{i}^{(d)}).$
+(21) $\mu_{\gamma^{\prime},\,i}^{(d)}\coloneqq\mu_{\gamma^{\prime}}^{(d)}(\cdot\mid A_{i}^{(d)}),\quad\text{and}\quad\mu_{\gamma,\,i}^{(d)}\coloneqq\mu_{\gamma}^{(d)}(\cdot\mid A_{i}^{(d)}).$
 If $\alpha_{i}^{(d)}=0$, we do not need to define $\mu_{\gamma^{\prime},\,i}^{(d)}$ since its contribution vanishes in Equation 23 through the factor $\big(\alpha_{i}^{(d)}\big)^{2}$.
@@
 The regime-wise $\mu_{\gamma^{\prime},\,i}^{(d)}\ll\mu_{\gamma,\,i}^{(d)}$ also holds for all regimes with $\beta_{i}>0$, since absolute continuity is preserved under restriction to measurable subsets.
-With these definitions, we obtain the following closed-form expression for the within-regime local marginal variance:
+With these definitions, we obtain the following closed-form estimation for the within-regime local marginal variance.
 Theorem B.1 (condPED-ANOVA (Restated)).
@@
 (23) $v_{\gamma,\mathrm{within}}^{(d)}=\left(\frac{\gamma^{\prime}}{\gamma}\right)^{\!2}\sum_{i:\,\beta_{i}^{(d)}>0}\frac{\big(\alpha_{i}^{(d)}\big)^{2}}{\beta_{i}^{(d)}}\,D_{\mathrm{PE}}\!\left(\mu_{\gamma^{\prime},i}^{(d)}\middle\|\mu_{\gamma,i}^{(d)}\right),$
+By normalizing the variance across all hyperparameters as defined in Equation 8, we obtain the conditional local HPI for the $d$-th hyperparameter.
 Here, the Pearson ($\chi^{2}$) divergence is defined as:
-$D_{\mathrm{PE}}(\nu\|\mu)\coloneqq\int\left(\frac{d\nu}{d\mu}-1\right)^{\!2}\,d\mu,$
+(24) $D_{\mathrm{PE}}(\nu\|\mu)\coloneqq\int\left(\frac{d\nu}{d\mu}-1\right)^{\!2}\,d\mu,$
 for probability measures $\nu$ and $\mu$ with $\nu\ll\mu$.
@@
 If $A_{i}^{(d)}$ corresponds to an inactive configuration (i.e., $\mathcal{Z}_{i}^{(d)}=\{\bot\}$), then $\mu_{\gamma,\,i}^{(d)}$ and $\mu_{\gamma^{\prime},\,i}^{(d)}$ are both degenerate on $A_{i}^{(d)}$, implying $D_{\mathrm{PE}}(\mu_{\gamma^{\prime},\,i}^{(d)}\|\mu_{\gamma,\,i}^{(d)})=0$.
-When $K^{(d)}=1$, we have $\alpha_{1}=\beta_{1}=1$ and Equation 14 reduces to the original PED-ANOVA expression.
+When the hyperparameter is not conditional, we have $K^{(d)}=1$ and $\alpha_{1}=\beta_{1}=1$, so Equation 23 reduces to the original PED-ANOVA expression.
 In practice, the sample size within each regime $A_{i}^{(d)}$ must be large enough for KDE to be stable.
@@
 For any measurable set $B\subset\mathsf{S}^{(d)}$, we have:
-$\mathbb{P}(S\in B,\,Y=1)$ $=\mathbb{P}(Y=1)\,\mathbb{P}(S\in B\mid Y=1)$
+(25) $\mathbb{P}(S\in B,\,Y=1)$ $=\mathbb{P}(Y=1)\,\mathbb{P}(S\in B\mid Y=1)$
 $=\kappa\,\mu_{\gamma^{\prime}}^{(d)}(B),$
 and
-$\mathbb{P}(S\in B)=\mu_{\gamma}^{(d)}(B).$
+(26) $\mathbb{P}(S\in B)=\mu_{\gamma}^{(d)}(B).$
 Since $\mathcal{X}_{\gamma^{\prime}}\subset\mathcal{X}_{\gamma}$, we have $\mu_{\gamma^{\prime}}^{(d)}\ll\mu_{\gamma}^{(d)}$, and thus the Radon–Nikodým derivative $d\mu_{\gamma^{\prime}}^{(d)}/d\mu_{\gamma}^{(d)}$ exists.
@@
 Using Equation 27 and Equation 28, for each $i$ with $\beta_{i}^{(d)}>0$, for $s\in A_{i}^{(d)}$, we obtain:
-$\mathbb{P}(Y=1\mid S=s)\;=\;\kappa\,\frac{\alpha_{i}^{(d)}}{\beta_{i}^{(d)}}\,\frac{d\mu_{\gamma^{\prime},\,i}^{(d)}}{d\mu_{\gamma,\,i}^{(d)}}(s)$
+(29) $\mathbb{P}(Y=1\mid S=s)\;=\;\kappa\,\frac{\alpha_{i}^{(d)}}{\beta_{i}^{(d)}}\,\frac{d\mu_{\gamma^{\prime},\,i}^{(d)}}{d\mu_{\gamma,\,i}^{(d)}}(s)$
 Recall that the within-regime local HPI is defined as the within-regime variance of the conditional probability of being in the tighter top set, i.e.,
-$v_{\gamma,\mathrm{within}}^{(d)}\coloneqq\mathbb{E}_{I}\!\left[\mathrm{Var}\!\left(\mathbb{P}(Y=1\mid S)\,\middle|\,I\right)\right].$
+(30) $v_{\gamma,\mathrm{within}}^{(d)}\coloneqq\mathbb{E}_{I}\!\left[\mathrm{Var}\!\left(\mathbb{P}(Y=1\mid S)\,\middle|\,I\right)\right].$
 Taking expectation over $I$ and using $\mathbb{P}(I=i)=\beta_{i}$ yields:
-$v_{\gamma,\mathrm{within}}^{(d)}$ $=\sum_{i:\,\beta_{i}^{(d)}>0}\beta_{i}^{(d)}\mathrm{Var}_{s\sim\mu_{\gamma,\,i}^{(d)}}\!\left(\kappa\frac{\alpha_{i}^{(d)}}{\beta_{i}^{(d)}}\frac{d\mu_{\gamma^{\prime},\,i}^{(d)}}{d\mu_{\gamma,\,i}^{(d)}}(s)\right)$
-$=\kappa^{2}\sum_{i:\,\beta_{i}>0}\frac{\big(\alpha_{i}^{(d)}\big)^{2}}{\beta_{i}^{(d)}}\mathrm{Var}_{s\sim\mu_{\gamma,\,i}^{(d)}}\!\left(\frac{d\mu_{\gamma^{\prime},\,i}^{(d)}}{d\mu_{\gamma,\,i}^{(d)}}(s)\right).$
+(31) $v_{\gamma,\mathrm{within}}^{(d)}$ $=\sum_{i:\,\beta_{i}^{(d)}>0}\beta_{i}^{(d)}\mathrm{Var}_{s\sim\mu_{\gamma,\,i}^{(d)}}\!\left(\kappa\frac{\alpha_{i}^{(d)}}{\beta_{i}^{(d)}}\frac{d\mu_{\gamma^{\prime},\,i}^{(d)}}{d\mu_{\gamma,\,i}^{(d)}}(s)\right)$
+$=\kappa^{2}\sum_{i:\,\beta_{i}^{(d)}>0}\frac{\big(\alpha_{i}^{(d)}\big)^{2}}{\beta_{i}^{(d)}}\mathrm{Var}_{s\sim\mu_{\gamma,\,i}^{(d)}}\!\left(\frac{d\mu_{\gamma^{\prime},\,i}^{(d)}}{d\mu_{\gamma,\,i}^{(d)}}(s)\right)$
 $=\kappa^{2}\sum_{i:\,\beta_{i}^{(d)}>0}\frac{\big(\alpha_{i}^{(d)}\big)^{2}}{\beta_{i}^{(d)}}\int\left(\frac{d\mu_{\gamma^{\prime},\,i}^{(d)}}{d\mu_{\gamma,\,i}^{(d)}}(s)-1\right)^{2}d\mu_{\gamma,\,i}^{(d)}(s)$
@@
 Here, we used the following equation, which holds since $\mu_{\gamma^{\prime},i}^{(d)}$ is a probability measure:
-$\mathbb{E}_{s\sim\mu_{\gamma,i}^{(d)}}\left[\frac{d\mu_{\gamma^{\prime},\,i}^{(d)}}{d\mu_{\gamma,\,i}^{(d)}}(s)\right]=\int\frac{d\mu_{\gamma^{\prime},\,i}^{(d)}}{d\mu_{\gamma,\,i}^{(d)}}\,d\mu_{\gamma,\,i}^{(d)}=\mu_{\gamma^{\prime},\,i}^{(d)}(\mathsf{S}^{(d)})=1.$
+(32) $\mathbb{E}_{s\sim\mu_{\gamma,i}^{(d)}}\left[\frac{d\mu_{\gamma^{\prime},\,i}^{(d)}}{d\mu_{\gamma,\,i}^{(d)}}(s)\right]=\int\frac{d\mu_{\gamma^{\prime},\,i}^{(d)}}{d\mu_{\gamma,\,i}^{(d)}}\,d\mu_{\gamma,\,i}^{(d)}=\mu_{\gamma^{\prime},\,i}^{(d)}(\mathsf{S}^{(d)})=1.$
 This completes the proof. ∎
-Appendix C Additional Results and Discussion on Numerical Experiments
-(a) Conditional activation (Equation 6) with disjoint domains (Equation 7)
-(b) Conditional activation (Equation 6) with overlapping domains (Equation 8)
-(c) Regime-dependent domains (Equation 9 and Equation 10)
-Figure 5.
-condPED-ANOVA HPI computed for the synthetic objectives.
-For the objective with conditional activation (Equation 6), the gating hyperparameter $c$ determines which branch is active: $x$ is present only when $c<0.5$, whereas $y$ is present only when $c\geq 0.5$.
-For the objective with regime-dependent domains (Equation 9), the gating hyperparameter $c$ determines the domain of the hyperparameters $x$ and $y$.
-The lines denote the mean, and the shaded regions denote the standard error, both computed over ten independent runs with different random seeds.
-C.1. condPED-ANOVA Results on Synthetic Objectives with Different $\gamma$ Values
-The results of condPED-ANOVA for the objective with conditional activation (Equation 6) and the objective with regime-dependent domains (Equation 9), with $\gamma$ set to $0.75$ and $0.5$, are shown in Figure 5.
-The results share some similarities with the higher-$\gamma^{\prime}$ range in Figure 1, which reflects the fact that varying $\gamma$ changes the region over which HPI is computed.
-However, the behavior is not a simple zoom-in because the definition of the top-$\gamma$ region itself changes.
-In particular, with $\gamma=0.5$ in the disjoint-domain setting (Figure 5, right panel), once we restrict attention to the better half of the evaluations (i.e., the region where $c<0.5$ is already selected), adjusting $c$ no longer changes the objective value.
-Accordingly, the importance of $c$ becomes small in this regime, similar to the importance of $y$.
-C.2. Comparison with Existing Methods
-Figure 6.
-Baseline HPIs computed with naive extensions of existing methods for the synthetic objective with conditional activation (Equation 6) under the overlapping domain setting (Equation 8).
-The lines and bars denote the mean, and the shaded regions and error bars denote the standard error, both computed over ten independent runs with different random seeds.
-Figure 7.
-Baseline HPIs computed with naive extensions of existing methods for the synthetic objective with regime-dependent domains (Equation 9 and Equation 10).
-The lines and bars denote the mean, and the shaded regions and error bars denote the standard error, both computed over ten independent runs with different random seeds.
-The results for the baseline methods described in Section 5.4 for the conditional-activation objective with the overlapping-domain setting (Equations 6 and 8) and the regime-dependent domain setting (Equations 9 and 10) are shown in Figures 6 and 7, respectively.
-First, for the conditional-activation objective, we observe the same trend as in the disjoint-domain setting (Equations 6 and 7).
-PED-ANOVA with filtering fails to distinguish between $x$ and $y$, while PED-ANOVA with imputation assigns $y$ an importance comparable to $c$ in the small-$\gamma^{\prime}$ regime, where $y$ should be inactive and thus cannot improve performance.
-Similarly, f-ANOVA with filtering and MDI with filtering again yield nearly identical HPI values for all hyperparameters, and the remaining baselines also tend to assign almost the same importance to $x$ and $y$.
-Next, for the regime-dependent domain setting, PED-ANOVA with expansion also behaves undesirably: although $c$ determines whether a configuration falls into the top half, the importance of $c$ around $\gamma^{\prime}\approx 0.5$ becomes smaller than that of $x$ and $y$.
-This failure is caused by expansion, which collapses the regime-specific ranges of $x$ and $y$ into a single unified domain and thereby obscures the effect of range switching, preventing accurate estimation of $x$ and $y$ in the regimes where their domains change.
-Moreover, for f-ANOVA, MDI, and SHAP with expansion, $y$ is assigned larger importance than $x$, even though $x$ should contribute more to the objective.
-This is because forcibly expanding the domain distorts the relationship between the variables and the objective across regimes, leading to misleading HPI estimates.
-These results indicate that the naive extensions do not provide meaningful HPI estimates that capture the difference between the contributions of $x$ and $y$ in this conditional setting, not only for the conditional-activation objective under the disjoint-domain setting (as shown in Section 5.4.2), but also under the overlapping-domain setting and in the regime-dependent domain setting.
-C.3. Runtime Comparison
-Figure 8.
-Runtime comparison against baseline HPI methods with naive extensions for the synthetic objective (Equation 9 and Equation 10).
-The lines denote the mean, and the shaded regions denote the standard error, both computed over ten independent runs with different random seeds.
-We compare the runtime of condPED-ANOVA against baseline HPI methods equipped with naive extensions on the synthetic objective with regime-dependent domains (Equation 9 and Equation 10).
-For each sample size from $2^{9}$ to $2^{17}$, we measure the wall-clock time required to compute HPI, and report the results in Figure 8.
-f-ANOVA is reported only up to $2^{11}$, since it could not be executed at larger sample sizes due to excessive memory consumption.
-Due to its substantial memory consumption, f-ANOVA could not be executed for larger sample sizes and is reported only up to $2^{11}$.
-Similarly, SHAP is reported only up to $2^{15}$ because it was prohibitively slow at larger sample sizes.
-In these results, condPED-ANOVA runs faster than all baseline methods across the evaluated ranges, indicating that the overhead introduced by our approach is small.
-Here, the fact that condPED-ANOVA is faster than PED-ANOVA indicates that the overhead of filtering or imputation is larger than the additional computation introduced by condPED-ANOVA.
-(a) Conditional activation (Equation 6) with disjoint domains (Equation 7)
-(b) Conditional activation (Equation 6) with overlapping domains (Equation 8)
-(c) Regime-dependent domains (Equation 9 and Equation 10)
-Figure 9.
-condPED-ANOVA ($\gamma=1.0$) HPI values before normalization, i.e., $v_{\gamma,\mathrm{within}}^{(d)}$, computed for the synthetic objectives.
-The lines denote the mean, and the shaded regions denote the standard error, both computed over ten independent runs with different random seeds.
-C.4. Ablation Study on the Effect of Regime-Level Weighting
-The ablation study results obtained by using a naive aggregation scheme (Equation 12) instead of the closed-form expression derived in Theorem 4.2 are shown in Figure 4.
-For reference, we also show the HPI values before normalization, i.e., $\tilde{v}_{\gamma,\mathrm{within}}^{(d)}$, in Figure 9.
-Appendix D Detailed Experimental Setup
-D.1. Details for Common Experimental Setup
+Appendix C Detailed Experimental Setup
+C.1. Details for Common Experimental Setup
 For estimating the PDFs $p_{\gamma^{\prime},\,i}^{(d)}$ and $p_{\gamma,\,i}^{(d)}$, we use KDE with Optuna’s default settings, which employ Scott’s rule for bandwidth selection (Scott, 1992).
@@
 We use Python version 3.13.11 and Optuna v4.7.0.
-D.2. Implementation Details for Baselines
+C.2. Implementation Details for Baselines
 For all four baseline methods, we use the implementations provided in Optuna (Akiba et al., 2019) and follow Optuna’s default hyperparameter settings.
 The pseudocode for naive extension for these baselines is provided in Algorithms 2, 3 and 4
-Algorithm 2 Naive Extension: Filtering (Conditional Presence)
+Algorithm 2 Naive Extension: Filtering for Conditional Presence
 1:Evaluation set $\mathcal{D}=\{(\bm{x}_{n},f(\bm{x}_{n}))\}_{n=1}^{N}$, baseline evaluator $\textsc{BaselineHPI}(\cdot)$
@@
 7:  else
-8:   $\text{HPI}\leftarrow d\text{-th value of }\textsc{BaselineHPI}(\mathcal{D}_{\mathrm{act}}^{(d)})$
+8:   $\text{HPI}^{(d)}\leftarrow d\text{-th value of }\textsc{BaselineHPI}(\mathcal{D}_{\mathrm{act}}^{(d)})$
 9:  end if
@@
 11:return $\{\text{HPI}^{(d)}\}_{d=1}^{D}$
-Algorithm 3 Naive Extension: Imputation (Conditional Presence)
+Algorithm 3 Naive Extension: Imputation for Conditional Presence
 1:Evaluation set $\mathcal{D}=\{(\bm{x}_{n},f(\bm{x}_{n}))\}_{n=1}^{N}$, hyperparameter domains $\mathcal{X}^{(1)}\times\cdots\times\mathcal{X}^{(D)}=[\ell^{(1)},u^{(1)}]\times\cdots\times[\ell^{(D)},u^{(D)}]$, baseline evaluator $\textsc{BaselineHPI}(\cdot)$
@@
 […]
-Algorithm 4 Naive Extension: Expansion (Regime-Specific Ranges)
+Algorithm 4 Naive Extension: Expansion for Regime-Dependent Domains
 1:Evaluation set $\mathcal{D}=\{(\bm{x}_{n},f(\bm{x}_{n}))\}_{n=1}^{N}$, regime-specific domains for each $d$-th hyperparameter $\{\mathcal{X}^{(d)}_{i}=[\ell^{(d)}_{i},u^{(d)}_{i}]\}_{i=1}^{K^{(d)}}$, baseline evaluator $\textsc{BaselineHPI}(\cdot)$
@@
 13:return $\textsc{BaselineHPI}(\tilde{\mathcal{D}},d)$
-Appendix E Supplementary Analysis for the Ablation Study Using Standard Local HPI
-E.1. Leakage of Gating Effects in Standard Local HPI
+C.3. Detailed Setup for the Experiments on Real-World Benchmarks (Section 6)
+We use YAHPO Gym (Pfisterer et al., 2022) as a surrogate-based benchmark suite for HPO, which provides fast evaluations for diverse search spaces and datasets.
+In YAHPO Gym, a scenario specifies a specific search space and a set of target metrics, while different instances correspond to different datasets/tasks within that scenario.
+YAHPO Gym includes the following scenario families:
+  • rbv2: Classical ML pipelines with mixed and often hierarchical/conditional hyperparameter spaces across many OpenML datasets, supporting multi-fidelity via training-size/fraction.
+  • nb301: NAS-Bench-301 (Zela et al., 2022) based neural architecture search with a high-dimensional categorical space and epoch-based fidelity, targeting validation accuracy and runtime.
+  • lcbench: Learning-curve benchmark over many OpenML (Vanschoren et al., 2014) tasks with epoch-based fidelity, exposing multiple validation/test performance metrics and training time.
+  • iaml: Interpretable AutoML scenarios that combine predictive performance with resource-usage and interpretability objectives, with multi-fidelity via training-size fractions.
+Among these, we use the rbv2_super scenario because it is single-objective and contains conditional search spaces.
+We follow the benchmark suites proposed with YAHPO Gym, where instances are selected based on surrogate faithfulness and diversity across scenarios; in the single-objective suite, rbv2_super includes six instances (which correspond to OpenML dataset IDs): 1053, 1457, 1063, 1479, 15, and 1468.
+We conduct experiments on all six instances.
+We present results for instance ID 1053 in the main text (Section 6) and report the remaining results in Appendix G.
+The remaining experimental settings are the same as those in the synthetic experiments (Sections 5.1 and C.1).
+Figure 6.
+Runtime comparison against baseline HPI methods with naive extensions for the synthetic objective (Equation 15 and Equation 16).
+The lines denote the mean, and the shaded regions denote the standard error, both computed over ten independent runs with different random seeds.
+Appendix D Additional Results and Discussion on Experiments on Synthetic Problem (Section 5)
+In Section 5, we presented the main results and discussion of experiments on synthetic objectives. In this section, we provide additional results and discussion that complement those in Section 5.
+D.1. Runtime Comparison
+We compare the runtime of condPED-ANOVA against baseline HPI methods on the synthetic objective with regime-dependent domains (Equation 15 and Equation 16).
+For each sample size from $2^{9}$ to $2^{17}$, we measure the wall-clock time required to compute HPI, and report the results in Figure 6.
+f-ANOVA is reported only up to $2^{11}$, since it could not be executed at larger sample sizes due to excessive memory consumption.
+Similarly, SHAP is reported only up to $2^{15}$ because it was prohibitively slow at larger sample sizes.
+In these results, condPED-ANOVA runs faster than all baseline methods across the evaluated ranges, indicating that the overhead introduced by our approach is small.
+Here, the fact that condPED-ANOVA is faster than PED-ANOVA indicates that the overhead of filtering or imputation is larger than the additional computation introduced by condPED-ANOVA.
+(a) Conditional activation (Equation 12) with disjoint domains (Equation 13)
+(b) Conditional activation (Equation 12) with overlapping domains (Equation 14)
+(c) Regime-dependent domains (Equation 15 and Equation 16)
+Figure 7.
+condPED-ANOVA HPI computed for the synthetic objectives with varying $\gamma$.
+The lines denote the mean, and the shaded regions denote the standard error, both computed over ten independent runs with different random seeds.
+(a) Conditional activation (Equation 12) with disjoint domains (Equation 13)
+(b) Conditional activation (Equation 12) with overlapping domains (Equation 14)
+(c) Regime-dependent domains (Equation 15 and Equation 16)
+Figure 8.
+condPED-ANOVA ($\gamma=1.0$) HPI computed for the synthetic objectives with varying NN.
+The lines denote the mean, and the shaded regions denote the standard error, both computed over ten independent runs with different random seeds.
+Figure 9.
+Baseline HPIs computed with naive extensions of existing methods for the synthetic objective with conditional activation under the overlapping domain setting (Equations 12 and 14).
+The lines and bars denote the mean, and the shaded regions and error bars denote the standard error, both computed over ten independent runs with different random seeds.
+D.2. condPED-ANOVA Results with Different $\gamma$
+The results of condPED-ANOVA for the synthetic objectives, with $\gamma$ set to $0.75$ and $0.5$, are shown in Figure 7.
+The results share some similarities with the higher-$\gamma^{\prime}$ range in Figure 1, which reflects the fact that varying $\gamma$ changes the region over which HPI is computed.
+However, the behavior is not a simple zoom-in because the definition of the top-$\gamma$ region itself changes.
+In particular, with $\gamma=0.5$ in the disjoint-domain setting (Figure 7, right panel), once we restrict attention to the better half of the evaluations (i.e., the region where $c<0.5$ is already selected), adjusting $c$ no longer changes the objective value.
+Accordingly, the HPI of $c$ becomes small in this regime, similar to the HPI of $y$.
+D.3. condPED-ANOVA Results with Different NN
+The results of condPED-ANOVA for the synthetic objectives, with $N$ set to $100$ and $500$, are shown in Figure 8.
+From this, we observe that reducing $N$ slightly increases the standard error.
+However, the resulting HPI estimates remain reasonable and are largely consistent with those obtained when $N=1000$.
+These results suggest that condPED-ANOVA is robust even under smaller evaluation budgets.
+D.4. Comparison with Existing Methods
+Figure 10.
+Baseline HPIs computed with naive extensions of existing methods for the synthetic objective with regime-dependent domains (Equation 15 and Equation 16).
+The lines and bars denote the mean, and the shaded regions and error bars denote the standard error, both computed over ten independent runs with different random seeds.
+In Section 5.4, we reported results for baselines computed with naive extensions of existing methods on the conditional-activation objective with the disjoint-domain setting (Equations 12 and 13).
+In this section, we present these baseline results for the conditional-activation objective with the overlapping-domain setting (Equations 12 and 14) and for the regime-dependent domain setting (Equations 15 and 16).
+The results are shown in Figures 9 and 10.
+(a) Conditional activation (Equation 12) with disjoint domains (Equation 13)
+(b) Conditional activation (Equation 12) with overlapping domains (Equation 14)
+(c) Regime-dependent domains (Equation 15 and Equation 16)
+Figure 11.
+condPED-ANOVA ($\gamma=1.0$) HPI values before normalization, i.e., $v_{\gamma,\mathrm{within}}^{(d)}$, computed for the synthetic objectives.
+The lines denote the mean, and the shaded regions denote the standard error, both computed over ten independent runs with different random seeds.
+First, for the conditional-activation objective, we observe the same trend as in the disjoint-domain setting (Equations 12 and 13).
+PED-ANOVA with filtering fails to distinguish between $x$ and $y$, assigning them nearly identical HPI values across all $\gamma^{\prime}$.
+This is clearly problematic, since the contributions of $x$ and $y$ to the objective should differ substantially.
+PED-ANOVA with imputation assigns $y$ an importance comparable to $c$ in the small-$\gamma^{\prime}$ regime, where $y$ should be inactive and thus cannot improve performance; it also exhibits unnatural instability for $\gamma^{\prime}>0.5$.
+f-ANOVA with filtering and MDI with filtering again yield nearly identical HPI values for all hyperparameters, and the remaining baselines also tend to assign almost the same importance to $x$ and $y$, failing to capture the true difference in their contributions to the objective.
+Next, for the regime-dependent domain setting, PED-ANOVA with expansion also behaves undesirably: although $c$ determines whether a configuration falls into the top half, the importance of $c$ around $\gamma^{\prime}\approx 0.5$ becomes smaller than that of $x$ and $y$.
+This failure is caused by expansion, which collapses the regime-specific ranges of $x$ and $y$ into a unified domain, preventing accurate estimation.
+Moreover, for f-ANOVA, MDI, and SHAP with expansion, $y$ is assigned larger importance than $x$, even though $x$ should contribute more to the objective.
+This is because forcibly expanding the domain distorts the relationship between the variables and the objective across regimes, leading to misleading HPI estimates.
+These results indicate that the naive extensions do not provide meaningful HPI estimates, not only for the conditional-activation objective under the disjoint-domain setting (as shown in Section 5.4.2), but also under the overlapping-domain setting and in the regime-dependent domain setting.
+D.5. Ablation Study on the Effect of Regime-Level Weighting
+In Figure 4, we showed the behavior of the $\tilde{v}_{\gamma,\mathrm{within}}^{(d)}$ under the naive aggregation scheme (Equation 18) (bottom row of Figure 4).
+For reference, we plot $v_{\gamma,\mathrm{within}}^{(d)}$ used by condPED-ANOVA (which is normalized to obtain conditional local HPI) in Figure 11.
+From this, we see that condPED-ANOVA applies appropriate regime-probability weighting, which calibrates the scale of $v_{\gamma,\mathrm{within}}^{(d)}$ for each variable and keeps it stable across regime changes without large discontinuities.
+Appendix E Additional Experiments on Synthetic Problems
+In Section 5, we considered the simplest synthetic problems. In this section, we study more complex synthetic settings.
+E.1. Nested Conditional Activation
+Figure 12.
+condPED-ANOVA ($\gamma=1.0$) HPI computed for the synthetic objectives on nested conditional activations (Equations 33 and 34).
+The lines denote the mean, and the shaded regions denote the standard error, both computed over ten independent runs with different random seeds.
+E.1.1. Problem Setting
+We next consider a synthetic problem with a nested conditional search space, where the active variable is selected through a two-level gating hierarchy. The objective function is defined as:
+(33) $f(c_{0},c_{1},x,y,z)=\begin{cases}x,&\text{if }c_{0}<0.75\text{ and }c_{1}<0.5,\\ y,&\text{if }c_{0}<0.75\text{ and }c_{1}\geq 0.5,\\ z,&\text{if }c_{0}\geq 0.75,\end{cases}$
+with the following variable domains:
+(34) $\begin{aligned} c_{0}&\in[0,1],\\ c_{1}&\in[0,1],\end{aligned}\qquad\begin{aligned} x&\in[-7,-4],\\ y&\in[-3,0],\\ z&\in[2,5].\end{aligned}$
+Our goal is to minimize $f$.
+When $c_{0}<0.75$, a second gating variable $c_{1}$ becomes active and selects between $x$ (if $c_{1}<0.5$) and $y$ (if $c_{1}\geq 0.5$); when $c_{0}\geq 0.75$, only the variable $z$ is active, and $c_{1}$, $x$, and $y$ are inactive.
+E.1.2. Results and Discussion
+The results are shown in Figure 12.
+We observe that $c_{0}$ attains the highest HPI around $\gamma^{\prime}\approx 0.75$, taking an importance close to one.
+This is reasonable, as $c_{0}$ induces the primary split of the search space.
+We also see that $c_{1}$ becomes important around $\gamma^{\prime}\approx 0.375$.
+Although its peak is lower than that of $c_{0}$ (around 0.75), this correctly reflects that, conditional on $c_{0}<0.75$, $c_{1}$ provides the next major bifurcation.
+Moreover, the importance shifts to the active variable within each regime: $x$ becomes large for $\gamma^{\prime}<0.375$, $y$ for $0.375<\gamma^{\prime}<0.75$, and $z$ for $\gamma^{\prime}>0.75$, with their HPI increasing as $\gamma^{\prime}$ moves away from the corresponding boundaries.
+These patterns are consistent with the nested conditional structure, indicating that condPED-ANOVA yields sensible importances even in nested conditional settings.
+E.2. Combined Conditional Activation and Domain Shifts in a Three-Way Branching
+E.2.1. Problem Setting
+We next consider a synthetic problem with a three-way branching conditional search space, where the gating variable induces both conditional activation and regime-dependent domain shifts.
+The objective function is defined as:
+(35) $f(c,x,y)=\begin{cases}x,&\text{if }c<1/3,\hskip 26.03424ptx\in[-7,-4],\\ x,&\text{if }1/3\leq c<2/3,\hskip 2.84526ptx\in[-3,0],\\ y,&\text{if }2/3\leq c,\hskip 26.74556pty\in[4,7],\end{cases}\ c\in[0,1].$
+Our goal is to minimize $f$. Here, $x$ is always active when $c<2/3$, but its domain depends on the regime induced by $c$; when $c\geq 2/3$, $y$ becomes active and $x$ is inactive.
+Figure 13.
+condPED-ANOVA ($\gamma=1.0$) HPI computed for the synthetic objectives on the combined conditional activation and domain shifts in a three-way branching setting (Equation 35).
+The lines denote the mean, and the shaded regions denote the standard error, both computed over ten independent runs with different random seeds.
+E.2.2. Results and Discussion
+The results are shown in Figure 13.
+We observe that $c$ exhibits pronounced peaks around $\gamma^{\prime}\approx 0.33$ and $\gamma^{\prime}\approx 0.66$.
+This behavior is reasonable, since it is consistent with the role of $c$ in partitioning the objective-value range into three regimes.
+For $x$, the HPI is large in the regions $\gamma^{\prime}<0.33$ and $0.33<\gamma^{\prime}<0.66$, while it becomes nearly zero for $\gamma^{\prime}>0.66$.
+This pattern reflects that $x$ is influential when $c<1/3$ and when $1/3\leq c<2/3$.
+The near-zero HPI for $\gamma^{\prime}>0.66$ is also reasonable, as it corresponds to the regime where $x$ is inactive.
+Finally, the HPI of $y$ increases only in the $\gamma^{\prime}>0.66$ region and remains close to zero elsewhere.
+This is consistent with the fact that $y$ is active only in the regime $c\geq 2/3$.
+These results suggest that condPED-ANOVA correctly computes HPI for branching hyperparameters such as $c$ that induce three or more regimes, as well as for hyperparameters such as $x$ whose activation and domain both vary across regimes.
+Appendix F Supplementary Analysis for the Ablation Study Using Standard Local HPI
+In this section, we provide additional theoretical details for the ablation study on the effect of within-regime variance for HPI presented in Section 5.5.1.
+In Section F.1, we give a theoretical analysis showing that gating effects can leak when within-regime local marginal variance is not used.
+In addition, Section 5.5.1 compared condPED-ANOVA against a PED-ANOVA-style construction of the standard local HPI using $v_{\gamma}^{(d)}$ as an ablation baseline (Equation 17); we provide its derivation in Section F.3.
+F.1. Leakage of Gating Effects in Standard Local HPI
 Let $b_{\gamma^{\prime}}\coloneqq\mathbf{1}\{\bm{x}\in\mathcal{X}_{\gamma^{\prime}}\}$ be the top-set indicator.
-Consider a gating hyperparameter $x^{(c)}$ and a conditional (child) hyperparameter $x^{(d)}$ whose regime is determined by $x^{(c)}$.
+Consider a conditioning hyperparameter $x^{(c)}$ and a conditioned hyperparameter $x^{(d)}$ whose regime is determined by $x^{(c)}$.
 Formally, assume that for a conditional hyperparameter $x^{(d)}$ there exists a measurable mapping $\phi^{(d)}\colon\mathcal{X}^{(c)}\to\{1,\ldots,K^{(d)}\}$ such that:
@@
 where the expectations are taken under the empirical distribution restricted to the top-$\gamma$ region $\mathcal{X}_{\gamma}$.
-Here, $x^{(c)}$ corresponds to the gating variable $c$ in the synthetic objectives (Equations 6 and 9), while $x^{(d)}$ corresponds to the conditional variables $x$ and $y$.
+Here, $x^{(c)}$ corresponds to the gating variable $c$ in the synthetic objectives (Equations 12 and 15), while $x^{(d)}$ corresponds to the conditioned variables $x$ and $y$.
 These assumptions are satisfied by the synthetic objectives used in our experiments.
 Then, we have the following result regarding the standard local HPI of the conditional hyperparameter $x^{(d)}$.
-Theorem E.1 (Leakage of gating effects under the standard local HPI).
-Under the above assumptions (Equations 17 and 18), the standard local marginal variance of the conditional hyperparameter $x^{(d)}$ contains the local marginal variance of the gating hyperparameter $x^{(c)}$:
+Theorem F.1 (Leakage of gating effects under the standard local HPI).
+Under the above assumptions (Equations 36 and 37), the standard local marginal variance of the conditional hyperparameter $x^{(d)}$ contains the local marginal variance of the gating hyperparameter $x^{(c)}$:
 (38) $v_{\gamma}^{(d)}=v_{\gamma,\mathrm{within}}^{(d)}+v_{\gamma}^{(c)}.$
 In particular, $v_{\gamma}^{(c)}\leq v_{\gamma}^{(d)}$.
-The proof is provided in Section E.2.
+The proof is provided in Section F.2.
 Consequently, when HPI is computed from $v_{\gamma}^{(d)}$, the effect of the gating variable $x^{(c)}$ is additively inherited by conditional hyperparameters through the inter-regime term.
@@
 Moreover, in the regions of $\gamma^{\prime}$ where $x$ or $y$ is inactive, their HPI becomes almost identical to that of $c$.
-This is exactly what Theorem E.1 predicts: once $v_{\gamma,\mathrm{within}}^{(d)}$ vanishes for an inactive variable, the standard local variance $v_{\gamma}^{(d)}$ reduces to the additive gating term $v_{\gamma}^{(c)}$.
-E.2. Proof of Theorem E.1
+This is exactly what Theorem F.1 predicts: once $v_{\gamma,\mathrm{within}}^{(d)}$ vanishes for an inactive variable, the standard local variance $v_{\gamma}^{(d)}$ reduces to the additive gating term $v_{\gamma}^{(c)}$.
+F.2. Proof of Theorem F.1
 Proof.
@@
 $\mathbb{E}_{Z^{(d)}}\left[g_{\gamma}^{(d)}(I^{(d)},Z^{(d)})\,\middle|\,I^{(d)}\right]$ $=\mathbb{E}_{Z^{(d)}}\Big[\mathbb{E}\big[b_{\gamma^{\prime}}\,\big|\,I^{(d)},Z^{(d)}\big]\,\Big|\,I^{(d)}\Big]$
-$=\mathbb{E}\!\left[b_{\gamma^{\prime}}\mid I^{(d)}\right].$
+(40) $=\mathbb{E}\!\left[b_{\gamma^{\prime}}\mid I^{(d)}\right].$
 Thus, the inter-regime variance in Equation 39 equals:
@@
 On the other hand, the local marginal mean of the top-set indicator $b_{\gamma^{\prime}}$ with respect to the gating coordinate $x^{(c)}$ is:
-$f_{\gamma}^{(c)}(x^{(c)})$ $\coloneqq\frac{1}{Z_{\gamma}^{(c)}(x^{(c)})}\sum_{n=1}^{N}b_{\gamma^{\prime}}(x^{(c)},x^{(\bar{c}\,)}_{n})\,b_{\gamma}(x^{(c)},x^{(\bar{c}\,)}_{n})$,
+(42) $f_{\gamma}^{(c)}(x^{(c)})$ $\coloneqq\frac{1}{Z_{\gamma}^{(c)}(x^{(c)})}\sum_{n=1}^{N}b_{\gamma^{\prime}}(x^{(c)},x^{(\bar{c}\,)}_{n})\,b_{\gamma}(x^{(c)},x^{(\bar{c}\,)}_{n})$,
 $\,=\mathbb{E}\left[b_{\gamma^{\prime}}\mid x^{(c)}\right].$
@@
 (44) $\mathrm{Var}_{x^{(c)}}\!\left(\mathbb{E}\!\left[b_{\gamma^{\prime}}\mid I^{(d)}\right]\right)=\mathrm{Var}_{I^{(d)}}\!\left(\mathbb{E}\!\left[b_{\gamma^{\prime}}\mid I^{(d)}\right]\right).$
-Comparing Equations 21, 22 and 23, we conclude that the inter-regime variance of the conditional hyperparameter $x^{(d)}$ equals the marginal variance of the gating hyperparameter $x^{(c)}$:
-$v_{\gamma,\mathrm{inter}}^{(d)}=v_{\gamma}^{(c)}.$
+Comparing Equations 41, 43 and 44, we conclude that the inter-regime variance of the conditional hyperparameter $x^{(d)}$ equals the marginal variance of the gating hyperparameter $x^{(c)}$:
+(45) $v_{\gamma,\mathrm{inter}}^{(d)}=v_{\gamma}^{(c)}.$
 Substituting this into Equation 39 completes the proof. ∎
-E.3. Decomposition of PED over Regimes
+F.3. Decomposition of PED over Regimes
 Here, we use the notation defined in Sections B.1 and B.2.
@@
 We can then show the following decomposition of the Pearson divergence between $\mu_{\gamma}^{(d)}$ and $\mu_{\gamma^{\prime}}^{(d)}$ on $\mathsf{S}^{(d)}$.
-Lemma E.2 (Regime-wise decomposition of Pearson divergence).
+Lemma F.2 (Regime-wise decomposition of Pearson divergence).
 Assume $\mu_{\gamma^{\prime}}^{(d)}\ll\mu_{\gamma}^{(d)}$ and $\mu_{\gamma^{\prime},\,i}^{(d)}\ll\mu_{\gamma,\,i}^{(d)}$ for every $i$ with $\beta_{i}^{(d)}>0$.
@@
 where $\alpha^{(d)}\coloneqq\big(\alpha_{i}^{(d)}\big)_{i=1}^{K^{(d)}}$ and $\beta^{(d)}\coloneqq\big(\beta_{i}^{(d)}\big)_{i=1}^{K^{(d)}}$ are the discrete distributions over regimes.
-The proof is provided in Section E.4.
+The proof is provided in Section F.4.
 The absolute continuity $\mu_{\gamma^{\prime}}^{(d)}\ll\mu_{\gamma}^{(d)}$ holds for the extended marginals in the same way as in Equation 5.
@@
 Using this decomposition, we obtain Equation 17.
-E.4. Proof of Lemma E.2
+F.4. Proof of Lemma F.2
 Proof.
@@
 Since $\mu_{\gamma^{\prime}}\ll\mu_{\gamma}$ and $\mu_{\gamma^{\prime},\,i}\ll\mu_{\gamma,\,i}$ holds for $\beta_{i}>0$, the Radon–Nikodým derivative $d\mu_{\gamma^{\prime}}/d\mu_{\gamma}$ exists and is given by:
-$\frac{d\mu_{\gamma^{\prime}}}{d\mu_{\gamma}}(x)=\frac{d(\alpha_{i}\,\mu_{\gamma^{\prime},\,i})}{d(\beta_{i}\,\mu_{\gamma,\,i})}(x)=\frac{\alpha_{i}}{\beta_{i}}\,\frac{d\mu_{\gamma^{\prime},\,i}}{d\mu_{\gamma,\,i}}(x),$
+(47) $\frac{d\mu_{\gamma^{\prime}}}{d\mu_{\gamma}}(x)=\frac{d(\alpha_{i}\,\mu_{\gamma^{\prime},\,i})}{d(\beta_{i}\,\mu_{\gamma,\,i})}(x)=\frac{\alpha_{i}}{\beta_{i}}\,\frac{d\mu_{\gamma^{\prime},\,i}}{d\mu_{\gamma,\,i}}(x),$
 for $x\in A_{i}$. We denote $s_{i}\coloneqq d\mu_{\gamma^{\prime},\,i}/d\mu_{\gamma,\,i}$.
@@
 $=D_{\mathrm{PE}}\left(\mu_{\gamma^{\prime},\,i}\,\middle\|\,\mu_{\gamma,\,i}\right)+1.$
-By substituting Equations 26 and 27 into Equation 25, we have:
-$D_{\mathrm{PE}}(\mu_{\gamma^{\prime}}\|\mu_{\gamma})$ $=\sum_{i=1}^{\;K^{(d)}}\left(\frac{\alpha_{i}^{2}}{\beta_{i}}\left(D_{\mathrm{PE}}\left(\mu_{\gamma^{\prime},\,i}\,\middle\|\,\mu_{\gamma,\,i}\right)+1\right)-2\alpha_{i}+\beta_{i}\right)$
+By substituting Equations 49 and 50 into Equation 48, we have:
+(51) $D_{\mathrm{PE}}(\mu_{\gamma^{\prime}}\|\mu_{\gamma})$ $=\sum_{i=1}^{\;K^{(d)}}\left(\frac{\alpha_{i}^{2}}{\beta_{i}}\left(D_{\mathrm{PE}}\left(\mu_{\gamma^{\prime},\,i}\,\middle\|\,\mu_{\gamma,\,i}\right)+1\right)-2\alpha_{i}+\beta_{i}\right)$
 $=\sum_{i=1}^{\;K^{(d)}}\left(\frac{\alpha_{i}^{2}}{\beta_{i}}D_{\mathrm{PE}}\left(\mu_{\gamma^{\prime},\,i}\,\middle\|\,\mu_{\gamma,\,i}\right)+\frac{(\alpha_{i}-\beta_{i})^{2}}{\beta_{i}}\right)$
@@
 $=\sum_{i=1}^{\;K^{(d)}}\frac{\alpha_{i}^{2}}{\beta_{i}}D_{\mathrm{PE}}\left(\mu_{\gamma^{\prime},\,i}\,\middle\|\,\mu_{\gamma,\,i}\right)+D_{\mathrm{PE}}\left(\alpha\,\middle\|\,\beta\right).$
-This completes the proof.
+This completes the proof. ∎
+Appendix G Additional Results and Discussion on Experiments on Real-World Benchmarks (Section 6)
+In Section 6, we presented results on the YAHPO Gym rbv2_super scenario (instance ID 1053).
+Here, we report additional results on other instances of the rbv2_super benchmark suite provided by YAHPO Gym.
+The results are shown in Figure 14.
+The corresponding statistics of objective values (acc) for each learner_id are provided in Table 2.
+These results show that, across instances, learner_id, train_size, and hyperparameters associated with the top-performing learners occupy the top ranks in HPI.
+Moreover, the HPIs produced by condPED-ANOVA broadly match the learner-wise performance ordering in Table 2, supporting their validity.
+In instance ID 1468, although aknn attains the highest mean performance, svm and ranger achieve higher maximum performance.
+This is reflected in higher HPIs in svm and ranger, which is reasonable given that our goal is to achieve the highest possible score.
+In instance IDs 1063 and 1468, the performance gaps among the top learners (in terms of mean and maximum accuracy) are smaller than in instance 1457.
+Consistently, for these instances, the HPI of learner_id is slightly lower than that of hyperparameters associated with the top learners, which is reasonable.
+Moreover, across all instances, condPED-ANOVA consistently assigns high HPI to empirically known important hyperparameters within each learner, such as sample.fraction and min.node.size for ranger, the cost and kernel for svm, and k for aknn.
+This is reasonable given their well-known influence on model performance.
+These results demonstrate that condPED-ANOVA consistently yields meaningful and interpretable HPIs across multiple real-world instances of the rbv2_super suite, emphasizing its robustness and applicability in practical scenarios.
+Table 2.
+Summary statistics of the objective value (acc) for each learner_id in other instances (Pfisterer et al., 2022; Binder et al., 2020).
+Values are reported as mean $\pm$ standard error, both computed over 10 independent runs with different random seeds.
+The learner_ids are sorted from best to worst max performance.
+(a) Instance ID 1457
+Learner ID Min Mean Max
+ranger $0.014\pm 0.001$ $0.555\pm 0.004$ $0.832\pm 0.003$
+svm $0.000\pm 0.000$ $0.211\pm 0.008$ $0.803\pm 0.003$
+glmnet $0.007\pm 0.000$ $0.146\pm 0.002$ $0.657\pm 0.019$
+rpart $003\pm 0.000 0.0$$063\pm 0.002$ $0.362\pm 0.006$
+aknn $0.014\pm 0.000$ $0.132\pm 0.002$ $0.354\pm 0.003$
+xgboost $008\pm 0.001 0.0$$046\pm 0.001$ $0.354\pm 0.017$
+(b) Instance ID 1063
+Learner ID Min Mean Max
+svm $0.711\pm 0.006$ $0.816\pm 0.001$ $0.933\pm 0.005$
+xgboost $0.524\pm 0.008$ $0.746\pm 0.001$ $0.902\pm 0.002$
+rpart $0.733\pm 0.002$ $0.818\pm 0.001$ $0.897\pm 0.003$
+ranger $0.727\pm 0.001$ $0.818\pm 0.000$ $0.874\pm 0.002$
+glmnet $0.728\pm 0.003$ $0.802\pm 0.001$ $0.883\pm 0.002$
+aknn $0.187\pm 0.002$ $0.649\pm 0.006$ $0.846\pm 0.001$
+(c) Instance ID 1479
+Learner ID Min Mean Max
+aknn $0.449\pm 0.002$ $0.630\pm 0.003$ $0.987\pm 0.002$
+svm $0.418\pm 0.003$ $0.554\pm 0.003$ $0.820\pm 0.009$
+xgboost $0.406\pm 0.007$ $0.524\pm 0.001$ $0.652\pm 0.006$
+glmnet $0.430\pm 0.002$ $0.513\pm 0.001$ $0.646\pm 0.006$
+rpart $0.444\pm 0.002$ $0.498\pm 0.001$ $0.629\pm 0.008$
+ranger $0.446\pm 0.005$ $0.513\pm 0.001$ $0.583\pm 0.005$
+(d) Instance ID 15
+Learner ID Min Mean Max
+ranger $0.588\pm 0.003$ $0.905\pm 0.002$ $0.985\pm 0.000$
+svm $0.776\pm 0.008$ $0.921\pm 0.001$ $0.980\pm 0.001$
+aknn $0.312\pm 0.001$ $0.721\pm 0.006$ $0.972\pm 0.001$
+rpart $0.648\pm 0.005$ $0.886\pm 0.002$ $0.963\pm 0.001$
+xgboost $0.469\pm 0.007$ $0.714\pm 0.002$ $0.931\pm 0.003$
+glmnet $0.646\pm 0.004$ $0.795\pm 0.001$ $0.929\pm 0.001$
+(e) Instance ID 1468
+Learner ID Min Mean Max
+svm $0.049\pm 0.004$ $0.471\pm 0.010$ $0.968\pm 0.001$
+ranger $0.087\pm 0.003$ $0.782\pm 0.004$ $0.943\pm 0.002$
+aknn $0.252\pm 0.014$ $0.808\pm 0.002$ $0.934\pm 0.002$
+glmnet $0.102\pm 0.002$ $0.404\pm 0.005$ $0.890\pm 0.003$
+rpart $0.083\pm 0.001$ $0.309\pm 0.004$ $0.786\pm 0.007$
+xgboost $0.080\pm 0.003$ $0.206\pm 0.001$ $0.744\pm 0.010$
+(a) Instance ID 1457
+(b) Instance ID 1063
+(c) Instance ID 1479
+(d) Instance ID 15
+(e) Instance ID 1468
+Figure 14.
+condPED-ANOVA HPIs on real-world problems from the YAHPO Gym rbv2_super scenario with other instances (Pfisterer et al., 2022; Binder et al., 2020).
+The bars denote the mean, and the error bars denote the standard error, both computed over ten independent runs with different random seeds.